### v2

Abstract
For $p\in[1,\infty]$, we define a smooth manifold structure on the set $AC_{L^{p}}([a,b],N)$ of absolutely continuous functions $\gamma\colon[a,b]\to N$ with $L^{p}$-derivatives for all real numbers $a<b$ and each smooth manifold $N$ modeled on a sequentially complete locally convex topological vector space, such that $N$ admits a local addition. Smoothness of natural mappings between spaces of absolutely continuous functions is discussed, like superposition operators $AC_{L^{p}}([a,b],N_{1})\to AC_{L^{p}}([a,b],N_{2})$, $\eta\mapsto f\circ\eta$, for a smooth map $f\colon N_{1}\to N_{2}$. For $1\leq p<\infty$ and $r\in\mathbb{N}$ we show that the right half-Lie groups $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$ and $\text{Diff}^{\text{ }r}(M)$ are $L^{p}$-semiregular. Here $K$ is a compact subset of $\mathbb{R}^{n}$ and $M$ is a compact smooth manifold. An $L^{p}$-semiregular half-Lie group $G$ admits an evolution map $\operatorname{Evol}:L^{p}([0,1],T_{e}G)\to AC_{L^{p}}([0,1],G)$, where $e$ is the neutral element of $G$. For the preceding examples, the evolution map $\operatorname{Evol}$ is continuous.
Manifolds of absolutely continuous functions with values in
an infinite-dimensional manifold and regularity
properties of half-Lie groups
Matthieu F. Pinaud
MSC 2020 subject classification: 58D25 (primary); 22E65, 34A12, 46T10, 58D05.
Keywords: Infinite-dimensional manifolds, manifold of mappings, absolutely continuous functions, half-Lie group, regular Lie group, regularity, diffeomorphism group, evolution, initial value problem.
1 Introduction
Manifolds of mappings play an important role in various branches of global analysis, infinite-dimensional geometry and Lie theory. For a compact smooth manifold $M$, a finite-dimensional manifold $N$ and a non-negative integer $r$, a smooth manifold structure on the set $C^{r}(M,N)$ of $C^{r}$-functions $\phi\colon M\to N$ was first obtained by Eells (see [6] and the references therein). Later, manifold structures were also obtained on $C^{\infty}(M,N)$, and for infinite-dimensional manifolds $N$ as long as they admit a local addition, a concept recalled in Definition 1.1 (see [19, 15] and the references therein, also [2]). For a Hilbert manifold $M$, a smooth manifold structure on the set $AC_{L^{2}}([0,1],N)$ of absolutely continuous paths with $L^{2}$-derivatives in local charts was used by Flaschel and Klingenberg to study closed geodesics in Riemannian manifolds (see [7] and [14], cf. also [27]). For Banach manifolds $N$ admitting a smooth local addition and $p\in[1,\infty]$, a smooth manifold structure on $AC_{L^{p}}([0,1],N)$ may also be obtained using a method of Krikoian [16] which is similar to Palais’ use of Banach section functors [25]. A Lie group structure (and hence a smooth manifold structure) on $AC_{L^{p}}([0,1],G)$ was obtained in [23] for each Lie group $G$ modeled on a sequentially complete locally convex space (as in [21]), generalizing the case of Fréchet–Lie groups treated in [10].
We construct manifolds of absolutely continuous functions in higher generality. To formulate the main result, let us fix notation.
Definition 1.1
Let $N$ be a smooth manifold modeled on real a locally convex space, with tangent bundle $TN$ and its bundle projection $\pi_{TN}\colon TN\to N$. A local addition on $N$ is a map
$\Sigma\colon\Omega\to N,$
defined on an open set $\Omega\subseteq TN$ which contains the zero-vector $0_{p}\in T_{p}N$ for each $x\in N$, such that $\Sigma(0_{p})=p$ for all $p\in N$ and
$\theta\colon\Omega\to N\times N,\quad v\mapsto(\pi_{TN}(v),\Sigma(v))$
has open image and is a $C^{\infty}$-diffeomorphism onto its image $\Omega^{\prime}$.
Throughout the following, $a<b$ are real numbers and $p\in[1,\infty]$. For a definition of absolutely continuous functions with values in a sequentially complete locally convex space $E$ or a smooth manifold $N$ defined thereon, the reader is referred to Definitions 2.11 and 2.21, respectively. For $\eta\in AC_{L^{p}}([a,b],N)$, the pointwise operations make
$\Gamma_{AC}(\eta):=\{\tau\in AC_{L^{p}}([a,b],TN)\colon\pi_{TN}\circ\tau=\eta\}$
a vector space; we endow it with a natural topology making it a locally convex topological vector space (cf. Definition 3.1). We shall see that
${\mathcal{V}}_{\eta}\colon:=\{\tau\in\Gamma_{AC}(\eta)\colon\tau([a,b])\subseteq\Omega\}$
is an open 00-neighborhood in $\Gamma_{AC}(\eta)$. Setting
${\mathcal{U}}_{\eta}:=\{\gamma\in AC_{L^{p}}([a,b],N)\colon(t,\gamma(t))\in\Omega^{\prime}\;\mbox{for all $t\in[a,b]$}\},$
the map
$\Psi_{\eta}\colon{\mathcal{V}}_{\theta}\to{\mathcal{U}}_{\theta},\quad\tau\mapsto\Sigma\circ\tau$
is a bijection (see Remark 4.3). We show (see Theorem 4.5):
Theorem 1.2
For each smooth manifold $N$ modeled on a sequentially complete locally convex space and $p\in[1,\infty]$, the set $AC_{L^{p}}([a,b],N)$ of all $AC_{L^{p}}$-maps $\gamma\colon[a,b]\to N$ admits a smooth manifold structure such that for each local addition $\Sigma\colon\Omega\to N$, the sets ${\mathcal{U}}_{\eta}$ are open in $AC_{L^{p}}([a,b],N)$ for all $\eta\in AC_{L^{p}}([a,b],N)$ and $\Psi_{\eta}\colon{\mathcal{V}}_{\eta}\to{\mathcal{U}}_{\eta}$ is a $C^{\infty}$-diffeomorphism.
Using the smooth manifold structures just described, we find:
Theorem 1.3
Let $f\colon N_{1}\to N_{2}$ be a smooth map between smooth manifolds $N_{1}$ and $N_{2}$ modeled on sequentially complete locally convex spaces such that $N_{1}$ and $N_{2}$ admit a local addition and $p\in[1,\infty]$. We then have $f\circ\gamma\in AC_{L^{p}}([a,b],N_{2})$ for all $\gamma\in AC_{L^{p}}([a,b],N_{1})$; the map
$AC_{L^{p}}([a,b],f)\colon AC_{L^{p}}([a,b],N_{1})\to AC_{L^{p}}([a,b],N_{2}),\quad\gamma\mapsto f\circ\gamma$
is smooth.
More generally, $AC_{L^{p}}([a,b],f)$ is $C^{r}$ for $r\in\mathbb{N}\cup\{0,\infty\}$ whenever $f$ is $C^{r+2}$ (see Proposition 4.8).
If, in the situation of Definition 1.1, $N$ is $\mathbb{K}$-analytic manifold modeled on a locally convex topological $\mathbb{K}$-vector space over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and $\theta\colon\Omega\to\Omega^{\prime}$ is a diffeomorphism of $\mathbb{K}$-analytic manifolds, then $\Sigma$ is called a $\mathbb{K}$-analytic local addition. In this case, $AC_{L^{p}}([a,b],N)$ can be given a $\mathbb{K}$-analytic manifold structure modeled on the locally convex topological $\mathbb{K}$-vector spaces $\Gamma_{\eta}$ with properties as in Theorem 1.2, replacing $C^{\infty}$-diffeomorphisms with diffeomorphisms of $\mathbb{K}$-analytic manifolds there (see Corollary 4.6). If $N_{j}$ is a $\mathbb{K}$-analytic manifold modeled on a sequentially complete locally convex topological $\mathbb{K}$-vector space such that $N_{j}$ admits a $\mathbb{K}$-analytic local addition for $j\in\{1,2\}$, then the map $AC_{L^{p}}([a,b],f)$ described in Theorem 1.3 is $\mathbb{K}$-analytic for all $p\in[1,\infty]$ (see Corollary 4.9).
Manifolds of absolutely continuous paths occur in the regularity theory of infinite-dimensional Lie groups. Consider a Lie group $G$ modeled on a sequentially complete locally convex space, with tangent space ${\mathfrak{g}}:=T_{e}G$ at the neutral element $e\in G$. For $g\in G$, let $\rho_{g}\colon G\to G$, $x\mapsto xg$ be the right multiplication with $g$. Then
$TG\times G\to TG,\quad(v,g)\mapsto T\rho_{g}(v)=:v.g$
is a smooth map and a right action of $G$ in $TG$. The following concept strengthens “regularity” in the sense of Milnor [21].
Definition 1.4
Following [23] (cf. also [10]), $G$ is called $L^{p}$-semiregular if, for each $\gamma\in{\mathcal{L}}^{p}([0,1],\mathfrak{g})$, the initial value problem
$\dot{\eta}(t)=\gamma(t).\eta(t),\quad\eta(0)=e$ (1)
has an $AC_{L^{p}}$-solution $\eta\colon[0,1]\to G$. Then $\eta$ is necessarily unique; we call $\eta$ the evolution of $\gamma$ and write $\operatorname{Evol}([\gamma]):=\eta$. If $G$ is $L^{p}$-semiregular and $\operatorname{Evol}\colon L^{p}([0,1],\mathfrak{g})\to AC_{L^{p}}([0,1],G)$ is smooth, then $G$ is called $L^{p}$-regular.
Remark 1.5
If $G$ is modeled on a Fréchet space, (1) is required to hold for almost all $t\in[0,1]$ with respect to Lebesgue measure. In the general case, $\eta$ is required to be a Carathéodory solution to (1), i.e., it solves the corresponding integral equation piecewise in local charts. We mention that a Lie group $G$ is $L^{p}$-regular if and only if $G$ is $L^{p}$-semiregular and $\operatorname{Evol}$ is smooth as map
$L^{p}([0,1],\mathfrak{g})\to C([0,1],G)$
(see [23]). The latter holds whenever $\operatorname{Evol}\colon L^{p}([0,1],\mathfrak{g})\to C([0,1],G)$ is continuous at 00 (see [11, Theorem E]).
Now considetr a half-Lie group $G$ modeled on a sequentially complete locally convex space $E$. Thus $G$ is a group, endowed with a smooth manifold structure modeled on $E$ which makes $G$ a topological group and turns the right translation $\rho_{g}\colon G\to G$ into a smooth mapping for each $g\in G$ (cf. [5, 18]). Let us use notation as in the case of Lie groups.
Definition 1.6
We say that a half-Lie group $G$ is $L^{p}$-semiregular if the differential equation
$\dot{y}(t)=\gamma(t).y(t)$
satisfies local uniqueness of Carathéodory solutions for each $\gamma\in{\mathcal{L}}^{p}([0,1],\mathfrak{g})$ (in the sense of [13]) and the initial value problem (1) has a Carathéodory solution $\operatorname{Evol}(\gamma):=\eta\colon[0,1]\to G$.
The Lie group $\text{Diff}(M)$ of $C^{\infty}$-diffeomorphisms of a compact smooth manifold $M$ without boundary is known to be $L^{1}$-regular, and also the Lie group $\text{Diff}_{K}(\mathbb{R}^{n})$ of all $C^{\infty}$-diffeomorphisms $\phi\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that $\phi(x)=x$ for all $x\in\mathbb{R}^{n}\setminus K$, for each compact subset $K\subseteq\mathbb{R}^{n}$ (see [10]). For each positive integer $r$, the following analogues are obtained (see Theorems  7.5 and 6.3):
Theorem 1.7
Let $1\leq p<\infty$. For each compact smooth manifold $M$ without boundary and $r\in\mathbb{N}$, the half-Lie group $G:=\mbox{\emph{Diff}}^{r}(M)$ of all $C^{r}$-diffeomorphisms $\phi\colon M\to M$ is $L^{p}$-semiregular. Moreover, its evolution map
$\mbox{\emph{Evol}}\colon L^{p}([0,1],\mathfrak{g})\to AC_{L^{p}}([0,1],G)$
is continuous.
Here $\mathfrak{g}$ is the Banach space of $C^{r}$-vector fields on $M$.
Theorem 1.8
Let $1\leq p<\infty$ and $r\in\mathbb{N}$. For each positive integer $n$ and compact subset $K$ of $\mathbb{R}^{n}$, the half-Lie group $G:=\mbox{\emph{Diff}}^{r}_{K}(\mathbb{R}^{n})$ of all $C^{r}$-diffeomorphisms $\phi\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ with $\phi|_{\mathbb{R}^{n}\setminus K}=\text{id}_{\mathbb{R}^{n}\setminus K}$ is $L^{p}$-semiregular. Moreover, its evolution map
$\mbox{\emph{Evol}}\colon L^{p}([0,1],\mathfrak{g})\to AC_{L^{p}}([0,1],G)$
is continuous.
If we replace $L^{p}$ with $L_{rc}^{\infty}$ (see Definition 2.5) the result remains valid.
Here $\mathfrak{g}$ is the Banach space of all $C^{r}$-vector fields on $\mathbb{R}^{n}$ which vanish outside $K$.
For an $L^{p}$-semiregular half-Lie group $G$ admitting a local addition with $1\leq p<\infty$, the smooth manifold structure on $AC_{L^{p}}([0,1],G)$ provided by Theorem 1.2 makes it possible to discuss continuity properties and differentiability properties of the evolution map as a mapping
$\operatorname{Evol}\colon L^{p}([0,1],\mathfrak{g})\to AC_{L^{p}}([0,1],G).$
So far, we have one positive result in this regard:
Theorem 1.9
Let $1\leq p<\infty$. Let $G$ be a right half-Lie group modeled in a Fréchet space $E$ which admits local addition. Let $G$ be $L^{p}$-semiregular with continuous evolution map
$\operatorname{Evol}_{C}:L^{p}([0,1],T_{e}G),\to C([0,1],G),\quad\gamma\mapsto\operatorname{Evol}_{C}(\gamma).$
If the restriction of the right action
$\tau:T_{e}G\times G\to TG,\quad(v,g)\mapsto v.g$
is continuous, then the evolution map
$\operatorname{Evol}:L^{p}([0,1],T_{e}G),\to AC_{L^{p}}([0,1],G),\quad\gamma\mapsto\operatorname{Evol}(\gamma)$
is continuous.
If we replace $L^{p}$ with $L_{rc}^{\infty}$ the result remains valid.
So far, $C^{0}$-regularity has been investigated for the half-Lie group $\text{Diff}^{r}(M)$ in a suitable sense (see the sketch in [20]). Independently, related questions of regularity have been considered by Pierron and Trouvé (see [26]).
2 Preliminaries
Definition 2.1
Let $E$ and $F$ be real locally convex spaces, $U\subseteq E$ be open and $f:U\to F$ be a map. We say that $f$ is $C^{0}$ if it is continuous. We say that $f$ is $C^{1}$ if $f$ is continuous, the directional derivative
$df(x,y):=(D_{y}f)(x):=\lim_{t\to 0}\frac{1}{t}(f(x+ty)-f(x))$
(with $t\neq 0$) exists in $F$ for all $(x,y)\in U\times E$, and $df:U\times E\to F$ is continuous. Recursively, for $k\in\mathbb{N}$ we say that $f$ is $C^{k}$ if $f$ is $C^{1}$ and $df:U\times E\to F$ is $C^{k-1}$. We say that $f$ is $C^{\infty}$ (or smooth) if $f$ is $C^{k}$ for each $k\in\mathbb{N}$.
Definition 2.2
Let $E_{1},E_{2}$ and $F$ be real locally convex spaces, $U\subseteq E_{1}$ and $V\subseteq E_{2}$ be open subsets, $r,s\in\mathbb{N}\cup\{0,\infty\}$ and $f:U\times V\to F$ be a map. If the iterated directional derivatives
$d^{(i,j)}f((x,a),y_{1},...,y_{i},b_{1},...,b_{j}):=(D_{(y_{1},0)}...D_{(y_{i},0)}D_{(0,b_{1})}...D_{(0,b_{j})}f)(x,a)$
exists for all $i,j\in\mathbb{N}\cup\{0\}$ such that $i\leq r$ and $j\leq s$, and all $y_{1},...,y_{i}\in E_{1}$ and $b_{1},...,b_{j}\in E_{2}$. Moreover, we assume that the mappings
$d^{(i,j)}f:U\times V\times E_{1}^{i}\times E_{2}^{j}\to F$
are continuous. Then the map $f$ is called a $C^{r,s}$-map.
Definition 2.3
Let $X$ be a locally compact topological space, endowed with a measure
$\mu:\mathcal{B}(X)\to[0,\infty]$ on its $\sigma$-algebra of Borel sets and let $Y$ be a topological space. A function $\gamma:X\to Y$ is called Lusin $\mu$-measurable (or $\mu$-measurable) if for each compact subset $K\subseteq X$ there exists a sequence $(K_{n})_{n\in\mathbb{N}}$ of compact subsets $K_{n}\subseteq K$ such that each restriction $\gamma|_{K_{n}}$ is continuous and $\mu\left(K\backslash\cup_{n\in\mathbb{N}}K_{n}\right)=0$.
For details of the construction of Lebesgue spaces, we refer the reader to see [24].
Definition 2.4
Let $E$ be a locally convex space, $a<b$ be real numbers, $1\leq p<\infty$ and $\lambda:\mathcal{B}([a,b])\to[0,\infty)$ be the restriction of the Lebesgue measure on $\mathbb{R}$. We define the set $\mathcal{L}^{p}([a,b],E)$ as the set of all Lusin $\lambda$-measurable functions $\gamma:[a,b]\to E$ such that for each continuous seminorm $q$ on $E$ we have
$q\circ\gamma\in\mathcal{L}^{p}([a,b],\mathbb{R}).$
And we endow it with the locally convex topology defined by the family of seminorms
$\lVert\cdot\lVert_{\mathcal{L}^{p},q}:\mathcal{L}^{p}([a,b],E)\to[0,\infty[,\quad\lVert\gamma\lVert_{\mathcal{L}^{p},q}:=\lVert q\circ\gamma\lVert_{\mathcal{L}^{p}}.$
Let $\gamma\sim\eta$ if and only if $\gamma=\eta$ a.e. We define the Hausdorff locally convex space
$L^{p}([a,b],E):=\mathcal{L}^{p}([a,b],E)/[0]$
with seminorms
$\lVert\cdot\lVert_{L^{p},q}:L^{p}([a,b],E)\to[0,\infty[,\quad\lVert[\gamma]\lVert_{L^{p},q}:=\lVert\gamma\lVert_{\mathcal{L}^{p},q}.$
Definition 2.5
Let $E$ be a locally convex space, $a<b$ be real numbers and $\lambda:\mathcal{B}([a,b])\to[0,\infty)$ be the restriction of the Lebesgue measure on $\mathbb{R}$. We define the set $\mathcal{L}^{\infty}([a,b],E)$ of all Lusin $\lambda$-measurable, essentially bounded functions $\gamma:[a,b]\to E$. For each continuous seminorm $q:E\to\mathbb{R}^{n}$ we define the seminorm
$\lVert\gamma\lVert_{\mathcal{L}^{q}}:=\operatorname*{ess\,sup}_{t\in[a,b]}q\circ\gamma(t).$
We endow $\mathcal{L}^{\infty}([a,b],E)$ with the Hausdorff locally convex topology given by these seminorms.
Let $\gamma\sim\eta$ if and only if $\gamma=\eta$ a.e. We define the Hausdorff locally convex space
$L^{\infty}([a,b],E):=\mathcal{L}^{\infty}([a,b],E)/[0]$
with seminorms
$\lVert\cdot\lVert_{L^{\infty},q}:L^{\infty}([a,b],E)\to[0,\infty[,\quad\lVert[\gamma]\lVert_{L^{\infty},q}:=\lVert\gamma\lVert_{\mathcal{L}^{\infty},q}.$
We define the vector space $\mathcal{L}_{rc}^{\infty}([a,b],E)$ of all Borel measurable functions $\gamma:[a,b]\to E$ with compact and metrizable image, endowed with the topology induced by $\mathcal{L}^{\infty}([a,b],E)$. Thus
$L_{rc}^{\infty}([a,b],E):=\mathcal{L}_{rc}^{\infty}([a,b],E)/[0]$
is a Hausdorff locally convex space.
Remark 2.6
Let $E$ be a Frechet space and $p\in[0,1]$. If $L_{\mathcal{B}}^{p}([0,1],E)$ denotes the Lebesgue space constructed with the set of Borel measurable functions (see [10]), then $L_{\mathcal{B}}^{p}([0,1],E)$ coincides with $L^{p}([0,1],E)$ ([24, Proposition 2.10]).
Remark 2.7
Let $1\leq q\leq p<\infty$, then
$\mathcal{L}^{\infty}([a,b],E)\subseteq\mathcal{L}^{p}([a,b],E)\subseteq\mathcal{L}^{q}([a,b],E)\subseteq\mathcal{L}^{1}([a,b],E)$
as a consequence of Hölder’s inequality.
Definition 2.8
Let $E$ be a locally convex space and ${\gamma:[a,b]\to E}$ be a Lusin $\lambda$-measurable function such that $\alpha\circ\gamma\in\mathcal{L}^{1}([a,b],\mathbb{R})$ for each $\alpha\in E^{\prime}$. An element $z\in E$ is called the weak integral of $\gamma$ if
$\forall\alpha\in E^{\prime}:\quad\alpha(z)=\int_{a}^{b}(\alpha\circ\gamma)(s)ds.$
[…]
The following lemma [24, Proposition 2.26] allow us to define absolute continuous functions with vector values.
Lemma 2.9
[…]
is continuous.
A related important result of weak integrals is the Mean Value Theorem (see e.g. [9]).
Theorem 2.10
Let $E$ and $F$ be locally convex spaces, $U\subseteq E$ be an open subset, $f:U\to F$ a $C^{1}$-map and $x,y\in U$ such that the line segment $\{tx+(1-t)y\in E:t\in[0,1]\}$ is contained in $U$. Then
[…]
Definition 2.11
Let $E$ be a sequentially complete locally convex space and $p\in[1,\infty]$. For $t_{0}\in[a,b]$, we say that a function $\eta:[a,b]\to E$ is $L^{p}$-absolutely continuous (or just absolutely continuous if there is no confusion) if there exists a $[\gamma]\in L^{p}([a,b],E)$ such that
[…]
We denote the space of all $L^{p}$-absolutely continuous functions by $AC_{L^{p}}([a,b],E)$. Let $t_{0}\in[a,b]$ be fixed, since $\eta^{\prime}:=[\gamma]$ is necessarily unique, the map
$\Phi:AC_{L^{p}}([a,b],E)\to E\times L^{p}([a,b],E),\quad\eta\mapsto(\eta(t_{0}),\eta^{\prime})$
is an isomorphism of vector spaces. We endow $AC_{L^{p}}([a,b],E)$ with the Hausdorff locally convex vector topology which makes $\Phi$ an isomorphism of topological vector space, which is independent of the choice of $t_{0}$.
The following result will allow us to study differentiability of functions with values in $AC_{L^{p}}([a,b],E)$ [10, Lemma 3.9].
Theorem 2.12
Let $E$ be a Fréchet space and $p\in[1,\infty]$. Then the map
$\Psi:AC_{L^{p}}([a,b],E)\to C([a,b],E)\times{L^{p}}([a,b],E),\quad\eta\mapsto(\eta,\eta^{\prime})$
is a linear topological embedding with closed image.
Remark 2.13
Let $p\in[1,\infty]$. Since the inclusion map ${AC_{L^{p}}([a,b],E)\to C([a,b],E)}$ is continuous [24, Lemma 3.2], the topology on $AC_{L^{p}}([a,b],E)$ is independent of the choice of $t_{0}$ and finer than the compact-open topology. Hence the sets
$AC_{L^{p}}([a,b],V):=\{\eta\in AC_{L^{p}}([a,b],E):\eta([a,b])\subseteq V\}$
are open on $AC_{L^{p}}([a,b],E)$, for each open subset $V\subseteq E$.
For maps between absolute continuous function spaces, we have the following results (see [24]).
Lemma 2.14
Let $E$ be a sequentially complete locally convex space and $p\in[1,\infty]$. For $c,d\in\mathbb{R}$ with $c<d$ we define a map $g:[c,d]\to[a,b]$ via
$g(t)=a+\frac{t-c}{d-c}(b-a),\quad t\in[c,d].$
Then the map
$AC_{L^{p}}(g,E):AC_{L^{p}}([a,b],E)\to AC_{L^{p}}([c,d],E),\quad\eta\mapsto\eta\circ g$
is continuous linear.
Lemma 2.15
Let $E$ and $F$ be sequentially complete locally convex space, $p\in[1,\infty]$, $V\subseteq E$ be open subset and ${f:V\to F}$ be a $C^{1}$-map. Then $f\circ\eta\in AC_{L^{p}}([a,b],F)$ for each $\eta\in AC_{L^{p}}([a,b],V)$.
Lemma 2.16
Let $E$ and $F$ be sequentially complete locally convex space, $p\in[1,\infty]$ and $k\in\mathbb{N}\cup\{0,\infty\}$. Let $V\subseteq E$ be open subset and ${f:V\to F}$ be a $C^{k+2}$-map, then the map
$f^{*}:=AC_{L^{p}}([a,b],f):AC_{L^{p}}([a,b],V)\to AC_{L^{p}}([a,b],F),\quad\eta\mapsto f\circ\eta$
is $C^{k}$. Moreover, we have
$d(f^{*})(\eta,\eta_{1})=df\circ(\eta,\eta_{1})$
for all $(\eta,\eta_{1})\in AC_{L^{p}}([a,b],V)\times AC_{L^{p}}([a,b],E)$.
Remark 2.17
For sequentially complete locally convex space $E$ and $F$ we have
$AC_{L^{p}}([a,b],E\times F)\cong AC_{L^{p}}([a,b],E)\times AC_{L^{p}}([a,b],F).$
For the complex case of Lemma 2.16, we will use the following result [9, Lemma 2.5].
Lemma 2.18
Let $E$ and $F$ be complex locally convex spaces, and $f:U\to F$ be a mapping defined on an open subset of $E$. Then $f$ is complex analytic if and only if $f$ is smooth and the mapping $df(x,\cdot):E\to F$ is complex linear for each $x\in U$.
Definition 2.19
Let $E$ and $F$ be real locally convex spaces, $U\subseteq E$ be an open subset and $f:U\to F$ be a map. We say that $f$ is real analytic if it extends to a complex analytic map $f:V\to F_{\mathbb{C}}$ on some open neighborhood $V$ of $U$ in $E_{\mathbb{C}}$.
Lemma 2.20
Let $E$ and $F$ be sequentially complete locally convex spaces over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$, $p\in[1,\infty]$, $V\subseteq E$ be an open set and $f:V\to F$ be a $\mathbb{K}$-analytic map. Then the map
$f^{*}:=AC_{L^{p}}([a,b],f):AC_{L^{p}}([a,b],V)\to AC_{L^{p}}([a,b],F),\quad\eta\mapsto f\circ\eta$
is $\mathbb{K}$-analytic.
Proof.
First we consider the case $\mathbb{K}=\mathbb{C}$. By Lemma 2.16 the map $f^{*}$ is smooth and $d(f^{*})$ is complex linear in the second variable, hence by Lemma 2.18 the map $f$ is complex analytic.
If $\mathbb{K}=\mathbb{R}$, then by definition the map $f$ has a complex analytic extension $\tilde{f}$, hence $(\tilde{f})^{*}$ is the complex analytic extension of $f^{*}$, whence $f^{*}$ is real analytic. ∎
Definition 2.21
Let $N$ be a smooth manifold modeled with on a sequentially complete locally convex space $E$ and $p\in[1,\infty]$. We say that a function $\eta:[a,b]\to N$ is $L^{p}$-absolutely continuous if it is continuous and there exists a partition $\{t_{0},...,t_{n}\}$ of $[a,b]$ such that for each $j\in\{1,...,n\}$, exists a chart $\varphi_{j}:U_{j}\to V_{j}$ that verify
  • i)
    $\eta([t_{j-1},t_{j}])\subseteq U_{j}$.
  • ii)
    $\varphi_{j}\circ\eta|_{[t_{j-1},t_{j}]}\in AC_{L^{p}}([t_{j-1},t_{j}],E)$.
In this case, we say that these charts verify the definition of $L^{p}$-absolute continuity for $\eta$. If there is no confusion, we simple call $\eta$ absolutely continuous.
For the case of absolute continuous functions with valued in a manifold, we have the following results (see [24]).
Lemma 2.22
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ and $p\in[1,\infty]$. If $\eta\in AC_{L^{p}}([a,b],N)$, then
$\varphi\circ\eta|_{[\alpha,\beta]}\in AC_{L^{p}}([\alpha,\beta],E),$
for each chart $\varphi:U\to V$ of $N$ and each subinterval $[\alpha,\beta]\subseteq[a,b]$ such that $\eta([\alpha,\beta])\subseteq U$.
Lemma 2.23
Let $M$ and $N$ be smooth manifolds modeled on sequentially complete locally convex spaces and $p\in[1,\infty]$. If $f:M\to N$ is a $C^{1}$-map, then $f\circ\eta\in AC_{L^{p}}([a,b],N)$ for each $\eta\in AC_{L^{p}}([a,b],M)$.
3 The space of absolutely continuous sections
Definition 3.1
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$, $\pi_{TN}:TN\to N$ its vector bundle and $p\in[1,\infty]$. For $\eta\in AC_{L^{p}}([a,b],N)$ we define the set
${\Gamma_{AC}(\eta)}:=\{\sigma\in AC_{L^{p}}([a,b],TN):\pi_{TN}\circ\sigma=\eta\}$
and we endow it with the pointwise operations, making it a vector space.
For a partition $P_{n}=\{t_{0},...,t_{n}\}$ of $[a,b]$ and charts $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ of $N$ that verify the definition of absolute continuity for $\eta$. Since $\eta([t_{i-1},t_{i}])\subseteq U_{i}$, we have
$\sigma([t_{i-1},t_{i}])\subseteq TU_{i},\quad\text{for all }\sigma\in{\Gamma_{AC}(\eta)}$
for each $i\in\{1,...,n\}$. We endow ${\Gamma_{AC}(\eta)}$ with the Hausdorff locally convex vector topology which is generated by the initial topology with respect to the mappings
$h_{i}:{\Gamma_{AC}(\eta)}\to AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\sigma\mapsto h_{i}(\sigma)=d\varphi_{i}\circ\sigma|_{[t_{i-1},t_{i}]}$
with $i\in\{1,...,n\}$.
Proposition 3.2
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. Let $P=\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ be charts of $N$ that verify the definition of absolute continuity for $\eta$, then the map
$\Phi_{\eta,P}:{\Gamma_{AC}(\eta)}\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\sigma\mapsto\left(d\varphi_{i}\circ\sigma|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}$
is a linear topological embedding with closed image given by the set of all elements $(\tau_{i})_{i=1}^{n}$ such that
$\tau_{i}(t_{i})=d\varphi_{i}\circ(T\varphi_{i+1})^{-1}\Big{(}\varphi_{i+1}\circ\eta(t_{i}),\tau_{i+1}(t_{i})\Big{)},\quad\text{ for all }i\in\{1,...,n-1\}.$
Proof.
The linear map $\Phi_{\eta,P}$ is continuous by the previous definition and it is injective.
Let $i\in\{1,...,n\}$. If $W_{i}=\varphi_{i}(U_{i})\cap\varphi_{i+1}(U_{i+1})$, then the map
$g_{i}:W_{i}\times E\to E,\quad(x,y)\mapsto d\varphi_{i}\circ(T\varphi_{i+1})^{-1}(x,y),$
is continuous and linear in the second component. This enables us to define the closed vector subspace $K$ given by the elements
$(\tau_{i})_{i=1}^{n}\in\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E)$
which verify
$\tau_{i}(t_{i})=d\varphi_{i}\circ(T\varphi_{i+1})^{-1}\Big{(}\varphi_{i+1}\circ\eta(t_{i}),\tau_{i+1}(t_{i})\Big{)},\quad\text{ for all }i\in\{1,...,n-1\}$
We will show that the image set of $\Phi_{\eta,P}$, denoted by $\text{Im}(\Phi_{\eta,P})$, coincide with the closed subspace $K$. Indeed, the space $\text{Im}(\Phi_{\eta,P})$ is contained in $K$ by definition of $\Phi_{\eta,P}$. Let us consider now $\tau=(\tau_{i})_{i=1}^{n}\in K$. We define the maps
$\sigma_{i}:[t_{i-1},t_{i}]\to TN,\quad s\mapsto(T\varphi_{i})^{-1}\Big{(}\varphi_{i}\circ\eta(s),\tau_{i}(s)\Big{)}$
and
$\sigma_{\tau}:[a,b]\to TN,\quad t\mapsto\sigma_{i}(t),\text{ for }t\in[t_{i-1},t_{i}].$
By the Glueing Lemma, the map $\sigma_{\tau}$ is continuous. Moreover, by Lemma 2.23 each function $\sigma_{i}$ is absolutely continuous, hence $\sigma$ it is too. Since ${\pi_{TN}\circ\sigma_{\tau}=\eta}$ we have that $\sigma_{\tau}\in{\Gamma_{AC}(\eta)}$ and $\Phi_{\eta,P}(\sigma_{\tau})=(\tau_{i})_{i=1}^{n}$, and $\text{Im}(\Phi_{\eta,P})=K$.
It remains to show that the inverse map
$\Phi_{\eta,P}^{-1}:\text{Im}(\Phi_{\eta,P})\to\Gamma_{AC}(\eta),\quad(\tau_{i})_{i=1}^{n}\mapsto\sigma_{\tau}$
is continuous. If $h_{i}$ are the functions that define the topology (Definition 3.1), then for each $i\in\{1,...,n\}$ we have $h_{i}\circ\Phi_{\eta,P}^{-1}=q_{i}$, where $q_{i}$ is the continuous linear map
$q_{i}:\prod_{j=1}^{n}AC_{L^{p}}([t_{j-1},t_{j}],E)\to AC_{L^{p}}([t_{i-1},t_{i}],E),\quad(\eta_{j})_{j=1}^{n}\mapsto\eta_{i}.$
Hence $\Phi_{\eta,P}^{-1}$ is continuous. ∎
Remark 3.3
From now we consider the map $\Phi_{\eta,P}$ as the homeomorphism
$\Phi_{\eta,P}:{\Gamma_{AC}(\eta)}\to\text{Im}(\Phi_{\eta,P}).$
Corollary 3.4
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. If $E$ is a Banach space (resp. a Fréchet space), then ${\Gamma_{AC}(\eta)}$ is a Banach space (resp. a Fréchet space).
Proof.
This follows from the fact that each vector space $AC_{L^{p}}([t_{i-1},i],E)$ is a Banach space (resp. a Fréchet space). ∎
Proposition 3.5
Let $M$ and $N$ be smooth manifolds modeled on sequentially complete locally convex spaces, $k\in\mathbb{N}\cup\{0,\infty\}$, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],M)$. If $f:M\to N$ is a $C^{k+3}$-map, then $Tf\circ\sigma\in\Gamma_{AC}(f\circ\eta)$ for each $\sigma\in\Gamma_{AC}(\eta)$. Moreover, the map
$\widetilde{f}:\Gamma_{AC}(\eta)\to\Gamma_{AC}(f\circ\eta),\quad\sigma\mapsto Tf\circ\sigma$
is continuous linear.
Proof.
Let $E_{M}$ and $E_{N}$ be the modeled spaces of $M$ and $N$ respectively. By Lemma 2.15 we have $f\circ\eta\in AC_{L^{p}}([a,b],N)$ and $Tf\circ\sigma\in AC_{L^{p}}([a,b],TN)$ for each $\sigma\in\Gamma_{AC}(\eta)$. Since $T_{\eta(t)}f\left(\sigma(t)\right)\in T_{f\circ\eta(t)}N$ for each $t\in[a,b]$, we have
$\pi_{TN}\circ(Tf\circ\sigma)=f\circ\eta.$
Thus $\widetilde{f}(\sigma)\in\Gamma_{AC}(f\circ\eta)$ for each $\sigma\in{\Gamma_{AC}(\eta)}$. The linearity of $\widetilde{f}$ is trivial.
Without loss of generality, we can chose a partition $P=\{t_{0},...,t_{n}\}$ of $[a,b]$ such that there is exists the families of charts $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ and $\{(\phi_{i},V_{i}):i\in\{1,..,n\}\}$ that verify the definition of absolutely continuous for $\eta$ and $f\circ\eta$ respectively, such that
$f(U_{i})\subseteq V_{i},\quad\text{for each }i\in\{1,...,n\}.$
For $\sigma\in{\Gamma_{AC}(\eta)}$ and for each $i\in\{1,...,n\}$ we denote $\sigma_{i}=\sigma|_{[t_{i-1},t_{i}]}$ and we define the maps
$F_{i}:AC_{L^{p}}([t_{i-1},t_{i}],E_{M})\to AC_{L^{p}}([t_{i-1},t_{i}],E_{N}),\quad\tau\mapsto\left(d\phi_{i}\circ Tf\circ d\varphi_{i}^{-1}\right)\circ\tau$
and
$F:\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{M})\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{N}),\quad\tau\mapsto\left(\left(d\phi_{i}\circ Tf\circ d\varphi_{i}^{-1}\right)\circ\tau\right)_{i=1}^{n}$
which are continuous by Lemma 2.16. We will show that
$F\left(\text{Im}\left(\Phi_{\eta,P}\right)\right)\subseteq\text{Im}\left(\Phi_{f\circ\eta,P}\right).$
Let $i\in\{1,..,n-1\}$, if
$\tau_{i}:=F_{i}(d\varphi_{i}\circ\sigma_{i})=d\phi_{i}\circ Tf\circ\sigma_{i}$
then
$d\varphi_{i}\circ(T\varphi_{i+1})^{-1}\Big{(}\varphi_{i+1}\circ\eta(t_{i}),\tau_{i+1}(t_{i})\Big{)}$ $=d\varphi_{i}\circ(T\varphi_{i+1})^{-1}\Big{(}\varphi_{i+1}\circ\eta(t_{i}),d\phi_{i+1}\circ Tf\circ\sigma_{i+1}(t_{i})\Big{)}$
$=d\varphi_{i}\Big{(}\eta(t_{i}),Tf\circ\sigma_{i+1}(t_{i})\Big{)}$
$=d\varphi_{i}\circ Tf\circ\sigma_{i+1}(t_{i})$
$=d\varphi_{i}\circ Tf\circ\sigma_{i}(t_{i})$
$=\tau_{i}(t_{i})$
Hence $F\circ\Phi_{\eta,P}(\sigma)\in\text{Im}(\Phi_{f\circ\eta,P})$ and in consequence
$\widetilde{f}=\Phi_{f\circ\eta,P}^{-1}\circ F\circ\Phi_{\eta,P}.$
Thus $\widetilde{f}$ is continuous. ∎
Remark 3.6
The topology of ${\Gamma_{AC}(\eta)}$ does not depend on the partition or charts chosen. Indeed, since the identity map $\text{id}_{M}:M\to M$ is smooth, by previous proposition the map
$\widetilde{\text{id}_{M}}:{\Gamma_{AC}(\eta)}\to\Gamma_{AC}(\text{id}_{M}\circ\eta),\quad\sigma\mapsto T\text{id}_{M}\circ\sigma$
is smooth regardless of the partition or charts chosen. Moreover, this map coincide with the identity map $\text{id}_{\Gamma}:{\Gamma_{AC}(\eta)}\to{\Gamma_{AC}(\eta)}$, $\sigma\mapsto\sigma$.
Remark 3.7
For $\eta\in C([a,b],N)$ we endow the vector space
$\Gamma_{C}(\eta)=\{\sigma\in C([a,b],TN):\pi_{TN}\circ\sigma=\gamma\}$
with the compact-open topology. Since each inclusion
$AC_{L^{p}}([t_{i-1},t_{i}],E)\to C([t_{i-1},t_{i}],E)$
are continuous [24, Lemma 3.2], the inclusion map $J_{\Gamma}:\Gamma_{AC}(\gamma)\to\Gamma_{C}(\gamma)$ is also continuous.
This implies that set
$\mathcal{V}:=\{\sigma\in{\Gamma_{AC}(\eta)}:\sigma([a,b])\subseteq V\}$
is open in ${\Gamma_{AC}(\eta)}$ for each open subset $V\subseteq TN$.
Proposition 3.8
Let $N_{1}$ and $N_{2}$ be smooth manifolds modeled on sequentially complete locally convex spaces, $p\in[1,\infty]$ and $\operatorname{pr}_{i}:N_{1}\times N_{2}\to N_{i}$ be the $i$-projection for $i\in\{1,2\}$. If $\eta_{1}\in AC_{L^{p}}([a,b],N_{1})$ and $\eta_{2}\in AC_{L^{p}}([a,b],N_{2})$ then the map
$\widetilde{(\operatorname{pr}_{1},\operatorname{pr}_{2})}:\Gamma_{AC}(\eta_{1}\times\eta_{2})\to\Gamma_{AC}(\eta_{1})\times\Gamma_{AC}(\eta_{1}),\quad\sigma\mapsto(T\operatorname{pr}_{1},T\operatorname{pr}_{2})(\sigma)$
is a linear homeomorphism.
Proof.
This is consequence of Proposition 3.5, with $f=(\operatorname{pr}_{1},\operatorname{pr}_{2})$ and considering that the map $Tf=(T\operatorname{pr}_{1},T\operatorname{pr}_{2}):T(N_{1}\times N_{2})\to TN_{1}\times TN_{2}$ is a smooth diffeomorphism. ∎
Proposition 3.9
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. For a partition $P=\{t_{0},...,t_{n}\}$ of $[a,b]$, the map
$\rho:{\Gamma_{AC}(\eta)}\to\prod_{i=1}^{n}\Gamma_{AC}(\eta|_{[t_{i-1},t_{i}]}),\quad\sigma\mapsto(\sigma|_{[t_{i-1},t_{i}]})_{i=1}^{n}$
is a linear topological embedding with closed image.
Proof.
Let $j\in\{1,...,n\}$, If $P_{j}=\{t_{i-1},t_{i}\}$, since $\eta_{j}=\eta|_{[t_{j-1},t_{j}]}\in AC_{L^{p}}([t_{j-1},t_{j}],N)$, we have
$\rho=\Big{(}\Phi_{\eta_{i},P_{i}}^{-1}\Big{)}_{i=1}^{n}\circ\Phi_{\eta,P}.$
The image is given by the closed subspace
$\text{Im}(\rho)=\left\{(\tau_{i})_{i=1}^{n}\in\prod_{i=1}^{n}\Gamma_{AC}(\eta|_{[t_{i-1},t_{i}]}):\tau_{i}(t_{i})=\tau_{i+1}(t_{i})\text{ for all }i\in\{1,...,n-1\}\right\}.$
Thus $(\rho|^{\text{Im}(\rho)})^{-1}:\text{Im}(\rho)\mapsto{\Gamma_{AC}(\eta)}$ is well defined and
$(\rho|^{\text{Im}(\rho)})^{-1}=\Phi_{\eta,P}^{-1}\circ\Big{(}\Phi_{\eta_{i},P_{i}}\Big{)}_{i=1}^{n}.$
Proposition 3.10
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ and $p\in[1,\infty]$. If $g:[c,d]\to[a,b]$ a the map as in Lemma 2.14, then $\eta\circ g\in AC_{L^{p}}([c,d],N)$ for each $\eta\in AC_{L^{p}}([a,b],N)$. Moreover, if $\eta\in AC_{L^{p}}([a,b],N)$, then the map
$L_{g}:{\Gamma_{AC}(\eta)}\to\Gamma_{AC}(\eta\circ g),\quad\sigma\mapsto\sigma\circ g$
is continuous linear.
Proof.
Let $\eta\in AC_{L^{p}}([a,b],N)$, we will show first that $\eta\circ g\in AC_{L^{p}}([c,d],N)$. Let $P=\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ charts of $N$ that verify the definition of absolute continuity for $\eta$. Since $g$ is a strictly increasing function, we can define a partition $Q=\{s_{1},...,s_{n}\}$ of $[c,d]$ such that $g(s_{i})=t_{i}$, for each $i\in\{1,...,n\}$. Moreover, given that
$\varphi_{i}\circ\left(\eta\circ g\right)|_{[s_{i-1},s_{i}]}=\left(\varphi_{i}\circ\eta|_{[t_{i-1},t_{i}]}\right)\circ g|_{[s_{i-1},s_{i}]}$
we have $\eta\circ g\in AC_{L^{p}}([a,b],E)$ by Lemma 2.14. Analogously, we have that $\sigma\circ g\in AC_{L^{p}}([a,b],TN)$ for each $\sigma\in{\Gamma_{AC}(\eta)}$ and
$\pi_{TN}\circ(\sigma\circ g)=\eta\circ g.$
Hence $L_{g}(\sigma)\in\Gamma_{AC}(\eta\circ g)$. To see the continuity of $L_{g}$, for each $i\in\{1,...,n\}$ we define the maps
$G_{i}:AC_{L^{p}}([t_{i+1},t_{i}],E)\to AC_{L^{p}}([s_{i+1},s_{i}],E),\quad\tau\mapsto\tau\circ g|_{[s_{i+1},s_{i}]}$
which are continuous by Lemma 2.14. Considering the topological embeddings $\Phi_{\eta,P}$ and $\Phi_{\eta\circ g,Q}$ (as Proposition 3.2) with the same family of charts, if $(\tau_{i})_{i=1}^{n}\in\text{Im}(\Phi_{\eta,P})$ we have
$\tau_{i}\circ g(s_{i})$ $=\tau_{i}(t_{i})$
$=d\varphi_{i}\circ(T\varphi_{i+1})^{-1}\Big{(}\varphi_{i+1}\circ\eta(t_{i}),\tau_{i+1}(t_{i})\Big{)}$
$=d\varphi_{i}\circ(T\varphi_{i+1})^{-1}\Big{(}\varphi_{i+1}\circ\eta\circ g(s_{i}),\tau_{i+1}\circ g(s_{i})\Big{)}$
for each $i\in\{1,...,n-1\}$. If $G=(G_{1},...,G_{n})$, then $\left(G\circ\Phi_{\eta,P}\right)(\sigma)\in\text{Im}(\Phi_{\eta\circ g,Q})$ and
$L_{g}=\Phi_{\eta\circ g,Q}^{-1}\circ G\circ\Phi_{\eta,Q}.$
Hence $L_{g}$ is continious and clearly linear. ∎
Proposition 3.11
Let $N$ be a smooth manifold modeled on sequentially complete locally convex space $E$, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. Then the evaluation map
$\epsilon:{\Gamma_{AC}(\eta)}\times[a,b]\to TN,\quad(\sigma,t)\mapsto\sigma(t)$
is continuous and linear in the first argument. Moreover, for each $t\in[a,b]$ the point evaluation map
$\epsilon_{t}:{\Gamma_{AC}(\eta)}\to TN,\quad\sigma\mapsto\sigma(t)$
is smooth.
Proof.
The evaluation map
$\tilde{\epsilon}:\Gamma_{C}(\eta)\times[a,b]\to TN,\quad(\sigma,t)\mapsto\sigma(t)$
is continuous and the evaluation map $\tilde{\epsilon}_{t}:\Gamma_{C}(\eta)\to TN$, $\sigma\mapsto\sigma(t)$ is smooth for each $t\in[a,b]$ (see [2]). Then $\epsilon=\tilde{\epsilon}\circ(J_{\Gamma},\text{Id}_{\mathbb{R}})$ and $\epsilon_{t}=\tilde{\epsilon}_{t}\circ J_{\Gamma}$, where $J_{\Gamma}:{\Gamma_{AC}(\eta)}\to\Gamma_{C}(\eta)$ is the inclusion map, which is continuous linear by Remark 3.7. ∎
4 Manifolds of absolutely continuous functions
Definition 4.1
Let $N$ be a smooth manifold and $\pi_{TN}:TN\to N$ its tangent bundle. A local addition is a smooth map ${\Sigma:\Omega\to N}$ defined on a open neightborhood $\Omega\subseteq TN$ of the zero-section ${0_{N}:=\{0_{p}\in T_{p}N:p\in N\}}$ such that
  • a)
    $\Sigma(0_{p})=p$ for all $p\in N$.
  • b)
    The image $\Omega^{\prime}:=\big{(}\pi_{TN},\Sigma\big{)}(\Omega)$ is open in $N\times N$ and the map
    $\theta_{N}:\Omega\to\Omega^{\prime},\quad v\mapsto\big{(}\pi_{TN}(v),\Sigma(v)\big{)}$
    is a $C^{\infty}$-diffeomorphism.
Moreover, if $T_{0_{p}}(\Sigma|_{T_{p}N})=id_{T_{p}N}$ for all $p\in N$, we say that the local addition $\Sigma$ is normalized. We denote the local addition as the pair $(\Omega,\Sigma)$.
If $\theta_{N}:\Omega\to\Omega^{\prime}$ is a diffeomorphism of $\mathbb{K}$-analytic manifolds, we call ${\Sigma:\Omega\to N}$ a $\mathbb{K}$-analytic local addition.
Remark 4.2
If a smooth manifold $N$ admits a local addition, then also its tangent manifold $TN$ admits a local addition [2, Lemma A.11]. Moreover, each manifold which admits local addition also admits a normalized local addition [2, Lemma A.14]. From now we will assume that each local addition is normalized.
Remark 4.3
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. We define the sets
$\mathcal{V}_{\eta}:=\{\sigma\in{\Gamma_{AC}(\eta)}:\sigma([a,b])\subseteq\Omega\}.$
which is open in ${\Gamma_{AC}(\eta)}$ by Remark 3.7 and
$\mathcal{U}_{\eta}:=\{\gamma\in AC_{L^{p}}([a,b],N):(\eta,\gamma)([a,b])\subseteq\Omega^{\prime}\}.$
Lemma 2.23 enable us to define the map
$\Psi_{\eta}:=AC_{L^{p}}([a,b],\Sigma):\mathcal{V}_{\eta}\to\mathcal{U}_{\eta},\quad\sigma\mapsto\Sigma\circ\sigma.$
with inverse given by
$\Psi_{\eta}^{-1}:\mathcal{U}_{\eta}\to\mathcal{V}_{\eta},\quad\gamma\mapsto\theta_{N}^{-1}\circ(\eta,\gamma).$
The following Lemma is just an application of [3, Lemma 10.1] to our particular case.
Lemma 4.4
Let $E$ and $F$ sequentially complete locally convex spaces, $U\subseteq E$ open and $f:U\to F$ a map. If $F_{0}\subseteq F$ is a closed subset and $f(U)\subseteq F_{0}$, then $f:U\to F$ is smooth if and only if $f|^{F_{0}}:U\to F_{0}$ is smooth.
Theorem 4.5
For each smooth manifold $N$ modeled on a sequentially complete locally convex space $E$ which admits a local addition and $p\in[1,\infty]$, the set $AC_{L^{p}}([a,b],N)$ admits a smooth manifold structure such that the sets ${\mathcal{U}}_{\eta}$ are open in $AC_{L^{p}}([a,b],N)$ for all $\eta\in AC_{L^{p}}([a,b],N)$ and $\Psi_{\eta}\colon{\mathcal{V}}_{\eta}\to{\mathcal{U}}_{\eta}$ is a $C^{\infty}$-diffeomorphism.
Proof.
We endow $AC_{L^{p}}([a,b],N)$ with the final topology with respect to the family of maps $\Psi_{\eta}\colon{\mathcal{V}}_{\eta}\to{\mathcal{U}}_{\eta}$, for each $\eta\in AC_{L^{p}}([a,b],N)$. If we define the maps $\Psi_{\eta}^{C}\colon{\mathcal{V}}_{\eta}^{C}\to{\mathcal{U}}_{\eta}^{C}$ on the space of continuous functions $C([a,b],N)$ for each $\eta\in C([a,b],N)$ as Remark 4.3, then the final topology on $C([a,b],N)$ coincide with its topology. Hence the inclusion map
$J:AC_{L^{p}}([a,b],N)\to C([a,b],N),\quad\gamma\mapsto\gamma$
is continuous. Moreover, for each $\eta\in AC_{L^{p}}([a,b],N)$ the set
$\mathcal{U}_{J(\eta)}^{C}:=\{\gamma\in C([a,b],N):(J(\eta),\gamma)([a,b])\subseteq\Omega^{\prime}\}$
is open in $C([a,b],N)$, thus
$\mathcal{U}_{\eta}=\mathcal{U}_{\eta}^{C}\cap AC_{L^{p}}([a,b],N)$
is open in $AC_{L^{p}}([a,b],N)$.
The goal is make the family $\{(\mathcal{U}_{\eta},\Psi_{\eta}^{-1}):\eta\in AC_{L^{p}}([a,b],N)\}$ an atlas for $AC_{L^{p}}([a,b],N)$ for a smooth manifold structure on $AC_{Lp}([a,b],N)$. We need to show that the charts are compatibles, i.e., the smoothness of the map
$\Lambda_{\xi,\eta}:=\Psi_{\xi}^{-1}\circ\Psi_{\eta}:\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})\subseteq{\Gamma_{AC}(\eta)}\to\Gamma_{AC}(\xi),\quad\sigma\mapsto\theta_{N}^{-1}\circ(\xi,\Sigma\circ\sigma).$ (2)
Let $R=\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$, let $\{(U_{\varphi_{i}},\varphi_{i}):i\in\{1,...n\}\}$ and $\{(U_{\phi_{i}},\phi_{i}):i\in\{1,...n\}\}$ be charts that verify the definition of absolutely continuous for $\eta$ and $\xi$ respectively. For $\sigma:[a,b]\to TN$, we denote $\sigma_{i}:=\sigma|_{[t_{i-1},t_{i}]}$. We will study the smoothness of the composition
$\Phi_{\xi,R}\circ\Lambda_{\xi,\eta}:\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})\to\text{Im}(\Phi_{\xi,R})\subset\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\sigma\mapsto\left(d\phi_{i}\circ\Lambda_{\xi,\eta}(\sigma_{i})\right)_{i=1}^{n}$
that by Lemma 4.4, is equivalent to the smoothness of $\Lambda_{\xi,\eta}$. For each $i\in\{1,...,n\}$ we have
$d\tilde{\varphi}_{i}\circ\Lambda_{\xi,\eta}(\sigma_{i})$ $=d\phi_{i}\circ\theta_{N}^{-1}\circ(\xi,\Sigma\circ\sigma_{i})$
$=d\phi_{i}\circ\theta_{N}^{-1}\circ(\phi_{i}^{-1}\left(\phi_{i}\circ\xi\right),\Sigma\circ T\varphi_{i}^{-1}\circ T\varphi_{i}\circ\sigma_{i})$
$=d\phi_{i}\circ\theta_{N}^{-1}\circ(\phi_{i}^{-1}\left(\phi_{i}\circ\xi\right),\Sigma\circ T\varphi_{i}^{-1}\left(\varphi_{i}\circ\eta,d\varphi_{i}\circ\sigma_{i}\right)).$
Lets $H_{i}:\phi_{i}(U_{\phi_{i}})\times\varphi_{i}(U_{i})\times E\to E$ defined by
$H_{i}(x,y,z)=d\phi_{i}\circ\theta_{N}^{-1}\circ(\phi_{i}^{-1}\left(x\right),\Sigma\circ T\varphi_{i}^{-1}(y,z)).$ (3)
Since the function $H_{i}$ is smooth (because is composition of smooth functions) the map
$AC([t_{i-1},t_{i}],H_{i}):AC_{L^{p}}([t_{i-1},t_{i}],E\times E\times E)\to AC_{L^{p}}([t_{i-1},t_{i}],E)\quad\alpha\mapsto H_{i}\circ\alpha$
is smooth by Lemma 2.16. Doing the identification of products of $AC_{L^{p}}$ spaces (Remark 2.17), if we fix the functions $\phi_{i}\circ\xi\in AC_{L^{p}}([t_{i-1},t_{i}],E)$ and $\varphi_{i}\circ\eta\in AC_{L^{p}}([t_{i-1},t_{i}],E)$, we see that the map
$\Theta_{i}:AC_{L^{p}}([t_{i-1},t_{i}],TU_{i})\to AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\tau\mapsto H_{i}\circ(\tilde{\varphi}_{i}\circ\xi,\varphi_{i}\circ\eta,d\varphi_{i}\circ\tau)$
is also smooth. Since the maps $h_{i}:{\Gamma_{AC}(\eta)}\to AC_{L^{p}}([t_{i-1},t_{i}],E),\sigma\mapsto d\phi_{i}\circ\sigma_{i}$ are continuous by definition of the topology, rewriting we have
$\Phi_{\xi,R}\circ\Lambda_{\xi,\eta}(\sigma)=\left(\Theta_{i}\circ h_{i}(\sigma)\right)_{i=1}^{n}$
for each $\sigma\in\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})$, hence $\Lambda_{\xi,\eta}$ is smooth. ∎
Proceding in the same way, using the fact that composition of $\mathbb{K}$-analytics maps is $\mathbb{K}$-analytic and using the analytic version of Lemma 4.4 (see [9]), we can obtain the analogous case.
Corollary 4.6
For each $\mathbb{K}$-analytic manifold $N$ modeled on a sequentially complete locally convex space $E$ which admits a $\mathbb{K}$-analytic local addition and $p\in[1,\infty]$, the set $AC_{L^{p}}([a,b],N)$ admits a $\mathbb{K}$-analytic manifold structure such that the sets ${\mathcal{U}}_{\eta}$ are open in $AC_{L^{p}}([a,b],N)$ for all $\eta\in AC_{L^{p}}([a,b],N)$ and $\Psi_{\eta}\colon{\mathcal{V}}_{\eta}\to{\mathcal{U}}_{\eta}$ is a $\mathbb{K}$-analytic diffeomorphism.
Proposition 4.7
Let $N$ be smooth manifolds modeled on sequentially complete locally convex spaces which admits a local addition and $1\leq q\leq p\leq\infty$. Then
$AC_{L^{\infty}}([a,b],N)\subseteq AC_{L^{p}}([a,b],N)\subseteq AC_{L^{q}}([a,b],N)\subseteq AC_{L^{1}}([a,b],N)$
with smooth inclusion maps.
Proof.
Let $\eta\in AC_{L^{p}}([a,b],N)$. Let $\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ be charts of $N$ such that verify the definition of absolutely continuous for $\eta$. Then, by Proposition 1.2, the smoothness of each inclusion in $\eta$ is equivalent to the smoothness of each inclusion map
$\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],N)\to\prod_{i=1}^{n}AC_{L^{q}}([t_{i-1},t_{i}],N),\quad(\eta)_{i=1}^{n}\mapsto(\eta)_{i=1}^{n}.$
Since each inclusion $AC_{L^{p}}([t_{i-1},t_{i}],N)\to AC_{L^{q}}([t_{i-1},t_{i}],N)$ is continuous linear [24, Remark 3.2], the inclusions are smooth. ∎
Proposition 4.8
Let $M$ and $N$ be smooth manifolds modeled on sequentially complete locally convex spaces $E$ which admits a local addition, $p\in[1,\infty]$ and $k\in\mathbb{N}\cup\{0,\infty\}$. If $f:M\to N$ is a $C^{k+2}$-map, then the map
$AC_{L^{p}}([a,b],f):AC_{L^{p}}([a,b],M)\to AC_{L^{p}}([a,b],N),\quad\eta\mapsto f\circ\eta$
is $C^{k}$.
Proof.
Let $(\Omega_{M},\Sigma_{M})$ and $(\Omega_{N},\Sigma_{N})$ be local additions on $M$ and $N$ respectively. The map make sense by Lemma 2.23. Let $\eta\in AC_{L^{p}}([a,b],N)$, $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ and $(\mathcal{U}_{f\circ\eta},\Psi_{f\circ\eta}^{-1})$ be charts in $\eta\in AC_{L^{p}}([a,b],N)$ and $f\circ\eta\in AC_{L^{p}}([a,b],N)$ respectively, if $\theta_{N}=(\pi_{TN},\Sigma_{N})$ then we define
$F(\sigma):=\Psi_{f\circ\eta}^{-1}\circ AC_{L^{p}}([a,b],f)\circ\Psi_{\eta}=\theta_{N}^{-1}\circ\Big{(}(f\circ\eta),(f\circ\Sigma_{M})\circ\sigma\Big{)}$
for all $\sigma\in\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}([a,b],f)^{-1}(\mathcal{U}_{f\circ\eta})\right)$.
Proceeding as the proof of Theorem 3.1, choosing the corresponding partition $P=\{t_{0},...,t_{n}\}$ of $[a,b]$ and the families of charts $\{(U_{\varphi_{i}},\varphi_{i}):i\in\{1,...,n\}\}$ and $\{(U_{\phi_{i}},{\phi}_{i}):i\in\{1,...,n\}\}$ that verify the definition of absolute continuity for $\eta$ and $f\circ\eta$ respectively, we denote $\sigma_{i}=\sigma|_{[t_{i-1},t_{i}]}$. We will study the continuity of the map
$\Phi_{f\circ\eta}\circ F:\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}([a,b],f)^{-1}(\mathcal{U}_{f\circ\eta})\right)\to\text{Im}(\Phi_{f\circ\eta}),\quad\sigma\mapsto\left(d\phi_{i}\circ F(\sigma_{i})\right)_{i=1}^{n}$
where $\Phi_{f\circ\eta}$ is the topological embedding as Proposition 3.2. For each $i\in\{1,...,n\}$ and $\sigma\in\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}([a,b],f)^{-1}(\mathcal{U}_{f\circ\eta})\right)$ we have
$d\phi_{i}\circ F(\sigma)$ $=d\phi_{i}\circ\theta_{N}^{-1}\circ\Big{(}f\circ\eta,f\circ\Sigma_{M}\circ\sigma\Big{)}$
$=d\phi_{i}\circ\theta_{N}^{-1}\circ\Big{(}\phi_{i}^{-1}\circ\phi_{i}\circ f\circ\eta,f\circ\Sigma_{M}\circ T\varphi_{i}^{-1}\left(\varphi_{i}\circ\eta,d\varphi_{i}\circ\sigma_{i}\right)).$
We define the map $H_{i}:\phi_{i}(U_{\phi_{i}})\times\varphi_{i}(U_{\varphi_{i}})\times E_{M}\to E_{N}$ given by
$H_{i}(x,y,z)=d\phi_{i}\circ\theta_{N}^{-1}\circ\Big{(}\phi_{i}^{-1}(x),f\circ\Sigma_{M}\circ T\varphi_{i}^{-1}(y,z)\Big{)}$
which is a $C^{k+2}$-map. Since $\phi_{i}\circ f\circ\eta\in AC_{Lp}([t_{i-1},t_{i}],E_{N})$ and $\varphi_{i}\circ\eta\in AC_{Lp}([t_{i-1},t_{i}],E_{M})$, then the map
$\Theta_{i}:AC_{L^{p}}([t_{i-1},t_{i}],E_{M})\to AC_{L^{p}}([t_{i-1},t_{i}],E_{N})\quad\tau\mapsto H_{i}\circ(\phi_{i}\circ f\circ\eta,\varphi_{i}\circ\eta,\tau)$
is $C^{k}$. Since the maps $h_{i}:{\Gamma_{AC}(\eta)}\to AC_{L^{p}}([t_{i-1},t_{i}],E_{M})$, $d\sigma\mapsto\varphi_{i}\circ\sigma_{[t_{i-1},t_{i}]}$ are continuous linear by definition, rewriting we have
$\Phi_{f\circ\eta}\circ F(\sigma)=\left(\Theta_{i}\circ h_{i}(\sigma)\right)_{i=1}^{n}$
for each $\sigma\in\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})$, hence the map $AC_{L^{p}}([a,b],f)$ is $C^{k}$. ∎
Proceeding in the same way we have the analogous case.
Corollary 4.9
If $f:M\to N$ is a $\mathbb{K}$-analytic map between $\mathbb{K}$-analytic manifolds modeled on sequentially complete locally convex spaces which admits a $\mathbb{K}$-analytic local addition and $p\in[1,\infty]$. Then the map
$AC_{L^{p}}([a,b],f):AC_{L^{p}}([a,b],M)\to AC_{L^{p}}([a,b],N),\quad\eta\mapsto f\circ\eta$
is $\mathbb{K}$-analytic.
Remark 4.10
The manifold structures for $AC_{L^{p}}([a,b],N)$ given by different local additions are diffeomorphic. Indeed, since the identity map $\text{id}_{N}:N\to N$ is smooth regardless of the chosen local addition, the map
$AC([a,b],\text{id}_{M}):AC_{L^{p}}([a,b],N)\to AC_{L^{p}}([a,b],N),\quad\eta\to\text{id}_{M}\circ\eta$
is, again, smooth regardless of the chosen local addition in each space.
Remark 4.11
The inclusion map $J:AC_{L^{p}}([a,b],N)\to C([a,b],N)$ is smooth. Indeed, let $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ and $(\mathcal{U}_{J(\eta)}^{C},\Psi_{J(\eta)}^{-1})$ be charts in $\eta\in AC_{L^{p}}([a,b],N)$ and $J(\eta)\in C([a,b],N)$ respectively, then
$\Psi_{J(\eta)}^{-1}\circ J\circ\Psi_{\eta}^{-1}(\sigma):\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap J^{-1}(\mathcal{U}_{j(\eta)})\right)\subseteq{\Gamma_{AC}(\eta)}\to\Gamma_{C}(\eta)$
is a restriction of the inclusion map ${\Gamma_{AC}(\eta)}\to\Gamma_{C}(\eta)$.
Moreover, if $U\subseteq N$ be a open subset, then the manifold structure induced by $AC_{L^{p}}([a,b],N)$ on the open subset
$AC_{L^{p}}([a,b],U):=\{\eta\in AC_{L^{p}}([a,b],N):\eta([a,b])\subseteq U\}.$
coincides with the manifold structure on $AC_{L^{p}}([a,b],U)$ constructed with the local addition
$\Sigma|_{\Omega\cap TU}:\Omega\cap TU\to U.$
Proposition 4.12
Let $N_{1}$ and $N_{2}$ be smooth manifolds with local addition modeled on a sequentially complete locally convex space which admits a local addition, $p\in[1,\infty]$ and let $\operatorname{pr}_{i}:N_{1}\times N_{2}\to N_{i}$ be the i-th projection for $i\in\{1,2\}$. Then the map
$\mathcal{P}:AC_{L^{p}}([a,b],N_{1}\times N_{2})\to AC_{L^{p}}([a,b],N_{1})\times AC_{L^{p}}([a,b],N_{2}),\quad\eta\mapsto(\operatorname{pr}_{1},\operatorname{pr}_{2})\circ\eta$
is a diffeomorphism.
Proof.
By previous remark, if $(\Omega_{1},\Sigma_{1})$ and $(\Omega_{1},\Sigma_{1})$ are the local addition on $N_{1}$ and $N_{2}$ respectively, then we can assume that the local additions on $N_{1}\times N_{2}$ is
$\Sigma:=\Sigma_{1}\times\Sigma_{2}:\Omega_{1}\times\Omega_{2}\to N_{1}\times N_{2}$
where $\Omega_{1}\times\Omega_{2}\subseteq TN_{1}\times TN_{2}\cong T(N_{1}\times N_{2})$. The map $\mathcal{P}$ is smooth as consequence of the smoothness of the maps
$AC_{L^{p}}([a,b],\operatorname{pr}_{i}):AC_{L^{p}}([a,b],N_{1}\times N_{2})\to AC_{L^{p}}([a,b],N_{i}),$
for each ${i\in\{1,2\}}$.
Let $(\mathcal{U}_{\eta_{1}}\times\mathcal{U}_{\eta_{2}},\Psi_{\eta_{1}}^{-1}\times\Psi_{\eta_{2}}^{-1})$ and $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ be charts in $(\eta_{1},\eta_{2})\in AC_{L^{p}}([a,b],N_{1})\times AC_{L^{p}}([a,b],N_{2})$ and $\mathcal{P}^{-1}(\eta_{1},\eta_{2})=\eta\in AC_{L^{p}}([a,b],N_{1}\times N_{2})$ respectively. Since the map
$\mathcal{Q}:\Gamma_{AC}(\eta)\to\Gamma_{AC}(\eta_{1})\times\Gamma_{AC}(\eta_{2}),\quad\tau\mapsto(\text{q}_{1},\text{q}_{2})\circ\tau$
is an diffeomorphism of vector spaces, where $\text{q}_{1}$ and $\text{q}_{2}$ are the corresponding projection of the space, we have
$\Psi_{\eta}^{-1}\circ\mathcal{P}^{-1}\circ(\Psi_{\eta_{1}}\times\Psi_{\eta_{2}})(\sigma_{1},\sigma_{2})$ $=(\pi_{N_{1}\times N_{2}},\Sigma)^{-1}\circ\left(\eta,\mathcal{P}^{-1}\circ(\Sigma_{1}\times\Sigma_{2})(\sigma_{1},\sigma_{2})\right)$
$=(\pi_{N_{1}\times N_{2}},\Sigma)^{-1}\circ\left(\eta,\Sigma\circ\mathcal{Q}^{-1}(\sigma_{1},\sigma_{2})\right)$
$=\mathcal{Q}^{-1}(\sigma_{1},\sigma_{2})$
for all $(\sigma_{1},\sigma_{2})\in(\Psi_{\eta_{1}}^{-1}\times\Psi_{\eta_{2}}^{-1})\left(\mathcal{U}_{\eta_{1}}\times\mathcal{U}_{\eta_{2}}\cap\mathcal{P}(\mathcal{U}_{\eta})\right)$. Hence $\mathcal{P}^{-1}$ is smooth. ∎
Proposition 4.13
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits a local addition and $p\in[1,\infty]$. For a partition $P=\{t_{0},...,t_{n}\}$ of $[a,b]$ the map
$T:AC_{L^{p}}([a,b],N)\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],N),\quad\eta\mapsto\left(\eta|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}$
is smooth and a smooth diffeomorphism onto a submanifold of $\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],N)$.
Proof.
It is clear that the map $T$ is well defined and injective. Let denote $\text{Im}(T)$ the image set of the map $T$, then
$\text{Im}(T)=\{(\gamma_{i})_{i=1}^{n}\in\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],N):\gamma_{i}(t_{i})=\gamma_{i+1}(t_{i})\text{ for all }i\in\{1,...,n-1\}\}.$
Let $\gamma:=(\gamma_{i})_{i=1}^{n}\in\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],N)$. For each $i\in\{1,...,n\}$ let $\Psi_{\gamma_{i}}^{-1}:\mathcal{U}_{i}\to\mathcal{V}_{i}$ be charts of $\gamma_{i}$, then the map
$\Psi_{\gamma}^{-1}:=\prod_{i=1}^{n}\Psi_{\gamma_{i}}^{-1}:\prod_{i=1}^{n}\mathcal{U}_{i}\to\prod_{i=1}^{n}\mathcal{V}_{i}$
is a chart for $\gamma$. Let $\eta\in AC_{L^{p}}([a,b],N)$ and $\tilde{\eta}=T(\eta)$, then $\Psi_{\tilde{\eta}}\circ T\circ\Psi_{\eta}^{-1}$ is just restriction of the product of restriction of the smooth maps
${\Gamma_{AC}(\eta)}\to\prod_{i=1}^{n}\Gamma_{AC}(\eta_{i}),\quad\sigma\mapsto\left(\sigma|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}$
thus $T$ is a smooth. Let $\tilde{T}:AC_{L^{p}}([a,b],N)\to\text{Im}(T)$, $\eta\mapsto T(\eta)$, then analogously to Proposition 3.9, we can see that $\tilde{T}^{-1}$ is smooth. Hence $T$ it is a diffeomorphism onto its image. Now we will show that the image $\text{Im}(T)$ is a submanifold. Let $\gamma=(\gamma_{i})_{i=1}^{n}\in\text{Im}(T)$ with charts as before, then for each $i\in\{1,...,n-1\}$ and $\xi=\left(\xi_{i}\right)_{j=1}^{n}\in\text{Im}(T)\cap\prod_{j=1}^{n}\mathcal{U}_{j}$ we have
$\Psi_{\gamma_{i}}^{-1}(\xi_{i})(t_{i})$ $=\theta_{N}^{-1}\circ(\gamma_{i},\xi_{i})(t_{i})$
$=\theta_{N}^{-1}\circ(\gamma_{i}(t_{i}),\xi_{i}(t_{i}))$
$=\theta_{N}^{-1}\circ(\gamma_{i+1}(t_{i}),\xi_{i+1}(t_{i}))$
$=\Psi_{\gamma_{i+1}}^{-1}(\xi_{i+1})(t_{i}).$
This implies that if $K$ denotes the vector space
$K:=\{\left(\sigma_{i}\right)_{i=1}^{n}\in\prod_{i=1}^{n}\Gamma_{AC}(\gamma_{i}):\sigma_{i}(t_{i})=\sigma_{i+1}(t_{i})\text{ for all }i\in\{1,...,n-1\}\}.$
Then $\Psi_{\gamma}^{-1}|_{\mathcal{U}_{\gamma}\cap\text{Im}(T)}:\text{Im}(\Psi)\cap\prod_{i=1}^{n}\mathcal{U}_{i}\to K\cap\prod_{i=1}^{n}\mathcal{V}_{i}$ is a chart of $\text{Im}(T)$, making $\text{Im}(T)$ a smooth submanifold.
Remark 4.14
Let $N$ is a $\mathbb{K}$-analytic manifold modeled on a sequentially complete locally convex space which admits a $\mathbb{K}$-analytic local addition and $p\in[1,\infty]$. Since every continuous linear operator is analytic, the isomorphism
${\Gamma_{AC}(\eta)}\to K,\quad\sigma\mapsto\left(\sigma|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}$
is $\mathbb{K}$-analytic, which implies that $T$ in Proposition 4.13 is a $\mathbb{K}$-analytic diffeomorphism onto the submanifold $\text{Im}(T)$.
Proposition 4.15
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits a local addition, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. If
$g:[c,d]\to[a,b]$ is a map as in Lemma 2.14, then the map
$AC_{L^{p}}(g,N):AC_{L^{p}}([a,b],N)\to AC_{L^{p}}([c,d],N),\quad\eta\mapsto\eta\circ g$
is smooth.
Proof.
By Proposition 3.10 we know that the map is well defined. Let $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ and $(\mathcal{U}_{\eta\circ g},\Psi_{\eta\circ g}^{-1})$ be charts in $\eta\in AC_{L^{p}}([a,b],N)$ and $\eta\circ g\in AC_{L^{p}}([c,d],N)$ respectively, then we have
$\Psi_{\eta\circ g}^{-1}\circ AC_{L^{p}}(g,M)\circ\Psi_{\eta}(\sigma)=\theta_{N}^{-1}\circ(\eta\circ g,\Sigma\circ(\sigma\circ g))$
for all $\sigma\in\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}(g,N)^{-1}(\mathcal{U}_{\eta\circ g})\right)$.
Let $\alpha=\eta\circ g:[c,d]\to N$ and $\tau=\sigma\circ g:[c,d]\to TN$, then both are absolutely continuous, with $\pi_{TN}\circ\tau=\alpha$, then $\tau\in\Gamma_{AC}(\alpha)$ and
$\Psi_{\eta\circ g}^{-1}\circ AC_{L^{p}}(g,M)\circ\Psi_{\eta}(\sigma)$ $=\theta_{N}^{-1}\circ(\alpha,\Sigma\circ\tau)$
$=\Psi_{\alpha}^{-1}\circ\Psi_{\alpha}(\tau)$
$=\tau$
$=\sigma\circ g.$
Hence, $\Psi_{\eta\circ g}^{-1}\circ AC_{L^{p}}(g,M)\circ\Psi_{\eta}$ is a restriction of the map
${\Gamma_{AC}(\eta)}\to{\Gamma_{AC}(\eta\circ g)},\quad\sigma\to\sigma\circ g$
which is continuous linear by Proposition 3.10. ∎
Proposition 4.16
Let $M$, $N$ and $L$ be smooth manifolds modeled on sequentially complete locally convex spaces which admit a local addition and $p\in[1,\infty]$. If $f:L\times M\to N$ is a $C^{k+2}$-map and $\gamma\in AC_{L^{p}}([a,b],L)$ is fixed, then
$f_{*}:AC_{L^{p}}([a,b],M)\to AC_{L^{p}}([a,b],N),\quad\eta\mapsto f\circ(\gamma,\eta)$
is a $C^{k}$-map.
Proof.
Define the smooth map
$C_{\gamma}:AC_{L^{p}}([a,b],N)\to AC_{L^{p}}([a,b],L)\times AC_{L^{p}}([a,b],N),\quad\eta\mapsto(\gamma,\eta)$
Identifying $AC_{L^{p}}([a,b],L)\times AC_{L^{p}}([a,b],N)$ with $AC_{L^{p}}([a,b],L\times N)$, we have
$f_{*}=AC_{L^{p}}([a,b],f)\circ C_{\gamma}.$
Hence $f_{*}$ is $C^{k}$. ∎
Proposition 4.17
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits a local addition, $p\in[1,\infty]$ and $t\in[a,b]$. Then the evaluation map
$\varepsilon:AC_{L^{p}}([a,b],N)\times[a,b]\to N,\quad(\eta,t)\mapsto\eta(t)$
is continuous and the point evaluation map
$\varepsilon_{t}:AC_{L^{p}}([a,b],N)\to N,\quad\eta\mapsto\eta(t)$
is smooth.
Proof.
The evaluation map
$\varepsilon_{c}:C([a,b],N)\times[a,b]\to N,\quad(\eta,t)\mapsto\eta(t)$
is continuous and the point evaluation $(\varepsilon_{c})_{t}:C([a,b],N)\to N$, $\eta\mapsto\eta(t)$ is smooth for each $t\in[a,b]$ (see [1]). Since the inclusion map $J:AC_{L^{p}}([a,b],N)\to C([a,b],N)$ is smooth, we have $\varepsilon=\varepsilon_{c}\circ(J,\text{Id}_{\mathbb{R}})$ and $\varepsilon_{t}=(\varepsilon_{c})_{t}\circ J$ for each $t\in[a,b]$. ∎
Proposition 4.18
Let $M$, $N$ and $L$ smooth manifolds modeled on a sequentially complete locally convex space such that $L$ admits a local addition and $p\in[1,\infty]$. If $f:L\times M\to N$ is a $C^{2}$-map and $\gamma\in AC_{L^{p}}([a,b],L)$ is fixed, then the map
$F:[a,b]\times M\to N,\quad F(t,m)=f(\gamma(t),m)$
is continuous.
Proof.
For each $(t,m)\in[a,b]\times M$ we have
$F(t,m)=AC_{L^{p}}([a,b],f)\circ(\varepsilon(\gamma,\cdot)\times\text{id}_{M})(t,m)$
where $\varepsilon$ is the evaluation map as before. Therefore, the map $F$ is composition of continuous maps. ∎
Proposition 4.19
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits a local addition and $p\in[1,\infty]$. For each $q\in N$ the function $\zeta_{q}:[a,b]\to N$, $t\mapsto q$ is absolutely continuous and the map
$\zeta:N\to AC_{L^{p}}([a,b],N),\quad q\mapsto\zeta_{q}$
is smooth and a topological embedding.
Proof.
Let consider the local addition $\Sigma:\Omega\to N$ and $\theta_{N}:\Omega\to\Omega^{\prime}$ as in Definition 4.1. Let $(U,\varphi)$ be a chart in $q\in N$ such that $\{q\}\times U\subseteq\Omega^{\prime}$ and $(\mathcal{U}_{\zeta_{q}},\Psi_{\zeta_{q}}^{-1})$ be a chart in $\zeta_{q}\in AC_{L^{p}}([a,b],N)$.
If $x\in\varphi\left(U\cap\zeta^{-1}(\mathcal{U}_{\zeta_{q}})\right)$, then for each $t\in[a,b]$ we have
$\Psi_{\zeta_{q}}^{-1}\circ\zeta\circ\varphi^{-1}(x)(t)$ $=\theta_{N}^{-1}\left(\zeta_{q}(t),\zeta_{\varphi^{-1}(x)}(t)\right)$
$=\theta_{N}^{-1}\left(q,\varphi^{-1}(x)\right)$
$=\theta_{N}^{-1}\circ(q,\varphi^{-1}\circ\tilde{\zeta}_{x}(t))$
where $\tilde{\zeta}_{x}:[a,b]\to E$, $t\mapsto x$ is the constant function. Since the map
$\tilde{\zeta}:E\to AC_{L^{p}}([a,b],E),\quad x\mapsto\tilde{\zeta}_{x}$
is a continuous linear, setting the smooth map
$h:\varphi(U)\to TN,\quad z\to\theta_{N}^{-1}\circ(q,\varphi^{-1}(z))$
we have
$\Psi_{\zeta_{p}}^{-1}\circ\zeta\circ\varphi^{-1}=AC_{L^{p}}([a,b],h)\circ\tilde{\zeta}|_{\phi(U)}.$
Hence $\zeta$ is smooth. Moreover, if $t\in[a,b]$, then $\varepsilon_{t}\circ\zeta=\text{id}_{N}:N\to N$. ∎
Remark 4.20
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits a local addition, $p\in[1,\infty]$ and let $TAC_{L^{p}}([a,b],N)$ be the tangent bundle of $AC_{L^{p}}([a,b],N)$. Since the point evaluation map $\varepsilon_{t}:AC_{L^{p}}([a,b],N)\to N$ is smooth for each $t\in[a,b]$, we have
$T\varepsilon_{t}:TAC_{L^{p}}([a,b],N)\to TN.$
For each $v\in TAC_{L^{p}}([a,b],N)$ we define the function
$\Theta_{N}(v):[a,b]\to TN,\quad\Theta_{N}(v)(t)=T\varepsilon_{t}(v).$
Proposition 4.21
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits local addition, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. Then $\Theta_{N}(v)\in{\Gamma_{AC}(\eta)}$ for each $v\in T_{\eta}AC_{L^{p}}([a,b],N)$ and the map
$\Theta_{\eta}:T_{\eta}AC_{L^{p}}([a,b],N)\to{\Gamma_{AC}(\eta)},\quad v\mapsto\Theta_{\eta}(v):=\Theta_{N}|_{T_{\eta}AC_{L^{p}}([a,b],N)}(v)$
is an isomorphism of topological vector spaces.
Proof.
Let $\Sigma:\Omega\to N$ be a normalized local addition of $N$. Since ${\Gamma_{AC}(\eta)}$ is a vector space, we identify its tangent bundle with ${\Gamma_{AC}(\eta)}\times{\Gamma_{AC}(\eta)}$. Let $\Psi_{\eta}:\mathcal{V}_{\eta}\to\mathcal{U}_{\eta}$ be a chart on $\eta$ such that $\Psi_{\eta}(0)=\eta$, then
$T\Psi_{\eta}:T\mathcal{V}_{\eta}\simeq\mathcal{V}_{\eta}\times{\Gamma_{AC}(\eta)}\to TAC_{L^{p}}([a,b],N)$
is a diffeomorphism onto its image. Moreover,
$T\Psi_{\eta}:\{0\}\times{\Gamma_{AC}(\eta)}\to T_{\eta}AC_{L^{p}}([a,b],N)$
is an isomorphism of topological vector spaces. We will show that
$\Theta_{\eta}\circ T\Psi_{\eta}(0,\sigma)=\sigma$
for each $\sigma\in{\Gamma_{AC}(\eta)}$. Which is equivalent to show that
$T\varepsilon_{t}\circ T\Psi_{\eta}(0,\sigma)=\sigma(t)\quad\text{ for all }t\in[a,b].$
Working with the geometric point of view of tangent vectors, we see that $(0,\sigma)$ is equivalent to the curve $[s\mapsto s\sigma]$. Hence, for each $t\in[a,b]$ we have
$T\varepsilon_{t}\circ T\Psi_{\eta}(0,\sigma)$ $=T\varepsilon_{t}\circ T\Psi_{\eta}([s\mapsto s\sigma])$
$=T\varepsilon_{t}([s\mapsto\Psi_{\eta}(s\sigma)])$
$=T\varepsilon_{t}([s\mapsto\Sigma(s\sigma)])$
$=[s\mapsto\Sigma|_{T_{\eta(t)}N}(s\sigma(t))]$
$=T_{0}\Sigma|_{T_{\eta(t)}N}([s\mapsto s\sigma(t)]).$
Since $\Sigma$ is normalized we have $T_{0}\Sigma|_{T_{\eta(t)}N}=\text{id}_{T_{\eta(t)}N}$ and
$T\varepsilon_{t}\circ T\Psi_{\eta}(0,\sigma)=\sigma(t).$
In consequence, for each $\sigma\in{\Gamma_{AC}(\eta)}$, there exists a $v\in T_{\eta}AC_{L^{p}}([a,b],N)$ with $v=T\Psi_{\eta}(0,\sigma)$ such that
$\Theta_{\eta}(v)=\sigma.$
Moreover, the function
$\Theta_{\eta}(v):[a,b]\to TN,\quad t\mapsto\Theta_{N}(v)(t)=\sigma(t)\in T_{\eta(t)}N$
is absolutely continuous with $\pi_{TN}\circ\Theta_{\eta}(v)=\eta$, making the map $\Theta_{\eta}$ an isomorphism of topological vector spaces. ∎
Remark 4.22
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits local addition and $p\in[1,\infty]$. Since $TN$ admits local addition and the vector bundle $\pi_{TN}:TN\to N$ is smooth, the map
$AC([a,b],\pi_{TN}):AC_{L^{p}}([a,b],TN)\to AC_{L^{p}}([a,b],N),\quad\tau\mapsto\pi_{TN}\circ\tau$
is smooth. Moreover, if $\eta\in AC_{L^{p}}([a,b],N)$, then
$AC([a,b],\pi_{TN})^{-1}(\{\eta\})={\Gamma_{AC}(\eta)}.$
The following result follows the same steps as for the case of $C^{\ell}$-maps (with $\ell\geq 0$) from a compact manifold (possibly with rough boundary) to a smooth manifold which admits local addition [2, Theorem A.12].
Proposition 4.23
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits a local addition, $p\in[1,\infty]$ and $\pi_{TN}:TN\to N$ its tangent bundle. Then the map
$AC([a,b],\pi_{TN}):AC_{L^{p}}([a,b],TN)\to AC_{L^{p}}([a,b],N),\quad\tau\mapsto\pi_{TN}\circ\tau$
is a smooth vector bundle with fiber ${\Gamma_{AC}(\eta)}$ over $\eta\in AC_{L^{p}}([a,b],N)$. Moreover, the map
$\Theta_{N}:TAC_{L^{p}}([a,b],N)\to AC_{L^{p}}([a,b],TN),\quad v\mapsto\Theta_{N}(v)$
is an isomorphism of vector bundles.
Proposition 4.24
Let $M$ and $N$ be a smooth manifolds modeled on sequentially complete locally convex spaces which admits a local addition and $p\in[1,\infty]$. If $f:M\to N$ is a $C^{3}$-map, then the tangent map of
$AC_{L^{p}}([a,b],f):AC_{L^{p}}([a,b],M)\to AC_{L^{p}}([a,b],N),\quad\eta\mapsto f\circ\eta$
is given by
$TAC_{L^{p}}([a,b],f)=\Theta_{N}^{-1}\circ AC_{L^{p}}([a,b],Tf)\circ\Theta_{M}.$
Proof.
By Proposition 4.8, since $f$ is $C^{3}$ we know that $AC_{L^{p}}([a,b],f)$ is $C^{1}$, thus $TAC_{L^{p}}([a,b],f)$ exists. Let consider the local addition $\Sigma_{M}:\Omega_{M}\to M$ and $\eta\in AC_{L^{p}}([a,b],M)$.
If $\Psi_{\eta}:\mathcal{V}_{\eta}\to\mathcal{U}_{\eta}$ is a chart on $\eta$ such that $\Psi_{\eta}(0)=\eta$. We consider the isomorphism of vector space
$T\Psi_{\eta}:\{0\}\times{\Gamma_{AC}(\eta)}\to T_{\eta}AC_{L^{p}}([a,b],M).$
For $t\in[a,b]$ we denote the point evaluation in $M$ and $N$ as $\varepsilon_{t}^{M}$ and $\varepsilon_{t}^{N}$ respectively, then for each $\sigma\in{\Gamma_{AC}(\eta)}$ we have
$T\varepsilon_{t}^{N}\circ TAC_{L^{p}}([a,b],f)\circ T\Psi_{\eta}(0,\sigma)$ $=T\varepsilon_{t}\circ TAC_{L^{p}}([a,b],f)\circ T\Psi_{\eta}([s\mapsto s\sigma])$
$=T\varepsilon_{t}^{N}\circ TAC_{L^{p}}([a,b],f)([s\mapsto\Sigma_{M}(s\sigma)])$
$=T\varepsilon_{t}^{N}([s\mapsto f\circ\Sigma_{M}(s\sigma)])$
$=[s\mapsto\varepsilon_{t}\left(f\circ\Sigma_{M}(s\sigma)\right)]$
$=[s\mapsto f\circ\Sigma_{M}(s\sigma(t))]$
$=Tf\circ T_{0}\Sigma_{M}|_{T_{\eta(t)}N}([s\sigma(t)])$
$=Tf([s\mapsto s\sigma(t)])$
$=AC_{L^{p}}([a,b],Tf)(\sigma(t))$
$=AC_{L^{p}}([a,b],Tf)\circ T\varepsilon_{t}^{M}\circ T\Psi_{\eta}(0,\sigma).$
Hence
$\Theta_{N}\circ TAC_{L^{p}}([a,b],f)=AC_{L^{p}}([a,b],Tf)\circ\Theta_{M}.$
Example 4.25
Let $p\in[1,\infty]$. If $G$ is a Lie group modeled in a sequentially locally convex space $E$, then we already know that the space $AC_{L^{p}}([a,b],G)$ is a Lie group with Lie algebra given by $AC_{L^{p}}([a,b],T_{e}G)$ (see [10, 24]). We will give an alternative proof of this.
Let $e\in G$ be the neutral element, let $L_{g}:G\to G$, $h\mapsto gh$ be the left translation by $g\in G$ and the action
$G\times TG\to TG,\quad(g,v_{h})\mapsto g.v_{h}:=TL_{g}(v_{h})\in T_{gh}G.$
If $\varphi:U\subseteq G\to V\subseteq T_{e}G$ is a chart in $e$ such that $\varphi(e)=0$, then the set
$\Omega_{\varphi}:=\bigcup\limits_{g\in G}g.V\subseteq TG$
and the map
$\Sigma_{\varphi}:\Omega_{\varphi}\to G,\quad v\mapsto\pi_{TG}(v)\left(\varphi^{-1}(\pi_{TG}(v)^{-1}.v)\right)$
defines a local addition for $G$ (see e.g. [15]); hence $AC_{L^{p}}([a,b],G)$ is a smooth manifold with charts constructed with $(\Omega_{\varphi},\Sigma_{\varphi})$.
Let $\mu_{G}:G\times G\to G$ and $\lambda_{G}:G\to G$ be the multiplication map and inversion maps on $G$ respectively, we define the multiplication map $\mu_{AC}$ and the inversion map $\lambda_{AC}$ on $AC_{L^{p}}([a,b],G)$ as
$\mu_{AC}:=AC_{L^{p}}([a,b],\mu_{G}):AC_{L^{p}}([a,b],G)\times AC_{L^{p}}([a,b],G)\to AC_{L^{p}}([a,b],G)$
and
$\lambda_{AC}:=AC_{L^{p}}([a,b],\lambda_{G}):AC_{L^{p}}([a,b],G)\to AC_{L^{p}}([a,b],G)$
that by Lemma 2.16 and Proposition 4.8 are smooth.
We observe that for the neutral element $\zeta_{e}:[a,b]\to G$, $t\mapsto e$ of $AC_{L^{p}}([a,b],G)$ we have
$\Gamma_{AC}(\zeta_{e})=AC_{L^{p}}([a,b],T_{e}G).$
If $\Psi_{\zeta_{e}}^{-1}:\mathcal{U}_{\zeta_{e}}\to\mathcal{V}_{\zeta_{e}}$ is a chart in ${\zeta_{e}}\in AC_{L^{p}}([a,b],G)$, then we have $\mathcal{U}_{\zeta_{e}}=AC([a,b],U)$ and ${\mathcal{V}_{\zeta_{e}}=AC([a,b],V)}$. Moreover, we see that
$\Psi_{\zeta_{e}}\circ AC_{L^{p}}([a,b],\varphi)(\eta)$ $=\Sigma_{\varphi}\circ(\varphi\circ\eta)$
$=\pi_{TG}(\varphi\circ\eta)\left(\varphi^{-1}(\pi_{TG}(\varphi\circ\eta)^{-1}.\varphi\circ\eta)\right)$
$=e\varphi^{-1}(e.\varphi\circ\eta)$
$=\eta.$
This enable us to say that for the neutral element ${\zeta_{e}}\in AC_{L^{p}}([a,b],G)$ the chart is given by
$AC_{L^{p}}([a,b],\varphi):AC_{L^{p}}([a,b],U)\to AC([a,b],V),\quad\eta\mapsto\varphi\circ\eta.$
5 Semiregularity of right half-Lie groups.
Definition 5.1
A group $G$, endowed with a smooth manifold structure modeled on a locally convex space, is called a right half-Lie group if it is a topological group and if for all $g\in G$, the right translations $\rho_{g}:G\to G,x\mapsto xg$ are smooth.
Remark 5.2
Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space. We define the right action
$TG\times G\to TG,\quad(v_{h},g)\mapsto v_{h}.g:=T\rho_{g}(v).$
And its restriction
$T_{e}G\times G\to TG,\quad(v,g)\mapsto v.g:=T\rho_{g}(v).$
Unlike on Lie groups, on half-Lie groups this actions may not be smooth. Hence we can not construct a local addition using a convenient chart on the identity (see Example 4.25).
The following proposition is direct application of Proposition 4.8.
Proposition 5.3
Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space which admits local addition and $p\in[1,\infty]$. For $\eta\in AC_{L^{p}}([a,b],G)$ and $g\in G$ we define the function
$\eta.g(t):=\eta(t)g,\quad\text{for all }t\in[a,b].$
Then $\eta.g\in AC_{L^{p}}([a,b],G)$ for each $g\in G$, the map
$AC_{L^{p}}([a,b],\rho_{g}):AC_{L^{p}}([a,b],G)\to AC_{L^{p}}([a,b],G),\quad\eta\mapsto\eta.g$
is smooth.
Remark 5.4
The smoothness of the map
$\mathcal{R}:AC_{L^{p}}([a,b],G)\times G\to AC_{L^{p}}([a,b],G),\quad(\eta,g)\mapsto\eta.g$
would imply the smoothness of the multiplication map on $G$. In fact, since the point evaluation map $\varepsilon_{a}:AC_{L^{p}}([a,b],G)\to G$, $\eta\mapsto\eta(a)$ and the map $\zeta:G\to AC_{L^{p}}([a,b],G)$, $g\mapsto[t\mapsto g]$, are smooth, the multiplication map on $G$ would be smooth as it coincides with the composition
$\varepsilon_{a}\circ\mathcal{R}\circ(\zeta,\text{id}_{G}):G\times G\to G,\quad(h,g)\mapsto hg.$
Definition 5.5
Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space $E$ which admits local addition, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. Let $P=\{t_{0},..,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ charts of $G$ such that verify the definition of absolutely continuous for $\eta$. For each $j\in\{1,...,n\}$ we denote
$\eta_{j}:=\varphi_{j}\circ\eta_{[t_{i-1},t_{i}]}\in AC_{L^{p}}([t_{i-1},t_{i}],E).$
Then $\eta_{j}^{\prime}\in L^{p}([t_{i-1},t_{i}],E)$. Let $\eta^{\prime}_{j}=[\gamma_{j}]$ with $\gamma_{j}\in\mathcal{L}^{p}([t_{i-1},t_{i}],E)$, we define $\gamma:[a,b]\to TG$ via
$\gamma(t):=T(\varphi_{j})^{-1}(\eta_{j}(t),\gamma_{j}(t)),\text{ if }t\in[t_{i-1},t_{i}[\text{ with }j\in\{1,...,n\},$
and $\gamma(b)=T(\varphi_{j})^{-1}(\eta_{j}(b),\gamma_{j}(b))$. Then $\gamma$ is measurable (see [10]) and we define
$\dot{\eta}:=\gamma.$
Definition 5.6
Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space and $p\in[1,\infty]$. We say that $G$ is $L^{p}$-semiregular if for each $[\gamma]\in L^{p}([0,1],T_{e}G)$, there exists an $AC_{L^{p}}$-Carathéodory solution $\eta_{\gamma}\in AC_{L^{p}}([0,1],G)$ of the equation
${\dot{\eta}}_{\gamma}=\gamma.\eta_{\gamma},\quad\eta_{\gamma}(0)=e$
such that the differential equation satisfies local uniqueness of Carathéodory solutions in the sense of [13]. In this case, we define the evolution map
$\operatorname{Evol}:L^{p}([0,1],T_{e}G)\to AC_{L^{p}}([a,b],G),\quad[\gamma]\mapsto\operatorname{Evol}(\gamma):=\eta_{\gamma}.$
Additionally, if $G$ admits a local addition, we say that $G$ is $L^{p}$-regular if $G$ is $L^{p}$-semiregular and if the evolution map is smooth. The definition for the case of $L_{rc}^{\infty}$-semiregularity and $L_{rc}^{\infty}$-regularity is analogous.
Definition 5.7
Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space and $p\in[1,\infty]$. We say that $G$ is locally $L^{p}$-semiregular if there is exists an open 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that for each $[\gamma]\in B$ there exists a $AC_{L^{p}}$-Carathéodory solution $\eta_{\gamma}\in AC_{L^{p}}([0,1],G)$ of the equation
$\dot{\eta}_{\gamma}=\gamma.\eta_{\gamma},\quad\eta_{\gamma}(0)=e,$
and the latter differential equation satisfies local uniqueness of solutions.
For our purpuse, we will use the subdivision property [24, Lemma 2.17].
Lemma 5.8
Let $E$ be a locally convex space, $p\in[1,\infty]$ and $[\gamma]\in L^{p}([0,1],E)$. For each $n\in\mathbb{N}$ and $k\in\{0,1,...,n-1\}$ we define
$\gamma_{n,k}:[0,1]\to E,\quad\gamma_{n,k}(t):=\frac{1}{n}\gamma\left(\frac{k+t}{n}\right).$
Then $[\gamma_{n,k}]\in L^{p}([0,1],E)$. Moreover, for each $[\gamma]\in L^{p}([0,1],E)$ and continuous seminorm $q$ on $L^{p}([0,1],E)$, we have
$\sup_{k\in\{1,...,n-1\}}q(\gamma_{n,k})\to 0,\quad\text{as }n\to\infty.$
Lemma 5.9
Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space which admits local addition and $p\in[1,\infty]$. Then $G$ is locally $L^{p}$-semiregular if and only if $G$ is $L^{p}$-semiregular.
Proof.
Let $G$ be locally $L^{p}$-semiregular. Then there is exists and 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that $\operatorname{Evol}(\beta):=\eta_{\beta}\in AC_{L^{p}}([a,b],G)$ exists of each $[\beta]\in B$. without loss of generalization we assume that
$B=\{[\beta]\in L^{p}([0,1],T_{e}G):\lVert[\beta]\lVert_{L^{p},q}<1\}$
for some continuous seminorm $q$ on $L^{p}([0,1],T_{e}G)$.
Let $[\gamma]\in L^{p}([0,1],T_{e}G)$. By Lemma 5.8, we find $n\in\mathbb{N}$ such that
$[\gamma_{n,k}]\in B,\quad\text{ for all }k\in\{0,1,...,n-1\}.$
Since each map
$\alpha_{k}:L^{p}([0,1],T_{e}G)\to L^{p}([0,1],T_{e}G),\quad[\tau]\mapsto[\tau_{n,k}]$
is continuous linear, there is exists an open $\gamma$-neighborhood $W_{\gamma}$ such that
$\alpha_{k}(W_{\gamma})\subseteq B,\quad\text{for all }k\in\{0,1,...,n-1\}.$
For $[\beta]\in W_{\gamma}$ and $j\in\{0,...,n-1\}$ we denote $\eta_{j}=\operatorname{Evol}(\beta_{n,j})$ and we define the function $\eta_{\beta}:[0,1]\to G$ via
$\eta_{\beta}(t):=\eta_{0}(nt),\quad\text{ if }t\in[0,1/n]$
and
$\eta_{\beta}(t):=\eta_{k}(nt-k).\Big{(}\eta_{k-1}(1)....\eta_{0}(1)\Big{)},\quad\text{ if }t\in[k/n,(k+1)/n]\text{ with }k\in\{1,..,n-1\}.$
Then the function $\eta_{\beta}$ is continuous and by Proposition 5.3 we have
$\eta_{\beta}|_{[k/n,(k+1)/n]}\in AC_{L^{p}}\left([k/n,(k+1)/n],G\right).$
Thus $\eta_{\beta}\in AC_{L^{p}}([a,b],G)$. If $t\in[0,\frac{1}{n}]$ we have $\eta_{\beta}(0)=e$ and
$\dot{\eta_{\beta}}(t)$ $=\dot{\eta_{0}}(nt)$
$=n\beta_{n,0}(nt).\eta_{0}(nt)$
$=n\frac{1}{n}\beta\left(\frac{0+nt}{n}\right).\eta_{0}(nt)$
$=\beta(t).\eta_{\beta}(t).$
For $k\in\{1,..,n-1\}$ we have
$\eta_{\beta}\left(\frac{k}{n}\right)=\eta_{k}\left(n\frac{k}{n}-k\right).\Big{(}\eta_{k-1}(1)....\eta_{0}(1)\Big{)}=\eta_{k-1}(1)....\eta_{0}(1)$
and for $t\in[\frac{k}{n},\frac{k+1}{n}]$
$\dot{\eta_{\beta}}(t)$ $=n\beta_{n,k}(nt-k)\Big{(}\eta_{k-1}(1)....\eta_{0}(1)\Big{)}.\eta_{k}(nt-k)$
$=\beta(t).\eta_{k}(nt-k)\Big{(}\eta_{k-1}(1)....\eta_{0}(1)\Big{)}$
$=\beta(t).\eta_{\beta}(t).$
Thus $\eta_{\beta}=\operatorname{Evol}(\beta)$ and in particular, $\eta_{\gamma}=\operatorname{Evol}(\gamma)$. Hence $G$ is $L^{p}$-semiregular.
The reciprocal is trivial. ∎
Definition 5.10
Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space which admits local addition and $p\in[1,\infty]$. We say that $G$ is locally $L^{p}$-regular if $G$ is $L^{p}$-semiregular and there is exists an 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that its restricted evolution map $\operatorname{Evol}|_{B}$ is smooth.
The following Lemma is just an application of [23, Proposition 4.11] to our case.
Lemma 5.11
Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space which admits local addition and $p\in[1,\infty]$. We consider the evolution map with continuous values
$\operatorname{Evol}_{C}:L^{p}([0,1],T_{e}G)\to C([0,1],G),\quad[\gamma]\mapsto\operatorname{Evol}_{C}(\gamma):=\eta_{\gamma}.$
Then, $\operatorname{Evol}_{C}$ is continuous if and only if there is exists an 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that its restricted evolution map $\operatorname{Evol}_{C}|_{B}$ is continuous.
Proof.
Since $G$ is a topological group, the map
$C([0,1],G)\times C([0,1],G)\to C([0,1],G),\quad(\eta,\xi)\mapsto\eta.(\xi(1))$
is continuous. Following Lemma 5.9, for each $[\beta]\in W_{\gamma}$, we have that the construction of the function $\operatorname{Evol}(\beta)$ implies that $Evol_{C}$ is just the product of composition of continuous maps. ∎
Lemma 5.12
Let $E_{1},E_{2}$ and $F$ be locally convex spaces, $1\leq p<\infty$ and $U\subseteq E_{2}$ an open subset. If $r\in\mathbb{N}\cup\{0,\infty\}$ and $f:E_{1}\times U\to F$ is a $C^{r}$-map such that for each $y\in U$ the map $f(\cdot,y):E_{1}\to F$ is linear, then the map
$\tilde{f}:L^{p}([0,1],E_{1})\times C([0,1],U)\to L^{p}([0,1],F),\quad([\gamma],\eta)\mapsto f\circ([\gamma],\eta)$
is $C^{r}$.
Proof.
Let $(\gamma_{0},\eta_{0})\in L^{p}([0,1],E_{1})\times C([0,1],U)$ and $\varepsilon>0$. Let $\beta:F\to[0,\infty)$ be a continuous semi norm and $K:=\eta_{0}([0,1])\subseteq U$. Since $f$ is continuous and linear in the first argument, for $y\in K$ there exists a $y$-neighborhood $V_{y}\subseteq U$ and a continuous seminorm $\kappa_{y}:E_{1}\to[0,\infty)$ such that
$f\left(B_{1}^{\kappa_{y}}(0)\times V_{y}\right)\subseteq B_{1}^{\beta}(0).$
By compactness of $K$, there exists a finite numbers of $y_{1},...,y_{n}\in K$ such that
$K\subseteq V:=\cup_{i=1}^{n}V_{y_{i}}.$
We define the continuous seminorm $\kappa:E_{1}\to[0,\infty)$ by $\kappa:=\kappa_{y_{1}}+...+\kappa_{y_{n}}$ and we have
$\beta\big{(}f(x,b)\big{)}\leq\kappa(x),\quad\text{for all }x\in E_{1},b\in V$
and $C([0,1],V)$ is open in $C([0,1],U)$. We estimate
[…]
$\leq\lVert\gamma_{0}\lVert_{L^{p},\kappa}.$
Hence $\tilde{f}(\gamma_{0},\eta_{0})\in L^{p}([0,1],F)$. Since the space $C([0,1],E_{1})$ is dense in $L^{p}([0,1],E_{1})$, there exists a $\gamma_{c}\in C([0,1],E_{1})$ such that
$\lVert\gamma_{c}-\gamma_{0}\lVert_{L^{p},p}\leq 2\varepsilon/5.$
Additionally, since the maps
$C([0,1],U)\to C([0,1],F),\quad\eta\mapsto f\circ(\gamma_{c},\eta)$
and the inclusion map $C([0,1],F)\to L^{p}([0,1],F)$ are continuous, we have that
$C([0,1],V)\to L^{p}([0,1],F),\quad\eta\mapsto f\circ(\gamma_{c},\eta)$
is continuous. Hence there exists an open neighborhood $W$ of $\eta_{0}\in C([0,1],V)$ such that
$\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{c},\eta_{0})\lVert_{L^{p},\beta}\leq 2\varepsilon/5,\quad\text{for all }\eta\in W.$
With this, for each $\eta\in W$ and $\gamma\in L^{p}([0,1],E_{1})$ with $\lVert\gamma-\gamma_{0}\lVert_{L^{p},\kappa}\leq 2\varepsilon/5$ we estimate
$\lVert\tilde{f}(\gamma,\eta)-\tilde{f}(\gamma_{0},\eta_{0})\lVert_{L^{p},\beta}$ $\leq\lVert\tilde{f}(\gamma-\gamma_{c},\eta)\lVert_{L^{p},\beta}+\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{0},\eta_{0})\lVert_{L^{p},\beta}$
$\leq\lVert\tilde{f}(\gamma-\gamma_{c},\eta)\lVert_{L^{p},\beta}+\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{c},\eta_{0})\lVert_{L^{p},\beta}+\lVert\tilde{f}(\gamma_{c}-\gamma_{0},\eta_{0})\lVert_{L^{p},\beta}$
$\leq\lVert\gamma-\gamma_{c}\lVert_{L^{p},\kappa}+\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{c},\eta_{0})\lVert_{L^{p},\beta}+\lVert\gamma_{c}-\gamma_{0}\lVert_{L^{p},\kappa}.$
Since $\lVert\gamma_{c}-\gamma_{0}\lVert_{L^{p},\kappa}\leq 2\varepsilon/10$ and $\lVert\gamma-\gamma_{0}\lVert_{L^{p},\kappa}\leq 2\varepsilon/10$ we have
$\lVert\gamma-\gamma_{c}\lVert_{L^{p},\kappa}\leq 2\varepsilon/5$
and
$\lVert\tilde{f}(\gamma,\eta)-\tilde{f}(\gamma_{0},\eta_{0})\lVert_{L^{p},\beta}\leq 2\varepsilon/5+2\varepsilon/5+2\varepsilon/10=\varepsilon$
Thus $\tilde{f}$ is continuous. By linearity in the first variable the function $\overline{f}$ has a continuous differential $d_{1}\tilde{f}:L^{p}([0,1],E_{1})\times L^{p}([0,1],E_{1})\times C([0,1],E_{2})\to L^{p}([0,1],F)$ in the first variable. Let consider the map $f$ as a $C^{1}$-map. If $x\in E_{1}$, $y,y_{1}\in E_{2}$ and $t\in\mathbb{R}^{\times}$, then we have
[…]
whenever $y+[0,1]ty_{1}\subseteq U$. Given that the map $d_{2}f:E_{1}\times U\times E_{2}\to F$ is continuous, identifying $C([0,1],E_{1}\times E_{2})$ with $C([0,1],E_{1})\times C([0,1],E_{2})$, we have that the map
$\widetilde{d_{2}f}:L^{p}([0,1],E_{1})\times C([0,1],U)\times C([0,1],E_{2})\to L^{p}([0,1],F),\quad(\gamma,\eta,\eta_{1})\mapsto d_{2}f(\gamma,\eta,\eta_{1})$
is continuous. For $t\in[0,1]$ we denote
$g_{t}:C([0,1],F)\to F,\quad\gamma\mapsto\gamma(t).$
Since the family of maps $\{g_{t}:t\in[0,1]\}$ separate points on $C([0,1],F)$, we have that the equality
[…]
is valid for each $\gamma_{c}\in C([0,1],E_{1})$, $\eta\in C([0,1],U)$, $\eta_{1}\in C([0,1],E_{2})$ and $t\in\mathbb{R}^{\times}$ such that
$\eta+[0,1]t\eta_{1}\in C([0,1],U).$
By density of $C([0,1],E_{1})$ on $L^{p}([0,1],E_{1})$ and continuity of $\tilde{f}$, for $\gamma\in L^{p}([0,1],E)$ the equation verifies
[…]
Let $\gamma$, $\eta$ and $\eta_{0}$ be fixed, then the map $(t,s)\mapsto\widetilde{d_{2}f}(\gamma,\eta+ts\eta_{1},\eta_{1})$ is continuous, including in $t=0$. Then, taken the limit $t\to 0$ in the equality we obtain
$d_{2}\tilde{f}(\gamma,\eta,\eta_{1})=\widetilde{d_{2}f}(\gamma,\eta,\eta_{1}).$
Hence the continuity of $\widetilde{d_{2}f}$ implies that the map $\overline{f}$ is $C^{1}$. Proceeding by induction, if $f$ is a $C^{k}$-map, since $\tilde{f}$ is linear in the first variable we have
$d\tilde{f}(\gamma,\eta,\gamma_{1},\eta_{1})$ $=d_{1}\tilde{f}(\gamma,\gamma_{1},\eta)+d_{2}\tilde{f}(\gamma,\eta,\eta_{1})$
$=\tilde{f}(\gamma_{1},\eta)+\widetilde{d_{2}f}(\gamma,\eta,\eta_{1}).$
By the induction hypothesis $\tilde{f}$ and $d_{2}f$ are $C^{r-1}$, hence $\widetilde{d_{2}f}$ is $C^{r-1}$ with $\widetilde{d_{2}f}=d_{2}\tilde{f}$, thus $\tilde{f}$ is $C^{r}$. ∎
For the case $L_{rc}^{\infty}$ we recall [10, Proposition 2.3].
Lemma 5.13
Let $E_{1},E_{2}$ and $F$ be integral complete locally convex spaces and $U\subseteq E_{2}$ an open subset. If $r\in\mathbb{N}\cup\{0,\infty\}$ and $f:E_{1}\times U\to F$ is a $C^{r}$-map such that for each $y\in U$ the map $f(\cdot,y):E_{1}\to F$ is linear, then the map
$\tilde{f}:L_{rc}^{\infty}([0,1],E_{1})\times C([0,1],U)\to L_{rc}^{\infty}([0,1],F),\quad([\gamma],\eta)\mapsto f\circ([\gamma],\eta)$
is $C^{r}$.
Theorem 5.14
Let $G$ be a right half-Lie group modeled in a Fréchet space $E$ which admits local addition and $1\leq p<\infty$. Let $G$ be $L^{p}$-semiregular with continuous evolution map
$\operatorname{Evol}_{C}:L^{p}([0,1],T_{e}G),\to C([0,1],G),\quad\gamma\mapsto\operatorname{Evol}_{C}(\gamma).$
If the restriction of the right action
$\tau:T_{e}G\times G\to TG,\quad(v,g)\mapsto v.g$
is continuous, then the evolution map
$\operatorname{Evol}:L^{p}([0,1],T_{e}G),\to AC_{L^{p}}([0,1],G),\quad\gamma\mapsto\operatorname{Evol}(\gamma)$
is continuous.
If we replace $L^{p}$ with $L_{rc}^{\infty}$ the result remains valid.
Proof.
Let $[\gamma]\in L^{p}([0,1],T_{e}G)$. Let $\mathcal{P}_{[}\gamma]$ be a open neighborhood of $\operatorname{Evol}_{C}(\gamma)$ in $C([0,1],G)$, then there exists a partition $P=\{t_{0},...,t_{n}\}$ of $[0,1]$ and a family of charts $\varphi_{i}:U_{i}\to V_{i}$ such that
$\mathcal{P}_{[\gamma]}=\bigcap_{i=1}^{n}\{\eta\in C([t_{i-1},t_{i}],U_{i}):\eta([t_{i-1},t_{i}])\subseteq U_{i}\}.$
Let $\mathcal{Q}_{[\gamma]}=\operatorname{Evol}_{C}^{-1}(\mathcal{P}_{[\gamma]})$, then $\mathcal{Q}_{[\gamma]}$ is an open neighborhood of ${[\gamma]}$ in $L^{p}([0,1],T_{e}G)$. We will show that the map
$\operatorname{Evol}|_{\mathcal{Q}_{[\gamma]}}:\mathcal{Q}_{[\gamma]}\to\mathcal{P}_{[\gamma]}\cap AC_{L^{p}}([0,1],G),\quad\gamma\mapsto\operatorname{Evol}(\gamma)$
is continuous. The map
$\Phi_{1}:\mathcal{Q}_{[\gamma]}\to\prod_{i=1}^{n}L^{p}([t_{i-1},t_{i}],T_{e}G),\quad[\beta]\mapsto\left([\beta|_{[t_{i-1},t_{i}]}]\right)_{i=1}^{n}$
is a topological embedding. For each $i\in\{1,...,n\}$ we have
$\operatorname{Evol}_{C}(\beta|_{[t_{i-1},t_{i}]})=\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]},\quad\text{ for all }[\beta]\in\mathcal{Q}_{[\gamma]}$
and each map $C([t_{i-1},t_{i}],U_{i})\to C([t_{i-1},t_{i}],V_{i})$, $\eta\mapsto\varphi_{i}\circ\eta$ is a homeomorphism. This allow us to define the continuous map
$\Phi_{2}:\mathcal{Q}_{[\gamma]}\to\prod_{i=1}^{n}C([t_{i-1},t_{i}],V_{i}),\quad[\beta]\mapsto\left(\varphi_{i}\circ\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}.$
We consider the map
$f_{i}:T_{e}G\times V_{i}\to E,\quad(v,x)\mapsto d\varphi_{i}\circ\tau(v,\varphi_{i}^{-1}(x)).$
Then each $f_{i}$ is continuous and for each $x\in V_{i}$ the function $f_{i}(\cdot,x):T_{e}G\to E$, $v\mapsto f_{i}(v,x)$ is linear. Hence by Lemma 5.12 the map
$\tilde{f}_{i}:L^{p}([t_{i-1},t_{i}],T_{e}G)\times C([t_{i-1},t_{i}],V_{i})\to L^{p}([t_{i-1},t_{i}],E),\quad([\beta],\eta)\mapsto f_{i}\circ([\beta],\eta)$
is continuous. We denote
$F:\prod_{i=1}^{n}L^{p}([t_{i-1},t_{i}],T_{e}G)\times C([t_{i-1},t_{i}],V_{i})\to\prod_{i=1}^{n}L^{p}([t_{i-1},t_{i}],E),\quad\left([\beta_{i}],\eta_{i}\right)_{i=1}^{n}\mapsto\left(f_{i}\circ([\beta_{i}],\eta_{i}\right)_{i=1}^{n}.$
Then $F$ is continuous, where for each $i\in\{1,...,n\}$ we have
$f_{i}\circ([\beta_{i}],\eta_{i})=d\varphi_{i}\circ\tau([\beta_{i}],\varphi_{i}^{-1}\circ\eta_{i}).$
This allow us to define the continuous map
$\Phi_{3}:\mathcal{Q}_{[\gamma]}$ $\to\prod_{i=1}^{n}L^{p}([t_{i-1},t_{i}],E)\times C([t_{i-1},t_{i}],V_{i}),$
$\beta$ $\mapsto\Big{(}f_{i}\circ\left(\left[\beta|_{[t_{i-1},t_{i}]}\right],\varphi_{i}\circ\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\right),\varphi_{i}\circ\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\Big{)}_{i=1}^{n}$
where for each $i\in\{1,...,n\}$ we have
$f_{i}\circ\left(\left[\beta|_{[t_{i-1},t_{i}]}\right],\varphi_{i}\circ\operatorname{Evol}_{C}|_{[t_{i-1},t_{i}]}\right)$ $=d\varphi_{i}\circ\tau\left(\left[\beta|_{[t_{i-1},t_{i}]}\right],\varphi_{i}^{-1}\circ\varphi_{i}\circ\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\right)$
$=d\varphi_{i}\circ\tau\left(\left[\beta|_{[t_{i-1},t_{i}]}\right],\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\right)$
$=d\varphi_{i}\circ\left(\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\right)^{\prime}$
$=\left(\varphi_{i}\circ\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\right)^{\prime}.$
Hence
$\Phi_{3}([\beta])=\Big{(}\left(\varphi_{i}\circ\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\right)^{\prime},\varphi_{i}\circ\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\Big{)}_{i=1}^{n},\quad\text{ for all }[\beta]\in\mathcal{Q}_{[\gamma]}.$
We set the topological embedding (see Lemma 2.12)
$\Psi_{i}:AC_{L^{p}}([t_{i-1},t_{i}],V_{i})\to L^{p}([t_{i-1},t_{i}],V_{i})\times C([t_{i-1},t_{i}],V_{i}),\quad\alpha\mapsto(\alpha^{\prime},\alpha)$
and
$\Psi:\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],V_{i})\to\prod_{i=1}^{n}L^{p}([t_{i-1},t_{i}],T_{e}G)\times C([t_{i-1},t_{i}],V_{i}),\quad(\alpha_{i})_{i=1}^{n}\mapsto(\alpha_{i}^{\prime},\alpha_{i})_{i=1}^{n}.$
Thus $\text{Im}(\Phi_{3})\subseteq\text{Im}(\Psi)$. For each $i\in\{1,...,n\}$ we note that
$\left(\Psi_{i}|^{\text{Im}(\Psi_{i})}\right)^{-1}\Big{(}(\varphi_{i}\circ\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]})^{\prime},\varphi_{i}\circ\operatorname{Evol}_{C}(\beta)|_{[t_{i-1},t_{i}]}\Big{)}=\varphi_{i}\circ\operatorname{Evol}(\beta)|_{[t_{i-1},t_{i}]}$
for all $[\beta]\in\mathcal{Q}_{[\gamma]}$. We set the continuous map
$\Phi_{4}:=\left(\Psi|^{\text{Im}(\Psi)}\right)^{-1}\circ\Phi_{3}:\mathcal{Q}_{[\gamma]}\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],V_{i}),\quad[\beta]\mapsto\left(\varphi_{i}\circ\operatorname{Evol}(\beta)|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}.$
Let denote the homeomorphism onto its image (see Proposition 3.9)
$\Phi_{5}:AC_{L^{p}}([0,1],G)\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],G),\quad\eta\mapsto(\eta|_{[t_{i-1},t_{i}]})_{i=1}^{n}.$
and the homeomorphism
$\Phi_{6}:\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],V_{i})\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],U_{i}),\quad\tau_{i}\mapsto(\varphi_{i}^{-1}\circ\tau_{i})_{i=1}^{n}.$
We see that
$\Phi_{6}\circ\Phi_{4}:\mathcal{Q}_{[\gamma]}\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],U_{i}),\quad[\beta]\mapsto\left(\operatorname{Evol}(\beta)|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}.$
Since each function $\operatorname{Evol}(\beta)$ is continuous, we have that $(\Phi_{6}\circ\Phi_{4})\left(\mathcal{Q}_{[\gamma]}\right)\subseteq\text{Im}(\Phi_{5})$, hence
$\operatorname{Evol}|_{\mathcal{Q}_{[\gamma]}}=\left(\Phi_{5}|^{\text{Im}(\Phi_{5})}\right)^{-1}\circ\Phi_{6}\circ\Phi_{4}.$
Thus the evolution map $\operatorname{Evol}$ is continuous.
For the case $L_{rc}^{\infty}$, by Lemma 5.13, the proof is analogous since the map
$\tilde{f}_{i}:L_{rc}^{\infty}([t_{i-1},t_{i}],T_{e}G)\times C([t_{i-1},t_{i}],V_{i})\to L_{rc}^{\infty}([t_{i-1},t_{i}],E),\quad([\beta],\eta)\mapsto f_{i}\circ([\beta],\eta)$
is continous for each $i\in\{1,...,n\}$. ∎
6 Semiregularity of $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$
Let $n,m,r\in\mathbb{N}$. We set the Fréchet space $C(\mathbb{R}^{n},\mathbb{R}^{m})$ of continuous functions $\phi:\mathbb{R}^{n}\to\mathbb{R}^{m}$, whose topology is generated by the famility of seminorms
$\lVert\cdot\lVert_{L}:C(\mathbb{R}^{n},\mathbb{R}^{m})\to[0,\infty),\quad\phi\mapsto\sup_{x\in L}\lvert\phi(x)\lvert$
for each non empty compact subset $L\subset\mathbb{R}^{n}$. Let $C^{r}(\mathbb{R}^{n},\mathbb{R}^{m})$ be the Fréchet space of all $C^{r}$-maps $\phi:\mathbb{R}^{n}\to\mathbb{R}^{n}$, whose topology is the compact-open $C^{r}$-topology, i.e., the initial topology with respect to the maps
$C^{r}(\mathbb{R}^{n},\mathbb{R}^{m})\to C(\mathbb{R}^{n},\mathbb{R}^{m}),\quad\phi\mapsto\frac{\partial^{\alpha}\phi}{\partial x^{\alpha}}$
for $\alpha\in\mathbb{N}_{0}^{n}$ with $\lvert\alpha\lvert\leq r$. Let $K\subseteq\mathbb{R}^{n}$ be a non empty compact subset, we define the Banach space of $C^{r}$-maps supported in $K$ as
$C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{m}):=\{\phi\in C^{r}(\mathbb{R}^{n},\mathbb{R}^{m}):\phi|_{\mathbb{R}^{n}\setminus K}=0\}$
endowed with the induced topology.
We denote derivative map
$D:C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))\to C_{K}^{r-1}(\mathbb{R}^{n},\mathbb{R}^{n\times n}),\quad f\mapsto Df:=f^{\prime}$
where $f^{\prime}(x)$ is the jacobian matrix of $f$.
Definition 6.1
Let $\text{Diff}^{r}(\mathbb{R}^{n})$ be the set of $C^{r}$-diffeomorphisms $\phi:\mathbb{R}^{n}\to\mathbb{R}^{n}$. The set $\text{Diff}^{r}(\mathbb{R}^{n})$ is a group under the composition and we define the subgroup $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$ of $C^{r}$-diffeomorphisms which are supported in a compact set $K\subset\mathbb{R}^{n}$ as
$\text{Diff}_{K}^{r}(\mathbb{R}^{n}):=\{\phi\in\text{Diff}^{r}(\mathbb{R}^{n}):\phi-\text{id}_{\mathbb{R}^{n}}\in C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\}.$
Let $\mathcal{V}_{K}:=\{\phi-\text{id}_{\mathbb{R}^{n}}\in C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}):\phi\in\text{Diff}_{K}^{r}(\mathbb{R}^{n})\}$, then $\mathcal{V}_{K}$ is open in $C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$ and the map
$\Phi:\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})\to\mathcal{V}_{K},\quad\phi\mapsto\phi-\text{id}_{\mathbb{R}^{n}}$
is a global chart for $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$, turning it into a right half-Lie group modeled on the Banach space $C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$ (See [12, Proposition 14.6]). On the set $\mathcal{V}_{K}$ we define a group multiplication by
$\phi\ast\psi:=\Phi\left(\Phi^{-1}(\phi)\circ\Phi^{-1}(\psi)\right)=\psi+\phi\circ(\text{id}_{\mathbb{R}^{n}}+\psi),\quad\text{ for each }\phi,\psi\in\mathcal{V}_{K}$
with the constant function 00 as neutral element.
Remark 6.2
We will study the $L^{1}$-semiregularity of $\mathcal{V}_{K}$ instead of $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$. Since $\mathcal{V}_{K}$ is an open set of a locally convex space, we have
$T\mathcal{V}_{K}=\mathcal{V}_{K}\times C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}).$
For $\psi\in\mathcal{V}_{K}$ fixed, we have the right translation
$\rho_{\psi}:\mathcal{V}_{K}\to\mathcal{V}_{K},\quad\phi\mapsto\psi+\phi\circ(\text{id}_{\mathbb{R}^{n}}+\psi)$
and its derivative
$d\rho_{\psi}:\mathcal{V}_{K}\times C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\to C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}),\quad(\phi,\varphi)\mapsto\varphi\circ(\text{id}_{\mathbb{R}^{n}}+\psi).$
Identifying $T_{0}\mathcal{V}_{K}$ with $\{0\}\times C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$, the restricted right action is given by
($\left(\{0\}\times C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\right)\times\mathcal{V}_{K}\to T\mathcal{V}_{K},\quad((0,\varphi),\psi)\mapsto d\rho_{\psi}(0,\varphi)=\varphi\circ(id_{\mathbb{R}^{n}}+\psi)$
Hence the right action is continuous (see e.g. [11]).
If $(\phi,\varphi)\in\mathcal{V}_{K}\times C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$, we have
$(\phi,\varphi).\psi:=T\rho_{\psi}(\phi,\varphi)=(\phi\ast\psi,\varphi\circ(\text{id}_{\mathbb{R}^{n}}+\psi)).$
If $(0,\gamma)\in\mathcal{L}^{1}([0,1],T_{0}\mathcal{V}_{K})$, we want to find a function $\eta\in AC_{L^{p}}([0,1],\mathcal{V}_{K})$ such that $\eta^{\prime}=\gamma.\eta$, i.e.
$(\eta(t),\eta^{\prime}(t))=\Big{(}\eta(t),\gamma(t)\circ(\text{id}_{\mathbb{R}^{n}}+\eta(t))\Big{)},\text{ for almost all }t\in[0,1].$
In other words, for $\gamma\in\mathcal{L}^{1}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))$, we need a function $\eta\in AC_{L^{p}}([0,1],\mathcal{V}_{K})$ such that $\eta(0)=0$ and
[…]
Setting $\zeta=\text{id}_{\mathbb{R}^{n}}+\eta(s)$, this is equivalent to
[…]
Let $\varepsilon_{x}:C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\to\mathbb{R}^{n}$ be the point evaluation map for $x\in\mathbb{R}^{n}$, then each $\varepsilon_{x}$ is continuous linear and the family of maps $(\varepsilon_{x})_{x\in\mathbb{R}^{n}}$ separate points on $C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$, hence the equation holds if and only if the functions $\zeta_{x}:=\varepsilon_{x}\circ\zeta\in AC_{L^{1}}([0,1],\mathbb{R}^{n})$ satisfy
[…]
Theorem 6.3
Let $1\leq p<\infty$. If $r\in\mathbb{N}$, then the right half-Lie group $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$ is $L^{p}$-semiregular. Moreover, the evolution map
$\operatorname{Evol}:L^{p}([0,1],T_{e}\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n}))\to AC_{L^{p}}([0,1],\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})),\quad\gamma\mapsto\eta_{\gamma}.$
is continuous.
Proof.
Following the discussion in Remark 6.2, we will show that $\mathcal{V}_{K}$ is locally $L^{p}$-semiregular.
We define the continuous seminorm
$\alpha:C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\to[0,\infty),\quad\alpha(f):=\lVert f\lVert_{\mathcal{L}^{\infty},\lVert\cdot\lVert_{op}}=\sup_{x\in\mathbb{R}^{n}}\lVert Df(x)\lVert_{op}.$
For $10<L<$ 1, we denote the open ball centered in $0\in L^{p}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))$ by
$B_{L}:=\left\{[\gamma]\in L^{p}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})):\lVert\gamma\lVert_{L^{p},\alpha}:=\left(\int_{0}^{1}(\alpha\circ\gamma)^{p}(t)dt\right)^{1/p}<L\right\}.$
For $x\in\mathbb{R}^{n}$, we define the smooth map
$c:\mathbb{R}^{n}\to C([0,1],\mathbb{R}^{n}),\quad x\mapsto[t\mapsto x].$
Let $J:L^{1}([0,1],\mathbb{R}^{n})\to C([0,1],\mathbb{R}^{n})$ be the continuous linear operator
[…]
Since the evaluation map $\varepsilon:C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ is $C^{r,\infty}$, by Lemma 5.12 the map
$\Phi_{n}:L^{p}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))\times C([0,1],\mathbb{R}^{n})\to L^{p}([0,1],\mathbb{R}^{n}),\quad\Phi_{n}([\gamma],\zeta):=\varepsilon\circ([\gamma],\zeta)$
is well defined and is $C^{r}$. We define the operator $T:B_{L}\times\mathbb{R}^{n}\times C([0,1],\mathbb{R}^{n})\to C([0,1],\mathbb{R}^{n})$ via
$T([\gamma],x,\zeta)(t)$ $:=c_{x}(t)+J(\Phi_{n}([\gamma],\zeta))(t)$
[…]
for $t\in[0,1]$. Then $T$ is $C^{\infty,\infty,r}$.
Let $[\gamma]\in B_{L}$ and $x\in\mathbb{R}^{n}$ be fixed, for $\zeta_{1},\zeta_{2}\in C([0,1],\mathbb{R}^{n})$, by the Mean Value Theorem we have
$\gamma(t)\circ\zeta_{2}(t)-\gamma(t)\circ\zeta_{1}(t)=\int_{0}^{1}\Big{(}D\gamma(t)\Big{)}(\zeta_{1}(t)+s(\zeta_{2}(t)-\zeta_{1}(t))).(\zeta_{2}(t)-\zeta_{1}(t))ds.$
Thus
$\lVert T([\gamma],x,\zeta_{2})-T([\gamma],x,\zeta_{1})\lVert_{\infty}$ $=\sup_{t\in[0,1]}\left\lVert\int_{0}^{t}(\gamma(s)\circ\zeta_{2}(s)-\gamma(s)\circ\zeta_{1}(s))ds\right\lVert_{\infty}$
$\leq\sup_{t\in[0,1]}\int_{0}^{1}\left\lVert\Big{(}D\gamma(t)\Big{)}(\zeta_{1}(t)+s(\zeta_{2}(t)-\zeta_{1}(t))).(\zeta_{2}(t)-\zeta_{1}(t))\right\lVert_{\infty}ds$
$\leq\sup_{t\in[0,1]}\int_{0}^{1}\lVert D\gamma(t)\lVert_{\mathcal{L}^{\infty},\lVert\cdot\lVert_{op}}\lVert\zeta_{2}-\zeta_{1}\lVert_{\infty}dt$
$=\sup_{t\in[0,1]}\int_{0}^{1}\alpha(\gamma(t))dt\lVert\zeta_{2}-\zeta_{1}\lVert_{\infty}$
$\leq\lVert\gamma\lVert_{L^{1},\alpha}\lVert\zeta_{2}-\zeta_{1}\lVert_{\infty}$
$\leq\lVert\gamma\lVert_{L^{p},\alpha}\lVert\zeta_{2}-\zeta_{1}\lVert_{\infty}$
$\leq L\lVert\zeta_{2}-\zeta_{1}\lVert_{\infty}.$
Hence $T([\gamma],x,\cdot)$ is a $C^{r}$-map and a $L$-contraction. By Banach’s Fixed Point Theorem, there exists $\zeta_{[\gamma],x}\in C([0,1],\mathbb{R}^{n})$ such that
$T([\gamma],x,\zeta_{[\gamma],x})=\zeta_{[\gamma],x}.$
By [10, Lemma 6.2] the map
$F_{1}:B_{L}\times\mathbb{R}^{n}\to C([0,1],\mathbb{R}^{n}),\quad([\gamma],x)\mapsto\zeta_{[\gamma],x}$
is $C^{r}$. By the exponential laws, we have:
  • a)
    $F_{2}:\left(B_{L}\times\mathbb{R}^{n}\right)\times[0,1]\to\mathbb{R}^{n},\quad([\gamma],x,t)\mapsto\zeta_{[\gamma],x}(t)$ is $C^{r,0}$.
  • b)
    $F_{3}:B_{L}\times[0,1]\to C^{r}(\mathbb{R}^{n},\mathbb{R}^{n}),\quad([\gamma],t)\mapsto\zeta_{[\gamma],\cdot}(t)$ is $C^{r,0}$.
  • c)
    $F_{4}:B_{L}\to C([0,1],C^{r}(\mathbb{R}^{n},\mathbb{R}^{n})),\quad[\gamma]\mapsto\zeta_{[\gamma]}$ is continuous.
Let $x\in\mathbb{R}^{n}\setminus K$ and $\zeta_{[\gamma,x}$ be the fixed point of
[…]
Since $\gamma(t)(x)=0$ for each $t\in[0,1]$, the constant map $c_{x}$ is also a fixed point and by uniqueness of solutions we have $\zeta_{[\gamma],x}=c_{x}$. Let $\text{I}:[0,1]\to C^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$, $t\mapsto\text{id}_{\mathbb{R}^{n}}$ the constant map mapping to the identity map. We define
$S:B\to C([0,1],C^{r}(\mathbb{R}^{n},\mathbb{R}^{n})),\quad[\gamma]\mapsto F_{4}([\gamma])-\text{I}.$
Then for each $t\in[0,1]$ we have $S([\gamma])(t)\in C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$. Moreover, since $S$ is continuous and $\mathcal{V}_{K}$ is open, there is exists an 0-neighborhood $B\subseteq B_{L}$ such that $S(B)\subseteq\mathcal{V}_{K}$. Hence, we can consider the map $S$ as the continuous map
$\tilde{S}:B\to C([0,1],\mathcal{V}_{K}),\quad[\gamma]\mapsto F_{4}([\gamma])-\text{I},$
where, by the discussion in Remark 6.2, if $[\gamma]\in B$ we have that $\eta:=\tilde{S}([\gamma])$ is solution of
[…]
Hence $\operatorname{Evol}(\gamma)=\eta$. In consequence, $\mathcal{V_{K}}$ is locally $L^{p}$-semiregular and thus, by Lemma 5.9, the half-Lie group $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$ is $L^{p}$-semiregular.
Moreover, the evolution map restricted to the 00-neighborhood $B$ is given by
$\operatorname{Evol}|_{B}:B\subseteq L^{p}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))\to AC_{L^{p}}([0,1],\mathcal{V}_{K}),\quad\gamma\mapsto\tilde{S}(\gamma).$
Since $\tilde{S}$ is continuous, by Lemma 5.11 the evolution map with continuous values $\operatorname{Evol}_{C}$ is continuous. Moreover, since the restricted right action of $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$ is continuous, by Theorem 5.14, the evolution map
$\operatorname{Evol}:L^{p}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))\to AC_{L^{p}}([0,1],\mathcal{V}_{K}),\quad\gamma\mapsto\operatorname{Evol}(\gamma)$
is continuous. ∎
Proceeding exactly as the case $L^{p}$, for the case $L_{rc}^{\infty}$ we have the same result.
Proposition 6.4
If $r\in\mathbb{N}$, then the right half-Lie group $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$ is $L_{rc}^{\infty}$-semiregular. Moreover, the evolution map
$\operatorname{Evol}:L_{rc}^{\infty}([0,1],T_{e}\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n}))\to AC_{L_{rc}^{\infty}}([0,1],\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})),\quad\gamma\mapsto\eta_{\gamma}$
is continuous.
Proof.
For $10<L<$ 1, we define the open ball centered in $0\in L_{rc}^{\infty}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))$ via
$B_{L}:=\{[\gamma]\in L_{rc}^{\infty}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})):\lVert\gamma\lVert_{L_{rc}^{\infty},\alpha}:=\operatorname*{ess\,sup}_{t\in[0,1]}\alpha(\gamma(t))<L\}$
Following the same notation as Proposition 1.7, since the evaluation map $\varepsilon$ is $C^{r,\infty}$, by Lemma 5.13 the map
$\Phi_{n}:L_{rc}^{\infty}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))\times C([0,1],\mathbb{R}^{n})\to L_{rc}^{\infty}([0,1],\mathbb{R}^{n}),\quad\Phi_{n}([\gamma],\zeta):=\varepsilon\circ([\gamma,\zeta])$
is $C^{r}$. Hence the operator $T:B_{L}\times\mathbb{R}^{n}\times C([0,1],\mathbb{R}^{n})\to C([0,1],\mathbb{R}^{n})$ given by
[…]
is a $C^{\infty,\infty,r}$-map. Let $[\gamma]\in L_{rc}^{\infty}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))$ and $x\in\mathbb{R}^{n}$ be fixed, for $\zeta_{1},\zeta_{2}\in C([0,1],\mathbb{R}^{n})$, by the Mean Value Theorem we have
$\gamma(t)\circ\zeta_{2}(t)-\gamma(t)\circ\zeta_{1}(t)=\int_{0}^{1}\Big{(}D\gamma(t)\Big{)}(\zeta_{1}(t)+s(\zeta_{2}(t)-\zeta_{1}(t))).(\zeta_{2}(t)-\zeta_{1}(t))ds$
and
$\lVert\Phi_{n}([\gamma],\zeta_{2})(t)-\Phi_{n}([\gamma],\zeta_{1})(t)\lVert_{\infty}$ $=\lVert\gamma(t)\circ\zeta_{2}(t)-\gamma(t)\circ\zeta_{1}(t)\lVert_{\infty}$
$\leq\int_{0}^{1}\lVert\Big{(}D\gamma(t)\Big{)}(\zeta_{1}(t)+s(\zeta_{2}(t)-\zeta_{1}(t))).(\zeta_{2}(t)-\zeta_{1}(t))\lVert_{\infty}ds$
$\leq\lVert D\gamma(t)\lVert_{\mathcal{L}^{\infty},\lVert\cdot\lVert_{op}}\lVert\zeta_{2}-\zeta_{1}\lVert_{\mathcal{L}^{\infty},\lVert\cdot\lVert_{\infty}}$
$=\alpha(\gamma(t))\lVert\zeta_{2}-\zeta_{1}\lVert_{\mathcal{L}^{\infty},\lVert\cdot\lVert_{\infty}}.$
Hence
$\lVert\Phi_{n}([\gamma],\zeta_{2})-\Phi_{n}([\gamma],\zeta_{1})\lVert_{L_{rc}^{\infty},\infty}\leq\lVert\gamma\lVert_{L_{rc}^{\infty},\alpha}\lVert\zeta_{2}-\zeta_{1}\lVert_{\mathcal{L}^{\infty},\lVert\cdot\lVert_{\infty}}.$
Thus $T([\gamma],x,\cdot)$ is a $C^{r}$-map and a $L$-contraction. With this, following the same steps as Proposition 1.7 we can show that $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$ is $L_{rc}^{\infty}$-semiregular and there exists a 00-neighborhood $B$ in $L_{rc}^{\infty}([0,1],T_{e}\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n}))$ such that the restricted map evolution map
$\operatorname{Evol}_{C}|_{B}:B\to AC_{L_{rc}^{\infty}}([0,1],\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})),\quad[\gamma]\mapsto\operatorname{Evol}_{C}(\gamma)$
is continuous. By Lemma 5.11 the evolution map with continuous values $\operatorname{Evol}_{C}$ is continuous. Moreover, since the restricted right action of $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$ is continuous, by Theorem 5.14, the evolution map
$\operatorname{Evol}:L_{rc}^{\infty}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))\to AC_{L_{rc}^{\infty}}([0,1],\mathcal{V}_{K}),\quad\gamma\mapsto\operatorname{Evol}(\gamma)$
is continuous. ∎
7 $L^{p}$-Semiregularity of $\text{Diff}^{\text{ }r}(M)$
Let $r\in\mathbb{N}$. Let $M$ be a compact smooth manifold without boundary, $g$ be a smooth Riemannian metric on $M$ and $\exp:\mathcal{D}\subseteq TM\to M$ be the Riemannian exponential map, with $\exp_{p}:=\exp|_{T_{p}M}$ for each $p\in M$. Let $\pi_{TM}:TM\to M$ be the bundle projection and $\mathcal{W}\subseteq\mathcal{D}$ be an open neighborhood of the zero-section such that the map
$(\pi_{TM},\exp):\mathcal{W}\subseteq TM\to M\times M,\quad v\mapsto(\pi_{TM}(v),\exp(v))$
is a diffeomorphism into its image. In particular, if $\mathcal{W}_{p}:=\mathcal{W}\cap T_{p}M$, the map
$\exp_{p}|_{\mathcal{W}_{p}}:\mathcal{W}_{p}\subseteq T_{p}M\to\exp_{p}\left(\mathcal{W}_{p}\right)\subseteq M$
is a $C^{\infty}$-diffeomorphism for each $p\in M$.
Let $\text{Diff}^{\text{ }r}(M)$ be the set of all $C^{r}$-diffeomorphism $\phi:M\to M$. For each $\phi\in\text{Diff}^{\text{ }r}(M)$ we define the Banach space of $C^{r}$-sections
$\Gamma_{C^{r}}(\phi):=\{X\in C^{r}(M,TM):\pi_{TM}\circ X=\phi\}$
and the open 00-neighborhood
$\mathcal{V}_{\phi}:=\{X\in\Gamma_{C^{r}}(\phi):X(M)\subseteq\mathcal{W}\}.$
If
$\mathcal{U}_{\phi}:=\{\exp\circ X\in\text{Diff}^{\text{ }r}(M):X\in\mathcal{V}_{\phi}\}$
then the map
$\Psi_{\phi}:\mathcal{U}_{\phi}\to\mathcal{V}_{\phi},\quad\Psi(\phi)(p):=\left(\exp_{p}|_{\mathcal{W}_{p}}\right)^{-1}(\phi(p))$
define a chart for $\text{Diff}^{\text{ }r}(M)$ on $\phi$, with inverse given by
$\Psi_{\phi}^{-1}:\mathcal{V}_{\phi}\to\mathcal{U}_{\phi},\quad X\mapsto\exp\circ X,$
turning $\text{Diff}^{\text{ }r}(M)$ in a Banach manifold. Moreover, under the composition the manifold $\text{Diff}^{\text{ }r}(M)$ becomes a right half-Lie group (see e.g. [12]). We see that for the neutral element $e=\text{id}_{M}$, the tangent space $T_{e}\text{Diff}^{\text{ }r}(M)$ coincide with space of $C^{r}$-vector fields $X:M\to TM$, denoted by $\mathcal{X}^{r}(M)$.
The restricted right action
$\mathcal{X}^{r}(M)\times\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})\to T\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n}),\quad(X,\phi)\mapsto X.\phi=X\circ\phi$
is continuous (see e.g. [11]).
Let $[\gamma]\in L^{p}([0,1],\mathcal{X}^{r}(M))$, we will study the equation
$\dot{\eta}=\gamma.\eta,\quad\eta(0)=\text{id}_{M}.$
Definition 7.1
Let $(E,\lVert\cdot\lVert_{E})$ be a normed space and $U\subseteq E$ be a subset. We say that a function $f:[a,b]\times U\to E$ satisfies a $L^{1}$-Lipschitz condition if there exists a measurable function $g\in L^{1}([a,b],\mathbb{R})$ such that
$\text{Lip}(f(t,\cdot))\leq g(t),\quad t\in[a,b]$
Where $\text{Lip}(f(t,\cdot))\in[0,\infty]$ denote the infimum of all Lipschitz constant for $f(t,\cdot):U\to E$.
Definition 7.2
Let $M$ be a $C^{1}$-manifold modeled on a normed space $E$, $J\subseteq\mathbb{R}$ be a non-degenerate interval and $f:J\times M\to TM$ be a function with $f(t,p)\in T_{p}M$ for all $(t,p)\in J\times M$. We say that $f$ satisfies a local $L^{1}$-Lipschitz condition if for all $t_{0}\in J$ and $P\in M$, there exists a chart $\kappa:U_{\kappa}\to V_{\kappa}$ of $M$ on $p$ and a relatively open subinterval $[a,b]\subseteq J$ which is a neighborhood of $t_{0}$ in $J$ such that the map
$f_{\kappa}:[a,b]\times V_{\kappa}\to E,\quad(t,x)\mapsto d\kappa\left(f(t,\kappa^{-1}(y))\right)$
satisfies an $L^{1}$-Lipschitz condition.
Remark 7.3
Let $(E,\lVert\cdot\lVert)$ be a normed space and $f:[a,b]\times M\to TM$ be a map with $f(t,q)\in T_{q}M$ for all $(t,q)\in[a,b]\times M$, which satisfies a local $L^{1}$-condition. If $\tau,\eta:[a,b]\to M$ are two $AC_{L^{1}}$-Carathéodory solutions to
$y^{\prime}=f(t,y)$
satisfying $\tau(t_{0})=\eta(t_{0})$ for some $t_{0}\in[a,b]$, then $\tau=\eta$ [10, Proposition 10.5]. If the initial value problem
$y^{\prime}(t)$ $=f(t,y(t)),\quad t\in[a,b]$
$y(0)$ $=q$
has a (necessarily unique) solution $\gamma_{q}:[a,b]\to M$ for each $q\in M$, then we say that $f$ admits a global flow for initial time $t_{0}$ and write
$F_{t,a}^{f}(q):=\gamma_{q}(t),$
for all $t\in[a,b]$ and $q\in M$.
The following result can be found in the appendix of [10] for the context of $AC_{L^{1}}$-functions with values in the Lie group of $C^{\infty}$-diffeomorphism with compact support $\text{Diff}_{c}^{\infty}(M)$. However, this result is also valid in our context.
Theorem 7.4
Let $1\leq p<\infty$ and $[\gamma]\in L^{p}([0,1],\mathcal{X}^{r}(M))$, then map
$\overline{\gamma}:[0,1]\times M\to TM,\quad\overline{\gamma}(t,q):=\gamma(t)(q)$
satisfies a local $L^{1}$-Lipschitz condition. Let $\eta\in AC_{L^{p}}([0,1],\text{Diff}^{\text{ }r}(M))$ with $\eta(0)=\text{id}_{M}$. Then $\eta=\operatorname{Evol}([\gamma])$ if and only if $f$ admits global flow $F$ for initial time $t_{0}=0$ and
$\eta(t)(p)=F_{t,0}^{\overline{\gamma}}(q),$
for all $t\in[0,1]$ and $q\in M$.
The following result follows the same steps as the proof of the $L^{1}$-regularity of $\text{Diff}_{c}^{\infty}(M)$ proved in [10, Section 11]. We recommend to the reader to read the reference to see some steps in detail.
Theorem 7.5
Let $M$ be a compact smooth manifold and $1\leq p<\infty$. If $r\in\mathbb{N}$, then the right half-Lie group $\text{Diff}^{\text{ }r}(M)$ is $L^{p}$-semiregular. Moreover, the evolution map
$\operatorname{Evol}:L^{p}([0,1],T_{e}\text{Diff}^{\text{ }r}(M))\to AC_{L^{p}}([0,1],\text{Diff}^{\text{ }r}(M)),\quad[\gamma]\mapsto\eta_{[\gamma]}$
is continuous.
Proof.
For $\phi\in\text{Diff}^{\text{ }r}(M)$, we can choose a finite locally finite cover $(U_{i})_{i=1}^{n}$ of $M$ by relatively compact, open subsets sets and charts $\kappa_{i}:U_{i}\to B_{5}(0)\subseteq\mathbb{R}^{n}$ such that the family of open sets $\kappa_{i}^{-1}(B_{1}(0))$ cover $M$ and $U_{i}\subseteq\phi^{-1}(U_{\psi_{i}})$ for some chart $\psi_{i}:U_{\psi_{i}}\to V_{\psi_{i}}$ of $M$ (see e.g. [17]). For $X\in\Gamma_{C^{r}}(\phi)$ we write
$X_{i}:=d\psi_{i}\circ X\circ\kappa_{i}^{-1}:B_{5}(0)\to\mathbb{R}^{n}.$
Then, for each $\ell\in[1,5]$ the map
$\rho_{\ell}:\Gamma_{C^{r}}(\phi)\to\prod_{i=1}^{n}C^{r}(B_{\ell}(0),\mathbb{R}^{n}),\quad X\mapsto\rho_{\ell}(X):=(X_{i}|_{B_{\ell}(0)})_{i=1}^{n}$
is a linear topological embedding with closed image (see [2]).
Let $\phi=\text{id}_{M}$, then $\Gamma_{C^{r}}(\text{id}_{M})=\mathcal{X}^{r}(M)$. Doing the corresponding identifications of product spaces, we have the linear topological embeddings with closed image
$R_{5}:=L^{p}([0,1],\rho_{5}):L^{p}([0,1],\mathcal{X}^{r}(M))\to\prod_{i=1}^{n}L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{n})),\quad\gamma\mapsto\rho_{5}\circ\gamma$
and
$R_{1}:=C([0,1],\rho_{1}):C([0,1],\mathcal{X}^{r}(M))\to\prod_{i=1}^{n}C([0,1],C^{r}(B_{1}(0),\mathbb{R}^{n})),\quad\gamma\mapsto\rho_{1}\circ\gamma.$
Before study the equation for $[\gamma]\in L^{p}([0,1],\mathcal{X}^{r}(M))$, the embedding $R_{1}$ allow us to study it first for
$[\gamma]\in L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{n})).$
We consider the continuous seminorm $q:C^{r}(B_{5}(0),\mathbb{R}^{n})\to[0,\infty)$ by
$q(\phi):=\sup_{x\in\overline{B}_{4}(0)}\left(\lVert\phi^{\prime}(x)\lVert_{op}+\lVert\phi(x)\lVert_{\infty}\right),\quad\text{for all }\phi\in C^{r}(B_{5}(0),\mathbb{R}^{n}).$
For $10<L<$ 1 we set the open ball
$B_{L}:=\{\gamma\in L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{n})):\lVert\gamma\lVert_{L^{p},q}<L\}.$
Just like the case for $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$, we need a $AC_{L^{p}}$-Carathéodory solution $\zeta:[0,1]\to B_{4}(0)$ that verifies the integral equation
[…]
for $[\gamma]\in B_{L}$ and $x\in B_{3}(0)$.
We define the operator $T:B_{L}\times B_{3}(0)\times C([0,1],B_{4}(0))\to C([0,1],B_{4}(0))$ by
[…]
By Lemma 5.12, the map $T$ is $C^{\infty,\infty,r}$ and for each $[\gamma]\in B_{L}$ and $x\in B_{3}(0)$ we have
$\text{Lip}(\Psi([\gamma],x,\cdot))\leq L.$
Thus, by Banach’s Fixed Point Theorem, for each $([\gamma],x)\in B_{L}\times B_{3}(0)$ the map $\Psi([\gamma],x,\cdot)$ has a unique fixed point $\zeta_{[\gamma],x}\in C([0,1],B_{4}(0))$ and
$F:B_{L}\times B_{3}(0)\to C([0,1],B_{4}(0)),\quad([\gamma],x)\mapsto\zeta_{[\gamma],x}$
is $C^{r}$ (see [10, Lemma 6.2]). By the exponential law, we have
  • a)
    $F_{2}:\left(B_{L}\times B_{3}(0)\right)\times[0,1]\to B_{4}(0),\quad([\gamma],t,x)\mapsto\zeta_{[\gamma],x}(t)$ is $C^{r,0}$.
  • b)
    $F_{3}:B_{L}\times[0,1]\to C^{r}(B_{3}(0),B_{4}(0)),\quad([\gamma],t)\mapsto\zeta_{[\gamma]}(t)$ is $C^{r,0}$.
  • c)
    $F_{4}:B_{L}\to C([0,1],C^{r}(B_{3}(0),B_{4}(0))),\quad[\gamma]\mapsto\zeta_{[\gamma]}$ is continuous.
Let $\rho_{2}:C^{r}(B_{3}(0),B_{4}(0))\to C^{r}(B_{2}(0),B_{4}(0))$, $\varphi\mapsto\varphi|_{B_{2}(0)}$, we denote
$H:=C([0,1],\rho_{2})\circ F_{4}:B_{L}\to C([0,1],C^{r}(B_{2}(0),B_{4}(0)),\quad[\gamma]\mapsto\rho_{2}\circ\zeta_{[\gamma]}.$
Then $H$ is continuous. We will show that $H$ has absolutely continuous values.
Let consider the open set
$\lfloor\overline{B_{2}(0)},B_{4}(0)\rfloor_{r}:=\{\varphi\in C^{r}(B_{3}(0),\mathbb{R}^{n}):\varphi(\overline{B_{2}(0)})\subseteq B_{4}(0)\}.$
Then the map
$S:C^{r}(B_{5}(0),\mathbb{R}^{n})\times\lfloor\overline{B_{2}(0)},B_{4}(0)\rfloor_{r}\to C^{r}(B_{2}(0),\mathbb{R}^{n}),\quad(\psi,\varphi)\mapsto\psi\circ\varphi$
is continuous. Moreover, since $S$ is linear in the first variable, by Lemma 5.12, the map
$\tilde{S}:L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{n}))\times C([0,1],\lfloor\overline{B_{2}(0)},B_{4}(0)\rfloor_{r})\to L^{p}([0,1],C^{r}(B_{2}(0),\mathbb{R}^{n})),\quad(\alpha,\beta)\mapsto S\circ(\alpha,\beta)$
is continuous. Hence, for $[\gamma]\in B_{L}$ we have
$S\circ(\gamma,H([\gamma]))=\gamma\left(H(\gamma)\right)=\gamma.H([\gamma])\in L^{p}([0,1],C^{r}(B_{2}(0),\mathbb{R}^{n})).$
This allow us to define the $L^{p}$-absolute continuous function $\tau:[0,1]\to C^{r}(B_{2}(0),\mathbb{R}^{n})$ given by
[…]
If we consider the point evaluation map $\varepsilon_{x}:C^{r}(B_{2}(0),\mathbb{R}^{n})\to\mathbb{R}^{n}$, $\varphi\mapsto\varphi(x)$, then for each $x\in B_{2}(0)$ we have
$\varepsilon_{x}\left(\tau(t)\right)=\varepsilon_{x}\left(H([\gamma])(t)\right),\quad\text{for all }t\in[0,1].$
Since the family of maps $(\varepsilon_{x})_{x\in B_{2}(0)}$ separate points in $C^{r}(B_{2}(0),\mathbb{R}^{n})$, we have $\tau=H([\gamma])$.
Hence $H[\gamma]$ is absolutely continuous and we can consider the map
$H_{AC}:B_{L}\to AC_{L^{p}}([0,1],C^{r}(B_{2}(0),B_{4}(0)),\quad[\gamma]\mapsto H([\gamma])$
which is the evolution map for $C^{r}(B_{2}(0),B_{4}(0))$.
Now we will see the case for $L^{p}([0,1],\mathcal{X}^{r}(M))$. We define the 00-neighborhood of $L^{p}([0,1],\mathcal{X}^{r}(M))$
$\mathcal{B}_{L}$ $:=R_{5}^{-1}\left(\prod_{i=1}^{n}B_{L}\right)$
=$=\{[\gamma]\in L^{p}([0,1],\mathcal{X}^{r}(M)):(\forall i\in\{1,...,n\})\text{ }[\gamma_{i}]\in B_{L}\}$
which is open by continuity of $R_{5}$.
Let $[\gamma]\in\mathcal{B}_{L}$, with $[\gamma_{i}]\in B_{L}$ for each $i\in\{1,...,n\}$. If $p\in M$, there is a chart $\kappa_{j}:U_{j}\to B_{5}(0)$ such that $p\in\kappa_{j}^{-1}(B_{3}(0))$ for some $j\in\{1,...,n\}$. Let $\zeta_{[\gamma],\kappa_{j}(p)}\in AC_{L^{P}}([0,1],C^{r}(B_{2}(0),B_{4}(0))$ be the solution of the integral equation
[…]
Then $\zeta_{[\gamma],\kappa_{j}(p)}$ is a $AC_{L^{p}}$-Carathéodory solution to
$x^{\prime}(t)$ $=\gamma_{j}(t)\left(x(t)\right),\quad t\in[0,1]$
$x(0)$ $=\kappa_{j}(p).$
Hence the function
$\eta_{[\gamma],p}:[0,1]\to M,\quad\eta_{[\gamma],p}(t):=\kappa_{j}^{-1}\circ\zeta_{[\gamma],\kappa_{j}(p)}(t)$
is a $AC_{L^{p}}$-Carathéodory solution of the equation
$y^{\prime}(t)$ $=\gamma(t)\left(y(t)\right),\quad t\in[0,1]$
$y(0)$ $=p.$
Since $\tilde{\gamma}$ satisfies a local $L^{1}$-Lipschitz condition, by Remark 7.3 the solution of this equation is unique and $\eta_{[\gamma],p}$ is well defined. Moreover, $\tilde{\gamma}$ admits a global flow for initial time $t_{0}=0$, given by
$F_{t,0}^{\overline{\gamma}}(p)=\eta_{[\gamma],p}(t)$
for all $t\in[0,1]$, $p\in M$.
Following point 11.16 of [10], for each $i\in\{1,...,n\}$ we can construct an exponential map
$\text{exp}_{i}:\mathcal{D}_{i}\subseteq TB_{5}(0)\to B_{5}(0)$
such that
$\kappa_{i}^{-1}\circ\text{exp}_{i}=\text{exp}\circ T\kappa_{i}^{-1}|_{\mathcal{D}_{i}}.$
Moreover, for an open subset $\mathcal{O}_{i}\subseteq\mathcal{D}_{i}$ we consider the map
$\psi_{i}:=(\text{pr}_{1},\text{exp}_{i})|_{\mathcal{O}_{i}}:\mathcal{O}_{i}\subseteq TB_{5}(0)\to\psi_{i}(\mathcal{O}_{i})\subseteq B_{5}(0)\times B_{4}(0).$
Assuming that
$T\kappa_{i}^{-1}(\mathcal{O}_{i})\subseteq\mathcal{W},$
since
$\{(x,x):x\in\overline{B_{5}(0)}\}\subseteq(\text{pr}_{1},\exp_{i})(\mathcal{O}_{i}),$
there is exists $s_{i}\in]0,1]$ such that
$\bigcup_{x\in B_{4}(0)}\{x\}\times B_{s_{i}}(x)\subseteq(\text{pr}_{1},\exp_{i})(\mathcal{O}_{i})$
is a smooth diffeomorphism and there exists a $s_{i}\in]0,1[$, such that $\psi_{i}^{-1}$ restrict to a smooth diffeomorphism of the form
$(\text{Id}_{B_{4}(0)},\theta_{i}):\bigcup_{x\in B_{4}(0)}\{x\}\times B_{s_{i}}(x)\subseteq B_{5}(0)\times B_{5}(0)\to\psi_{i}^{-1}\left(\bigcup_{x\in B_{4}(0)}\{x\}\times B_{s_{i}}(x)\right)\subseteq\mathcal{O}_{i}$
with open image in $\mathcal{O}_{i}$. For each $x\in B_{4}(0)$ we define the set
$\mathcal{O}_{i,x}:=\{y\in\mathbb{R}^{n}:(x,y)\in\mathcal{O}_{i}\}.$
Hence, we have
$\theta_{i}(x,\cdot)=\left(\exp(x,\cdot)|_{\mathcal{O}_{i,x}}\right)^{-1}|_{B_{s_{i}}(0)}.$
The set
$Z_{i,x}:=\left\{\varphi\in C^{r}(B_{2}(0),\mathbb{R}^{n}):(\forall x\in\overline{B_{1}(0)})\text{ }\varphi(x)\in B_{s_{i}}(x)\right\}$
is open in $C^{r}(B_{2}(0),\mathbb{R}^{n})$ and the map
$(\theta_{i})_{*}:Z_{i,x}\subseteq C^{r}(B_{2}(0),\mathbb{R}^{n})\to C^{r}(B_{1}(0),\mathbb{R}^{n}),\quad\varphi\mapsto(\theta_{i})_{*}(\varphi):=\theta_{i}\circ(\text{id}_{B_{1}(0)},\varphi)$
is smooth since $\theta$ is smooth [12, Proposition 4.23].
Let $\Psi:\mathcal{U}\to\mathcal{V}$ be a chart on $\text{id}_{M}\in\text{Diff}^{r}(M)$, then since $\rho_{1}$ is a topological embedding we can assume that
$\mathcal{V}:=\rho_{1}^{-1}\left(\prod_{i=1}^{n}\mathcal{V}_{i}\right)$
for suitable open 00-neighborhoods $\mathcal{V}_{i}\subseteq C^{r}(B_{1}(0),\mathbb{R}^{n})$.
Since $(\theta_{i})_{*}(\text{id}_{B_{2}(0)})=0$, by continuity of $(\theta_{i})_{*}$ there exists open $\text{id}_{B_{2}(0)}$-neighborhoods $\mathcal{Y}_{i}\subseteq Z_{i,x}$ such that
$(\theta_{i})_{*}(\mathcal{Y}_{i})\subseteq\mathcal{V}_{i}.$
Since $H(0)(t)=\text{id}_{B_{2}(0)}$ for all $t\in[0,1]$, by continuity of $H$ there exists open 00-neighborhoods $\mathcal{P}_{i}\subseteq\mathcal{B}_{L}$ such that
$H(\mathcal{P}_{i})\subseteq C([0,1],\mathcal{Y}_{i})$
If
$\mathcal{P}:=(R_{\text{id}_{M},t})^{-1}\left(\prod_{i=1}^{n}\mathcal{P}_{i}\right)$
then $\mathcal{P}$ is an open 00-neighborhood in $L^{p}([0,1],\mathcal{X}^{r}(M))$ with $\mathcal{P}\subseteq\mathcal{B}_{L}$.
This allow us to do the composition
$(\theta_{i})_{*}\circ H:B_{L}\subseteq L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{n})\to C([0,1],C^{r}(B_{1}(0),\mathbb{R}^{n})$
Following the same steps of point $1.18$ and $1.19$ of [10], we have that for each $[\gamma]\in\mathcal{P}$, there exists an unique $\theta_{[\gamma]}\in C([0,1],\mathcal{X}^{r}(M))$ such that
$R_{1}(\theta_{[\gamma]})=\Big{(}(\theta_{i})_{*}(H([\gamma_{i}]))\Big{)}_{i=1}^{n}.$
Then $\rho_{1}\left(\theta_{[\gamma]}(t)\right)\in\prod_{i=1}^{n}\mathcal{V}_{i}$, whence $\theta_{[\gamma]}(t)\in\mathcal{V}$. In consequence
$\Psi^{-1}(\theta_{[\gamma]}(t))=\exp\circ\theta_{[\gamma]}(t)\in\text{Diff}^{\text{ }r}(M),\quad\text{ for all }t\in[0,1].$
Let $p\in M$ with $x:=\kappa_{i}(p)$, then we have
$\exp\circ\theta_{[\gamma]}(t)(p)$ $=\left(\kappa_{i}^{-1}\circ\exp_{i}\circ T\kappa_{i}\right)\circ\theta_{[\gamma]}(t)(\kappa_{i}^{-1}(x))$
$=\kappa_{i}^{-1}\circ\exp_{i}\circ(\theta_{i})_{*}(H([\gamma_{i}]))(t)(x)$
$=\kappa_{i}^{-1}\circ\exp_{i}\circ\left(\exp_{i}|_{\mathcal{W_{i}}}\right)^{-1}\circ H([\gamma_{i}])(t)(x)$
$=\kappa_{i}^{-1}\circ H([\gamma_{i}])(t)(x)$
$=\kappa_{i}^{-1}\circ\zeta_{[\gamma_{i}],x}(t)$
$=F_{t,0}^{\tilde{\gamma}}(p).$
Hence, we define
$\eta_{[\gamma]}:=\exp\circ\theta_{[\gamma]}:[0,1]\to\text{Diff}^{\text{ }r}(M).$
We will show that $\eta_{[\gamma]}$ is absolutely continuous. Since the exponential map $\exp$ is smooth, we need to show that $\theta_{[\gamma]}$ is absolutely continuous, but this is equivalent to show that
$[0,1]\to\prod_{i=1}^{n}C^{r}(B_{1}(0),\mathbb{R}^{n}),\quad t\mapsto\rho_{1}\circ\theta_{[\gamma]}(t)$
But this maps coincide with the map
$[0,1]\to\prod_{i=1}^{n}C^{r}(B_{1}(0),\mathbb{R}^{n}),\quad t\mapsto\left((\theta_{i})_{*}\left(H([\gamma_{i}])\right)\right)_{i=1}^{n}(t).$
is absolutely continuous. Since $(\theta_{i})_{*}$ is smooth and $H([\gamma_{i}])\in AC_{L^{p}}([0,1],C^{r}(B_{2}(0),B_{4}(0)))$, by Lemma 2.16, the composition and hence the product are absolutely continuous. Then we have that
$\theta_{[\gamma]}\in AC_{L^{p}}([0,1],\mathcal{X}^{r}(M))$
whence
$\eta_{[\gamma]}\in AC_{L^{p}}([0,1],\text{Diff}^{\text{ }r}(M)).$
In consequence, Lemma 5.9 and Remark 7.3, the right half-Lie group $\text{Diff}^{\text{ }r}(M)$ is $L^{p}$-semiregular. Since the map
$B_{L}\to\prod_{i=1}^{n}C([0,1],C^{r}(B_{1}(0),\mathbb{R}^{n})),\quad[\gamma_{i}]\mapsto\left((\theta_{i})_{*}\left(H([\gamma_{i}])\right)\right)_{i=1}^{n}$
is continuous, the restricted evolution map with continuous values is given by
$\operatorname{Evol}_{C}|_{\mathcal{P}}:\mathcal{P}\to AC_{L^{p}}([0,1],\text{Diff}^{\text{ }r}(M)),\quad[\gamma]\mapsto\eta_{[\gamma]}=\exp\circ\theta_{[\gamma]}.$
Hence, by Lemma 5.11 the evolution map $\operatorname{Evol}_{C}$ is continuous and since the restricted right action of $\text{Diff}^{\text{ }r}(M)$ is continuous, by Theorem 5.14, the evolution map
$\operatorname{Evol}:L^{p}([0,1],\mathcal{X}^{r}(M))\to AC_{L^{p}}([0,1],\text{Diff}^{\text{ }r}(M)),\quad[\gamma]\mapsto\operatorname{Evol}(\gamma)$
is continuous. ∎
Acknowledgements: The author would like to thank Helge Glöckner for his guidance during the development of this work. The author was financially supported by ANID and DAAD (DAAD/BecasChile 2020, ID:91762237/62190017).

### v4

@@
 Smoothness of natural mappings between spaces of absolutely continuous functions is discussed, like superposition operators $AC_{L^{p}}([a,b],N_{1})\to AC_{L^{p}}([a,b],N_{2})$, $\eta\mapsto f\circ\eta$, for a smooth map $f\colon N_{1}\to N_{2}$.
-For $1\leq p<\infty$ and $r\in\mathbb{N}$ we show that the right half-Lie groups $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$ and $\text{Diff}^{\text{ }r}(M)$ are $L^{p}$-semiregular.
-Here $K$ is a compact subset of $\mathbb{R}^{n}$ and $M$ is a compact smooth manifold.
-An $L^{p}$-semiregular half-Lie group $G$ admits an evolution map $\operatorname{Evol}:L^{p}([0,1],T_{e}G)\to AC_{L^{p}}([0,1],G)$, where $e$ is the neutral element of $G$.
-For the preceding examples, the evolution map $\operatorname{Evol}$ is continuous.
+Let $G$ be a right half-Lie group.
+For $1\leq p<\infty$, we show certains properties of $L^{p}$-semiregularity of $G$ under certains condicions on $G$.
+For $r\in\mathbb{N}$, we prove that the right half-Lie groups $\text{Diff}_{K}^{r}(\mathbb{R})$ and $\text{Diff}^{r}(M)$ are $L^{p}$-semiregular.
+Here $K$ is a compact subset of $\mathbb{R}$ and $M$ is a compact smooth manifold (possible with boundary).
+An $L^{p}$-semiregular half-Lie group $G$ admits an evolution map $\text{Evol}:L^{p}([0,1],T_{e}G)\to AC_{L^{p}}([0,1],G)$, where $e$ is the neutral element of $G$.
+For the preceding examples, the evolution map Evol is continuous.
 Manifolds of absolutely continuous functions with values in
@@
 1 Introduction
-Manifolds of mappings play an important role in various branches of global analysis, infinite-dimensional geometry and Lie theory.
-For a compact smooth manifold $M$, a finite-dimensional manifold $N$ and a non-negative integer $r$, a smooth manifold structure on the set $C^{r}(M,N)$ of $C^{r}$-functions $\phi\colon M\to N$ was first obtained by Eells (see [6] and the references therein).
-Later, manifold structures were also obtained on $C^{\infty}(M,N)$, and for infinite-dimensional manifolds $N$ as long as they admit a local addition, a concept recalled in Definition 1.1 (see [19, 15] and the references therein, also [2]).
-For a Hilbert manifold $M$, a smooth manifold structure on the set $AC_{L^{2}}([0,1],N)$ of absolutely continuous paths with $L^{2}$-derivatives in local charts was used by Flaschel and Klingenberg to study closed geodesics in Riemannian manifolds (see [7] and [14], cf.
+The study of mapping manifolds is an important branch of geometry and global analysis.
+For example, for $k\in\mathbb{N}\cup\{0\}$, a well-known case is the manifold of the set of $C^{k}$-maps $f:M\to N$, denoted by $C^{k}(M,N)$, for a compact manifold $M$ and a smooth manifold $N$ admitting local addition (see e.g.
+[1, 9, 28]).
+In [24], Krikorian shows how to contruct smooth manifold structures for set of functions with Hölder conditions and Sobolev functions.
+Our main focus will be the study of absolutely continuous functions with values in a manifold $N$ modeled on a sequentially complete locally convex space which admits a local addition.
+For a Hilbert manifold $N$, a smooth manifold structure on the set $AC_{L^{2}}([0,1],N)$ of absolutely continuous paths with $L^{2}$-derivatives in local charts was used by Flaschel and Klingenberg to study closed geodesics in Riemannian manifolds (see [10] and [22], cf.
 also [39]).
-For Banach manifolds $N$ admitting a smooth local addition and $p\in[1,\infty]$, a smooth manifold structure on $AC_{L^{p}}([0,1],N)$ may also be obtained using a method of Krikoian [16] which is similar to Palais’ use of Banach section functors [25].
+For Banach manifolds $N$ admitting a smooth local addition and $p\in[1,\infty]$, Pierron and Trouvé show that a smooth manifold structure on $AC_{L^{p}}([0,1],N)$ may also be obtained using a method of Krikoian [24], which is similar to Palais’ use of Banach section functors [35].
 A Lie group structure (and hence a smooth manifold structure) on $AC_{L^{p}}([0,1],G)$ was obtained in [33] for each Lie group $G$ modeled on a sequentially complete locally convex space (as in [30]), generalizing the case of Fréchet–Lie groups treated in [14].
@@
 $\Sigma\colon\Omega\to N,$
-defined on an open set $\Omega\subseteq TN$ which contains the zero-vector $0_{p}\in T_{p}N$ for each $x\in N$, such that $\Sigma(0_{p})=p$ for all $p\in N$ and
+defined on an open set $\Omega\subseteq TN$ which contains the zero-vector $0_{p}\in T_{p}N$ for each $p\in N$, such that $\Sigma(0_{p})=p$ for all $p\in N$ and
 $\theta\colon\Omega\to N\times N,\quad v\mapsto(\pi_{TN}(v),\Sigma(v))$
-has open image and is a $C^{\infty}$-diffeomorphism onto its image $\Omega^{\prime}$.
+has open image and is a $C^{\infty}$-diffeomorphism onto its image.
 Throughout the following, $a<b$ are real numbers and $p\in[1,\infty]$.
-For a definition of absolutely continuous functions with values in a sequentially complete locally convex space $E$ or a smooth manifold $N$ defined thereon, the reader is referred to Definitions 2.11 and 2.21, respectively.
+For a definition of absolutely continuous functions with values in a sequentially complete locally convex space $E$ or a smooth manifold $N$ defined thereon, the reader is referred to Definitions 2.12 and 2.23, respectively (see also [33]).
 For $\eta\in AC_{L^{p}}([a,b],N)$, the pointwise operations make
@@
 a vector space; we endow it with a natural topology making it a locally convex topological vector space (cf. Definition 2.26). We shall see that
-${\mathcal{V}}_{\eta}\colon:=\{\tau\in\Gamma_{AC}(\eta)\colon\tau([a,b])\subseteq\Omega\}$
+${\mathcal{V}}_{\eta}:=\{\tau\in\Gamma_{AC}(\eta)\colon\tau([a,b])\subseteq\Omega\}$
 is an open 00-neighborhood in $\Gamma_{AC}(\eta)$. Setting
-${\mathcal{U}}_{\eta}:=\{\gamma\in AC_{L^{p}}([a,b],N)\colon(t,\gamma(t))\in\Omega^{\prime}\;\mbox{for all $t\in[a,b]$}\},$
+${\mathcal{U}}_{\eta}:=\{\gamma\in AC_{L^{p}}([a,b],N)\colon(\eta(t),\gamma(t))\in\Omega^{\prime}\;\mbox{for all $t\in[a,b]$}\},$
 the map
@@
 More generally, $AC_{L^{p}}([a,b],f)$ is $C^{r}$ for $r\in\mathbb{N}\cup\{0,\infty\}$ whenever $f$ is $C^{r+2}$ (see Proposition 3.8).
-If, in the situation of Definition 1.1, $N$ is $\mathbb{K}$-analytic manifold modeled on a locally convex topological $\mathbb{K}$-vector space over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and $\theta\colon\Omega\to\Omega^{\prime}$ is a diffeomorphism of $\mathbb{K}$-analytic manifolds, then $\Sigma$ is called a $\mathbb{K}$-analytic local addition.
-In this case, $AC_{L^{p}}([a,b],N)$ can be given a $\mathbb{K}$-analytic manifold structure modeled on the locally convex topological $\mathbb{K}$-vector spaces $\Gamma_{\eta}$ with properties as in Theorem 1.2, replacing $C^{\infty}$-diffeomorphisms with diffeomorphisms of $\mathbb{K}$-analytic manifolds there (see Corollary 4.6).
+If, in the situation of Definition 1.1, $N$ is a $\mathbb{K}$-analytic manifold modeled on a locally convex topological $\mathbb{K}$-vector space over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and $\theta\colon\Omega\to\Omega^{\prime}$ is a diffeomorphism of $\mathbb{K}$-analytic manifolds, then $\Sigma$ is called a $\mathbb{K}$-analytic local addition.
+In this case, $AC_{L^{p}}([a,b],N)$ can be given a $\mathbb{K}$-analytic manifold structure modeled on the locally convex topological $\mathbb{K}$-vector spaces $\Gamma_{AC}(\eta)$ with properties as in Theorem 1.2, replacing $C^{\infty}$-diffeomorphisms with diffeomorphisms of $\mathbb{K}$-analytic manifolds there (see Corollary 3.6).
 If $N_{j}$ is a $\mathbb{K}$-analytic manifold modeled on a sequentially complete locally convex topological $\mathbb{K}$-vector space such that $N_{j}$ admits a $\mathbb{K}$-analytic local addition for $j\in\{1,2\}$, then the map $AC_{L^{p}}([a,b],f)$ described in Theorem 1.3 is $\mathbb{K}$-analytic for all $p\in[1,\infty]$ (see Corollary 3.9).
+If $\pi_{TN}:TN\to N$ denotes the tangent bundle of $N$. Followings results on more general manifolds of mappings (see e.g. [3]), we can conclude that the map
+$AC_{L^{p}}([a,b],\pi_{TN}):AC_{L^{p}}([a,b],TN)\to AC_{L^{p}}([a,b],N),\quad\tau\mapsto\pi_{TN}\circ\tau$
+defines a vector bundle on $AC_{L^{p}}([a,b],N)$. The proofs are analogous to the reference given, but we will show them for the sake of completeness.
 Manifolds of absolutely continuous paths occur in the regularity theory of infinite-dimensional Lie groups.
@@
 $TG\times G\to TG,\quad(v,g)\mapsto T\rho_{g}(v)=:v.g$
-is a smooth map and a right action of $G$ in $TG$. The following concept strengthens “regularity” in the sense of Milnor [21].
+is a smooth map and a right action of $G$ on $TG$. The following concept strengthens “regularity” in the sense of Milnor [30].
 Definition 1.4
 Following [33] (cf. also [14]), $G$ is called $L^{p}$-semiregular if, for each $\gamma\in{\mathcal{L}}^{p}([0,1],\mathfrak{g})$, the initial value problem
-$\dot{\eta}(t)=\gamma(t).\eta(t),\quad\eta(0)=e$ (1)
+$\dot{\eta}(t)$ $=\gamma(t).\eta(t),\quad t\in[0,1]$ (1.1)
+$\eta(0)$ $=e$ (1.2)
 has an $AC_{L^{p}}$-solution $\eta\colon[0,1]\to G$.
@@
 Remark 1.5
-If $G$ is modeled on a Fréchet space, (1) is required to hold for almost all $t\in[0,1]$ with respect to Lebesgue measure.
-In the general case, $\eta$ is required to be a Carathéodory solution to (1), i.e., it solves the corresponding integral equation piecewise in local charts.
+If $G$ is modeled on a Fréchet space, (1.1) is required to hold for almost all $t\in[0,1]$ with respect to Lebesgue measure.
+In the general case, $\eta$ is required to be a Carathéodory solution to (1.1), i.e., it solves the corresponding integral equation piecewise in local charts.
 We mention that a Lie group $G$ is $L^{p}$-regular if and only if $G$ is $L^{p}$-semiregular and $\operatorname{Evol}$ is smooth as map
 $L^{p}([0,1],\mathfrak{g})\to C([0,1],G)$
-(see [23]). The latter holds whenever $\operatorname{Evol}\colon L^{p}([0,1],\mathfrak{g})\to C([0,1],G)$ is continuous at 00 (see [11, Theorem E]).
-Now considetr a half-Lie group $G$ modeled on a sequentially complete locally convex space $E$.
+(see [33]). The latter holds whenever $\operatorname{Evol}\colon L^{p}([0,1],\mathfrak{g})\to C([0,1],G)$ is continuous at 00 (see [15, Theorem E]).
+Now consider a half-Lie group $G$ modeled on a sequentially complete locally convex space $E$.
 Thus $G$ is a group, endowed with a smooth manifold structure modeled on $E$ which makes $G$ a topological group and turns the right translation $\rho_{g}\colon G\to G$ into a smooth mapping for each $g\in G$ (cf.
@@
 We say that a half-Lie group $G$ is $L^{p}$-semiregular if the differential equation
-$\dot{y}(t)=\gamma(t).y(t)$
-satisfies local uniqueness of Carathéodory solutions for each $\gamma\in{\mathcal{L}}^{p}([0,1],\mathfrak{g})$ (in the sense of [13]) and the initial value problem (1) has a Carathéodory solution $\operatorname{Evol}(\gamma):=\eta\colon[0,1]\to G$.
+$\dot{y}(t)=\gamma(t).y(t),\quad t\in[0,1]$ (1.3)
+satisfies local uniqueness of Carathéodory solutions for each $\gamma\in{\mathcal{L}}^{p}([0,1],\mathfrak{g})$ (in the sense of [18]) and the initial value problem (1.1) has a Carathéodory solution $\operatorname{Evol}(\gamma):=\eta\colon[0,1]\to G$.
 The Lie group $\text{Diff}(M)$ of $C^{\infty}$-diffeomorphisms of a compact smooth manifold $M$ without boundary is known to be $L^{1}$-regular, and also the Lie group $\text{Diff}_{K}(\mathbb{R}^{n})$ of all $C^{\infty}$-diffeomorphisms $\phi\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that $\phi(x)=x$ for all $x\in\mathbb{R}^{n}\setminus K$, for each compact subset $K\subseteq\mathbb{R}^{n}$ (see [14]).
-For each positive integer $r$, the following analogues are obtained (see Theorems  7.5 and 6.3):
+For each positive integer $r$, the following analogues are obtained (see Theorems  6.5 and 5.3):
 Theorem 1.7
@@
 is continuous.
-If we replace $L^{p}$ with $L_{rc}^{\infty}$ (see Definition 2.5) the result remains valid.
+If we replace $L^{p}$ with $L_{rc}^{\infty}$ (see Definition 2.5) the preceding theorem remains valid.
 Here $\mathfrak{g}$ is the Banach space of all $C^{r}$-vector fields on $\mathbb{R}^{n}$ which vanish outside $K$.
@@
 Theorem 1.9
-Let $1\leq p<\infty$. Let $G$ be a right half-Lie group modeled in a Fréchet space $E$ which admits local addition. Let $G$ be $L^{p}$-semiregular with continuous evolution map
+Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space space $E$ which admits a local addition and $1\leq p<\infty$.
+Let $G$ be $L^{p}$-semiregular with continuous evolution map
 $\operatorname{Evol}_{C}:L^{p}([0,1],T_{e}G),\to C([0,1],G),\quad\gamma\mapsto\operatorname{Evol}_{C}(\gamma).$
@@
 $\operatorname{Evol}:L^{p}([0,1],T_{e}G),\to AC_{L^{p}}([0,1],G),\quad\gamma\mapsto\operatorname{Evol}(\gamma)$
-is continuous.
-If we replace $L^{p}$ with $L_{rc}^{\infty}$ the result remains valid.
+is continuous. If $G$ is a right half-Lie group modeled in an integral complete locally convex space $E$, then if we replace $L^{p}$ with $L_{rc}^{\infty}$ the result remains valid.
 So far, $C^{0}$-regularity has been investigated for the half-Lie group $\text{Diff}^{r}(M)$ in a suitable sense (see the sketch in [29]).
@@
 $d^{(i,j)}f((x,a),y_{1},...,y_{i},b_{1},...,b_{j}):=(D_{(y_{1},0)}...D_{(y_{i},0)}D_{(0,b_{1})}...D_{(0,b_{j})}f)(x,a)$
-exists for all $i,j\in\mathbb{N}\cup\{0\}$ such that $i\leq r$ and $j\leq s$, and all $y_{1},...,y_{i}\in E_{1}$ and $b_{1},...,b_{j}\in E_{2}$. Moreover, we assume that the mappings
+exist for all $i,j\in\mathbb{N}\cup\{0\}$ such that $i\leq r$ and $j\leq s$, and all $y_{1},...,y_{i}\in E_{1}$ and $b_{1},...,b_{j}\in E_{2}$, and, we assume that the mappings
 $d^{(i,j)}f:U\times V\times E_{1}^{i}\times E_{2}^{j}\to F$
-are continuous. Then the map $f$ is called a $C^{r,s}$-map.
+are continuous, then $f$ is called a $C^{r,s}$-map.
 Definition 2.3
@@
 $\mu:\mathcal{B}(X)\to[0,\infty]$ on its $\sigma$-algebra of Borel sets and let $Y$ be a topological space.
-A function $\gamma:X\to Y$ is called Lusin $\mu$-measurable (or $\mu$-measurable) if for each compact subset $K\subseteq X$ there exists a sequence $(K_{n})_{n\in\mathbb{N}}$ of compact subsets $K_{n}\subseteq K$ such that each restriction $\gamma|_{K_{n}}$ is continuous and $\mu\left(K\backslash\cup_{n\in\mathbb{N}}K_{n}\right)=0$.
+A function $\gamma:X\to Y$ is called Lusin $\mu$-measurable (or $\mu$-measurable) if for each compact subset $K\subseteq X$, there exists a sequence $(K_{n})_{n\in\mathbb{N}}$ of compact subsets $K_{n}\subseteq K$ such that each restriction $\gamma|_{K_{n}}$ is continuous and $\mu\left(K\backslash\cup_{n\in\mathbb{N}}K_{n}\right)=0$.
 For details of the construction of Lebesgue spaces, we refer the reader to see [33].
 Definition 2.4
-Let $E$ be a locally convex space, $a<b$ be real numbers, $1\leq p<\infty$ and $\lambda:\mathcal{B}([a,b])\to[0,\infty)$ be the restriction of the Lebesgue measure on $\mathbb{R}$.
+Let $E$ be a locally convex space, $a<b$ be real numbers, $1\leq p<\infty$ and $\lambda:\mathcal{B}([a,b])\to[0,\infty)$ be the restriction of the Lebesgue-Borel measure on $\mathbb{R}$.
 We define the set $\mathcal{L}^{p}([a,b],E)$ as the set of all Lusin $\lambda$-measurable functions $\gamma:[a,b]\to E$ such that for each continuous seminorm $q$ on $E$ we have
@@
 $\lVert\cdot\lVert_{\mathcal{L}^{p},q}:\mathcal{L}^{p}([a,b],E)\to[0,\infty[,\quad\lVert\gamma\lVert_{\mathcal{L}^{p},q}:=\lVert q\circ\gamma\lVert_{\mathcal{L}^{p}}.$
-Let $\gamma\sim\eta$ if and only if $\gamma=\eta$ a.e. We define the Hausdorff locally convex space
+Let $\gamma\sim\eta$ if and only if $\gamma(t)=\eta(t)$ for almost all $t\in[a,b]$ and write $[\gamma]$ for the equivalence class of $\gamma$. We define the Hausdorff locally convex space
 $L^{p}([a,b],E):=\mathcal{L}^{p}([a,b],E)/[0]$
@@
 For each continuous seminorm $q:E\to\mathbb{R}^{n}$ we define the seminorm
-$\lVert\gamma\lVert_{\mathcal{L}^{q}}:=\operatorname*{ess\,sup}_{t\in[a,b]}q\circ\gamma(t).$
+$\lVert\gamma\lVert_{\mathcal{L}^{\infty},q}:=\operatorname*{ess\,sup}_{t\in[a,b]}q\circ\gamma(t).$
 We endow $\mathcal{L}^{\infty}([a,b],E)$ with the Hausdorff locally convex topology given by these seminorms.
@@
 $\lVert\cdot\lVert_{L^{\infty},q}:L^{\infty}([a,b],E)\to[0,\infty[,\quad\lVert[\gamma]\lVert_{L^{\infty},q}:=\lVert\gamma\lVert_{\mathcal{L}^{\infty},q}.$
-We define the vector space $\mathcal{L}_{rc}^{\infty}([a,b],E)$ of all Borel measurable functions $\gamma:[a,b]\to E$ with compact and metrizable image, endowed with the topology induced by $\mathcal{L}^{\infty}([a,b],E)$.
+We define the vector space $\mathcal{L}_{rc}^{\infty}([a,b],E)$ of all Borel measurable functions $\gamma:[a,b]\to E$ such that the image $\text{Im}(\gamma)$ has compact and metrizable closure, endowed with the topology induced by $\mathcal{L}^{\infty}([a,b],E)$.
 Thus
@@
 Remark 2.6
-Let $E$ be a Frechet space and $p\in[0,1]$.
-If $L_{\mathcal{B}}^{p}([0,1],E)$ denotes the Lebesgue space constructed with the set of Borel measurable functions (see [10]), then $L_{\mathcal{B}}^{p}([0,1],E)$ coincides with $L^{p}([0,1],E)$ ([24, Proposition 2.10]).
+Let $E$ be a Frechet space and $p\in[1,\infty]$.
+If $L_{\mathcal{B}}^{p}([0,1],E)$ denotes the Lebesgue space constructed with the set of Borel measurable functions (see [14]), then $L_{\mathcal{B}}^{p}([0,1],E)$ coincides with $L^{p}([0,1],E)$ ([33, Proposition 2.10]).
 Remark 2.7
@@
 Definition 2.8
-Let $E$ be a locally convex space and ${\gamma:[a,b]\to E}$ be a Lusin $\lambda$-measurable function such that $\alpha\circ\gamma\in\mathcal{L}^{1}([a,b],\mathbb{R})$ for each $\alpha\in E^{\prime}$.
+Let $E$ be a locally convex space and ${\gamma:[a,b]\to E}$ be such that $\alpha\circ\gamma\in\mathcal{L}^{1}([a,b],\mathbb{R})$ for each $\alpha\in E^{\prime}$.
 An element $z\in E$ is called the weak integral of $\gamma$ if
-$\forall\alpha\in E^{\prime}:\quad\alpha(z)=\int_{a}^{b}(\alpha\circ\gamma)(s)ds.$
 […]
-The following lemma [24, Proposition 2.26] allow us to define absolute continuous functions with vector values.
+for each $\forall\alpha\in E^{\prime}$. Then $z$ is called the weak integral of $\gamma$ from $a$ to $b$, and we write $z=:\int_{a}^{b}\gamma(s)ds$.
+Definition 2.9
+Let $E$ be a locally convex space.
+We say that a sequence $(x_{n})_{n}\subset E$ is a Cauchy sequence if for each $\varepsilon>0$ and each continuous seminorm $q$ of $E$, there exists an integer $N_{\varepsilon,q}\in\mathbb{N}$, such that for all $m,n\geq N_{\varepsilon,q}$ we have
+$q(x_{m}-x_{n})<\varepsilon.$
+We say that $E$ is sequentially complete if every Cauchy sequence converge in $E$.
+The following lemma [33, Proposition 2.26] allows us to define absolute continuous functions with vector values.
 Lemma 2.10
@@
 […]
-We denote the space of all $L^{p}$-absolutely continuous functions by $AC_{L^{p}}([a,b],E)$. Let $t_{0}\in[a,b]$ be fixed, since $\eta^{\prime}:=[\gamma]$ is necessarily unique, the map
-$\Phi:AC_{L^{p}}([a,b],E)\to E\times L^{p}([a,b],E),\quad\eta\mapsto(\eta(t_{0}),\eta^{\prime})$
+We denote the space of all $L^{p}$-absolutely continuous functions by $AC_{L^{p}}([a,b],E)$.
+Let $t_{0}\in[a,b]$ be fixed, since $\eta^{\prime}:=[\gamma]$ is necessarily unique (see [32, Lemma 2.28]), the map
+$\Phi:AC_{L^{p}}([a,b],E)\to E\times L^{p}([a,b],E),\quad\eta\mapsto(\eta(t_{0}),\eta^{\prime})$ (2.2)
 is an isomorphism of vector spaces.
-We endow $AC_{L^{p}}([a,b],E)$ with the Hausdorff locally convex vector topology which makes $\Phi$ an isomorphism of topological vector space, which is independent of the choice of $t_{0}$.
-The following result will allow us to study differentiability of functions with values in $AC_{L^{p}}([a,b],E)$ [10, Lemma 3.9].
-Theorem 2.12
-Let $E$ be a Fréchet space and $p\in[1,\infty]$. Then the map
-$\Psi:AC_{L^{p}}([a,b],E)\to C([a,b],E)\times{L^{p}}([a,b],E),\quad\eta\mapsto(\eta,\eta^{\prime})$
+We endow $AC_{L^{p}}([a,b],E)$ with the Hausdorff locally convex vector topology which makes $\Phi$ an isomorphism of topological vector space (see [32, Definition 3.1]).
+The following result will allow us to study differentiability of functions with values in $AC_{L^{p}}([a,b],E)$.
+Lemma 2.13
+Let $E$ be a sequentially complete locally convex space and $p\in[1,\infty]$. Then the map
+$\Psi:AC_{L^{p}}([a,b],E)\to C([a,b],E)\times{L^{p}}([a,b],E),\quad\eta\mapsto(\eta,\eta^{\prime})$ (2.3)
 is a linear topological embedding with closed image.
+Proof.
+We let $I:E\times L^{p}([a,b],E)\to C([a,b],E)$ be the continuous map given by
+$I((x,[\gamma])(t):=x+\int_{a}^{t}\gamma(s)ds,\quad t\in[a,b]$
+for each $x\in E$ and $[\gamma]\in L^{p}([a,b],E)$.
+Let $\Phi:AC_{L^{p}}([a,b],E)\to E\times L^{p}([a,b],E)$, with $t_{0}:=a$, be the isomorphism of topological vector spaces as above.
+We consider the map
+$\Theta:E\times L^{p}([a,b],E)\to C([a,b],E)\times{L^{p}}([a,b],E),\quad(x,[\gamma])\mapsto\big{(}I(x,[\gamma]),[\gamma]\big{)}$
+which is continuous.
+Moreover, since the evaluation map $\varepsilon_{a}:C([a,b],E)\to E$, $\eta\mapsto\eta(a)$ is continuous, the map $\left(\Theta|^{\text{Im}(\Theta)}\right)^{-1}=(\varepsilon_{a},\text{id}_{L^{p}})$ is also continuous.
+Hence $\Psi$ is a topological embedding.
+Let $(\eta_{\alpha},\eta_{\alpha}^{\prime})_{\alpha}$ be a net in $\text{Im}(\Theta)$ that converges to $(\eta,[\gamma])\in C([a,b],E)\times{L^{p}}([a,b],E)$.
+By continuity of $\varepsilon_{a}$, the net $(\eta_{\alpha}(a))_{\alpha}$ converges to $\eta(a)\in E$, and by continuity of $\Theta$, the net $\left(\Theta(\eta_{\alpha}(a),\eta_{\alpha}^{\prime})\right)_{\alpha}$ converges to $\left(I(\eta(a),[\gamma]),[\gamma]\right)\in\text{Im}(\Theta)$.
+Since the net $\left(\Theta(\eta_{\alpha}(a),\eta_{\alpha}^{\prime})\right)_{\alpha}$ also converges to $(\eta,[\gamma])$, we have
+$I(\eta(a),[\gamma])=\eta.$
+Therefore $\eta^{\prime}=[\gamma]$ and $(\eta,[\gamma])\in\text{Im}(\Theta)$. ∎
 Remark 2.14
 Let $p\in[1,\infty]$.
-Since the inclusion map ${AC_{L^{p}}([a,b],E)\to C([a,b],E)}$ is continuous [24, Lemma 3.2], the topology on $AC_{L^{p}}([a,b],E)$ is independent of the choice of $t_{0}$ and finer than the compact-open topology.
+Since the inclusion map ${AC_{L^{p}}([a,b],E)\to C([a,b],E)}$ is continuous [33, Lemma 3.2], the topology on $AC_{L^{p}}([a,b],E)$ is independent of the choice of $t_{0}$ and finer than the compact-open topology.
 Hence the sets
@@
 $g(t)=a+\frac{t-c}{d-c}(b-a),\quad t\in[c,d].$
-Then the map
+Then $\eta\circ g\in AC_{L^{p}}([c,d],E)$ for each $\eta\in AC_{L^{p}}([a,b],E)$ and the map
 $AC_{L^{p}}(g,E):AC_{L^{p}}([a,b],E)\to AC_{L^{p}}([c,d],E),\quad\eta\mapsto\eta\circ g$
@@
 Lemma 2.16
-Let $E$ and $F$ be sequentially complete locally convex space, $p\in[1,\infty]$, $V\subseteq E$ be open subset and ${f:V\to F}$ be a $C^{1}$-map.
+Let $E$ and $F$ be sequentially complete locally convex spaces, $p\in[1,\infty]$, $V\subseteq E$ be open subset and ${f:V\to F}$ be a $C^{1}$-map.
 Then $f\circ\eta\in AC_{L^{p}}([a,b],F)$ for each $\eta\in AC_{L^{p}}([a,b],V)$.
 Lemma 2.17
-Let $E$ and $F$ be sequentially complete locally convex space, $p\in[1,\infty]$ and $k\in\mathbb{N}\cup\{0,\infty\}$.
+Let $E$ and $F$ be sequentially complete locally convex spaces, $p\in[1,\infty]$ and $k\in\mathbb{N}\cup\{0,\infty\}$.
 Let $V\subseteq E$ be open subset and ${f:V\to F}$ be a $C^{k+2}$-map, then the map
-$f^{*}:=AC_{L^{p}}([a,b],f):AC_{L^{p}}([a,b],V)\to AC_{L^{p}}([a,b],F),\quad\eta\mapsto f\circ\eta$
+$f_{*}:=AC_{L^{p}}([a,b],f):AC_{L^{p}}([a,b],V)\to AC_{L^{p}}([a,b],F),\quad\eta\mapsto f\circ\eta$
 is $C^{k}$. Moreover, we have
-$d(f^{*})(\eta,\eta_{1})=df\circ(\eta,\eta_{1})$
+$d(f_{*})(\eta,\eta_{1})=df\circ(\eta,\eta_{1})$
 for all $(\eta,\eta_{1})\in AC_{L^{p}}([a,b],V)\times AC_{L^{p}}([a,b],E)$.
 Remark 2.18
-For sequentially complete locally convex space $E$ and $F$ we have
+For sequentially complete locally convex spaces $E$ and $F$ we have
 $AC_{L^{p}}([a,b],E\times F)\cong AC_{L^{p}}([a,b],E)\times AC_{L^{p}}([a,b],F).$
-For the complex case of Lemma 2.16, we will use the following result [9, Lemma 2.5].
+Definition 2.19
+Let $E$ and $F$ be complex topological vector spaces, where is $F$ is locally convex, $U\subseteq E$ be an open subset and $f:U\to F$ be a mapping.
+We say that $f$ is complex analytic if it is continuous and, for each $x\in U$, there exists a 00-neighbouhood $V\subseteq E$ such that $x+V\subseteq U$ and certain continuous homogeneous polynomials $\beta_{n}:E\to F$ of degree $n$, such that $f$ admits the expansion: $f(x+y)=\sum_{n=0}^{\infty}\beta_{n}(y)$, for all $y\in V$.
+For our context, we present an application of [4, Proposition 7.7] to our particular case.
 Lemma 2.20
@@
 Let $E$ and $F$ be real locally convex spaces, $U\subseteq E$ be an open subset and $f:U\to F$ be a map.
-We say that $f$ is real analytic if it extends to a complex analytic map $f:V\to F_{\mathbb{C}}$ on some open neighborhood $V$ of $U$ in $E_{\mathbb{C}}$.
+We say that $f$ is real analytic if it extends to an analytic map $f:V\to F_{\mathbb{C}}$ on some open neighborhood $V$ of $U$ in $E_{\mathbb{C}}$, where $E_{\mathbb{C}}$ and $F_{\mathbb{C}}$ denotes the complexification of $E$ and $F$, respectively.
 Lemma 2.22
@@
 Then the map
-$f^{*}:=AC_{L^{p}}([a,b],f):AC_{L^{p}}([a,b],V)\to AC_{L^{p}}([a,b],F),\quad\eta\mapsto f\circ\eta$
+$f_{*}:=AC_{L^{p}}([a,b],f):AC_{L^{p}}([a,b],V)\to AC_{L^{p}}([a,b],F),\quad\eta\mapsto f\circ\eta$
 is $\mathbb{K}$-analytic.
@@
 First we consider the case $\mathbb{K}=\mathbb{C}$.
-By Lemma 2.16 the map $f^{*}$ is smooth and $d(f^{*})$ is complex linear in the second variable, hence by Lemma 2.18 the map $f$ is complex analytic.
-If $\mathbb{K}=\mathbb{R}$, then by definition the map $f$ has a complex analytic extension $\tilde{f}$, hence $(\tilde{f})^{*}$ is the complex analytic extension of $f^{*}$, whence $f^{*}$ is real analytic.
+By Lemma 2.17 the map $f_{*}$ is smooth and $d(f_{*})$ is complex linear in the second variable, hence by Lemma 2.20 the map $f_{*}$ is complex analytic.
+If $\mathbb{K}=\mathbb{R}$, then by definition the map $f$ has a complex analytic extension $\tilde{f}$, hence $(\tilde{f})_{*}$ is the complex analytic extension of $f_{*}$, whence $f_{*}$ is real analytic.
 ∎
@@
 Let $N$ be a smooth manifold modeled with on a sequentially complete locally convex space $E$ and $p\in[1,\infty]$.
-We say that a function $\eta:[a,b]\to N$ is $L^{p}$-absolutely continuous if it is continuous and there exists a partition $\{t_{0},...,t_{n}\}$ of $[a,b]$ such that for each $j\in\{1,...,n\}$, exists a chart $\varphi_{j}:U_{j}\to V_{j}$ that verify
+We say that a function $\eta:[a,b]\to N$ is $L^{p}$-absolutely continuous if it is continuous and there exists a partition $\{t_{0},...,t_{n}\}$ of $[a,b]$ such that for each $j\in\{1,...,n\}$, there exists a chart $\varphi_{j}:U_{j}\to V_{j}$ that verifies
   • i)
@@
 In this case, we say that these charts verify the definition of $L^{p}$-absolute continuity for $\eta$. If there is no confusion, we simple call $\eta$ absolutely continuous.
-For the case of absolute continuous functions with valued in a manifold, we have the following results (see [24]).
+For the case of absolutely continuous functions with values in a manifold, the following facts are available (see [33]).
 Lemma 2.24
@@
 If $f:M\to N$ is a $C^{1}$-map, then $f\circ\eta\in AC_{L^{p}}([a,b],N)$ for each $\eta\in AC_{L^{p}}([a,b],M)$.
-3 The space of absolutely continuous sections
 Definition 2.26
-Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$, $\pi_{TN}:TN\to N$ its vector bundle and $p\in[1,\infty]$. For $\eta\in AC_{L^{p}}([a,b],N)$ we define the set
-${\Gamma_{AC}(\eta)}:=\{\sigma\in AC_{L^{p}}([a,b],TN):\pi_{TN}\circ\sigma=\eta\}$
+Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$, $\pi_{TN}:TN\to N$ its tangent bundle and $p\in[1,\infty]$.
+For $\eta\in AC_{L^{p}}([a,b],N)$ we define the set
+${\Gamma_{AC}(\eta)}:=\{\sigma\in AC_{L^{p}}([a,b],TN):\pi_{TN}\circ\sigma=\eta\}$ (2.4)
 and we endow it with the pointwise operations, making it a vector space.
-For a partition $P_{n}=\{t_{0},...,t_{n}\}$ of $[a,b]$ and charts $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ of $N$ that verify the definition of absolute continuity for $\eta$.
+Consider a partition $P_{n}=\{t_{0},...,t_{n}\}$ of $[a,b]$ and charts $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ of $N$ that verify the definition of absolute continuity for $\eta$.
 Since $\eta([t_{i-1},t_{i}])\subseteq U_{i}$, we have
 $\sigma([t_{i-1},t_{i}])\subseteq TU_{i},\quad\text{for all }\sigma\in{\Gamma_{AC}(\eta)}$
-for each $i\in\{1,...,n\}$. We endow ${\Gamma_{AC}(\eta)}$ with the Hausdorff locally convex vector topology which is generated by the initial topology with respect to the mappings
-$h_{i}:{\Gamma_{AC}(\eta)}\to AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\sigma\mapsto h_{i}(\sigma)=d\varphi_{i}\circ\sigma|_{[t_{i-1},t_{i}]}$
+and for each $i\in\{1,...,n\}$. We endow ${\Gamma_{AC}(\eta)}$ with the Hausdorff locally convex vector topology which is the initial topology with respect to the linear mappings
+$h_{i}:{\Gamma_{AC}(\eta)}\to AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\sigma\mapsto h_{i}(\sigma)=d\varphi_{i}\circ\sigma|_{[t_{i-1},t_{i}]}$ (2.5)
 with $i\in\{1,...,n\}$.
@@
 Let $P=\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ be charts of $N$ that verify the definition of absolute continuity for $\eta$, then the map
-$\Phi_{\eta,P}:{\Gamma_{AC}(\eta)}\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\sigma\mapsto\left(d\varphi_{i}\circ\sigma|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}$
+$\Phi_{\eta,P}:{\Gamma_{AC}(\eta)}\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\sigma\mapsto\left(d\varphi_{i}\circ\sigma|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}$ (2.6)
 is a linear topological embedding with closed image given by the set of all elements $(\tau_{i})_{i=1}^{n}$ such that
@@
 The linear map $\Phi_{\eta,P}$ is continuous by the previous definition and it is injective.
-Let $i\in\{1,...,n\}$. If $W_{i}=\varphi_{i}(U_{i})\cap\varphi_{i+1}(U_{i+1})$, then the map
+Let $i\in\{1,...,n\}$. If $W_{i}=\varphi_{i+1}(U_{i}\cap U_{i+1})$, then the map
 $g_{i}:W_{i}\times E\to E,\quad(x,y)\mapsto d\varphi_{i}\circ(T\varphi_{i+1})^{-1}(x,y),$
@@
 $\tau_{i}(t_{i})=d\varphi_{i}\circ(T\varphi_{i+1})^{-1}\Big{(}\varphi_{i+1}\circ\eta(t_{i}),\tau_{i+1}(t_{i})\Big{)},\quad\text{ for all }i\in\{1,...,n-1\}$
-We will show that the image set of $\Phi_{\eta,P}$, denoted by $\text{Im}(\Phi_{\eta,P})$, coincide with the closed subspace $K$.
+We will show that the image of $\Phi_{\eta,P}$, denoted by $\text{Im}(\Phi_{\eta,P})$, coincides with the closed subspace $K$.
 Indeed, the space $\text{Im}(\Phi_{\eta,P})$ is contained in $K$ by definition of $\Phi_{\eta,P}$.
@@
 Corollary 2.29
-Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$.
-If $E$ is a Banach space (resp.
-a Fréchet space), then ${\Gamma_{AC}(\eta)}$ is a Banach space (resp.
-a Fréchet space).
+Let $N$ be a smooth manifold modeled on a $E$ Banach space (resp.
+Frechet space), $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$.
+Then the vector space ${\Gamma_{AC}(\eta)}$ is a Banach space (resp.
+Frechet space).
 Proof.
-This follows from the fact that each vector space $AC_{L^{p}}([t_{i-1},i],E)$ is a Banach space (resp. a Fréchet space). ∎
+This follows from the fact that each vector space $AC_{L^{p}}([t_{i-1},i],E)$ is a Banach space (resp. Frechet space) ∎
 Proposition 2.30
@@
 Proof.
-Let $E_{M}$ and $E_{N}$ be the modeled spaces of $M$ and $N$ respectively.
+Let $E_{M}$ and $E_{N}$ be the modeling spaces of $M$ and $N$ respectively.
 By Lemma 2.16 we have $f\circ\eta\in AC_{L^{p}}([a,b],N)$ and $Tf\circ\sigma\in AC_{L^{p}}([a,b],TN)$ for each $\sigma\in\Gamma_{AC}(\eta)$.
@@
 $\pi_{TN}\circ(Tf\circ\sigma)=f\circ\eta.$
-Thus $\widetilde{f}(\sigma)\in\Gamma_{AC}(f\circ\eta)$ for each $\sigma\in{\Gamma_{AC}(\eta)}$. The linearity of $\widetilde{f}$ is trivial.
-Without loss of generality, we can chose a partition $P=\{t_{0},...,t_{n}\}$ of $[a,b]$ such that there is exists the families of charts $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ and $\{(\phi_{i},V_{i}):i\in\{1,..,n\}\}$ that verify the definition of absolutely continuous for $\eta$ and $f\circ\eta$ respectively, such that
+Thus $\widetilde{f}(\sigma)\in\Gamma_{AC}(f\circ\eta)$ for each $\sigma\in{\Gamma_{AC}(\eta)}$. The linearity of $\widetilde{f}$ is clear.
+Without loss of generality, we can choose a partition $P=\{t_{0},...,t_{n}\}$ of $[a,b]$ such that there exist families of charts $\{(\varphi_{i},U_{i}):i\in\{1,...,n\}\}$ and $\{(\phi_{i},V_{i}):i\in\{1,...,n\}\}$ that verify the definition of absolute continuity for $\eta$ and $f\circ\eta$ respectively, such that
 $f(U_{i})\subseteq V_{i},\quad\text{for each }i\in\{1,...,n\}.$
-For $\sigma\in{\Gamma_{AC}(\eta)}$ and for each $i\in\{1,...,n\}$ we denote $\sigma_{i}=\sigma|_{[t_{i-1},t_{i}]}$ and we define the maps
-$F_{i}:AC_{L^{p}}([t_{i-1},t_{i}],E_{M})\to AC_{L^{p}}([t_{i-1},t_{i}],E_{N}),\quad\tau\mapsto\left(d\phi_{i}\circ Tf\circ d\varphi_{i}^{-1}\right)\circ\tau$
+For $\sigma\in{\Gamma_{AC}(\eta)}$ and for each $i\in\{1,...,n\}$ we denote $\eta_{i}=\eta|_{[t_{i-1},t_{i}]}$ and $\sigma_{i}=\sigma|_{[t_{i-1},t_{i}]}$.
+Moreover, for each $i\in\{1,...,n\}$ we define the maps
+$F_{i}:AC_{L^{p}}([t_{i-1},t_{i}],E_{M})\to AC_{L^{p}}([t_{i-1},t_{i}],E_{N}),\quad\tau\mapsto\left(d\phi_{i}\circ Tf\circ T\varphi_{i}^{-1}\right)\circ(\varphi_{i}\circ\eta_{i},\tau)$
 and
-$F:\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{M})\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{N}),\quad\tau\mapsto\left(\left(d\phi_{i}\circ Tf\circ d\varphi_{i}^{-1}\right)\circ\tau\right)_{i=1}^{n}$
+$F:\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{M})\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{N}),\quad\left(\tau_{i}\right)_{i=1}^{n}\mapsto\Big{(}F_{i}(\tau_{i})\Big{)}_{i=1}^{n}$
 which are continuous by Lemma 2.17. We will show that
 $F\left(\text{Im}\left(\Phi_{\eta,P}\right)\right)\subseteq\text{Im}\left(\Phi_{f\circ\eta,P}\right).$
-Let $i\in\{1,..,n-1\}$, if
-$\tau_{i}:=F_{i}(d\varphi_{i}\circ\sigma_{i})=d\phi_{i}\circ Tf\circ\sigma_{i}$
+Let $i\in\{1,..,n-1\}$. We denote
+$\tau_{i}:$ $=F_{i}(d\varphi_{i}\circ\sigma_{i})$
+$=(d\phi_{i}\circ Tf\circ d\varphi_{i}^{-1})(\varphi_{i}\circ\eta_{i},d\varphi_{i}\circ\sigma_{i})$
+$=d\phi_{i}\circ Tf\circ\sigma_{i}$
 then
-$d\varphi_{i}\circ(T\varphi_{i+1})^{-1}\Big{(}\varphi_{i+1}\circ\eta(t_{i}),\tau_{i+1}(t_{i})\Big{)}$ $=d\varphi_{i}\circ(T\varphi_{i+1})^{-1}\Big{(}\varphi_{i+1}\circ\eta(t_{i}),d\phi_{i+1}\circ Tf\circ\sigma_{i+1}(t_{i})\Big{)}$
-$=d\varphi_{i}\Big{(}\eta(t_{i}),Tf\circ\sigma_{i+1}(t_{i})\Big{)}$
-$=d\varphi_{i}\circ Tf\circ\sigma_{i+1}(t_{i})$
-$=d\varphi_{i}\circ Tf\circ\sigma_{i}(t_{i})$
+$d\phi_{i}\circ(T\phi_{i+1})^{-1}\Big{(}\phi_{i+1}\circ(f\circ\eta)(t_{i}),\tau_{i+1}(t_{i})\Big{)}$ $=d\phi_{i}\circ(T\phi_{i+1})^{-1}\Big{(}\phi_{i+1}\circ(f\circ\eta)(t_{i}),d\phi_{i+1}\circ Tf\circ\sigma_{i+1}(t_{i})\Big{)}$
+$=d\phi_{i}\circ Tf\circ\sigma_{i+1}(t_{i})$
+$=d\phi_{i}\circ Tf\circ\sigma_{i}(t_{i})$
 $=\tau_{i}(t_{i})$
@@
 Remark 2.31
-The topology of ${\Gamma_{AC}(\eta)}$ does not depend on the partition or charts chosen. Indeed, since the identity map $\text{id}_{M}:M\to M$ is smooth, by previous proposition the map
+The topology of ${\Gamma_{AC}(\eta)}$ does not depend on the partition or charts chosen. Indeed, since the identity map $\text{id}_{M}:M\to M$ is smooth, by the previous proposition the map
 $\widetilde{\text{id}_{M}}:{\Gamma_{AC}(\eta)}\to\Gamma_{AC}(\text{id}_{M}\circ\eta),\quad\sigma\mapsto T\text{id}_{M}\circ\sigma$
-is smooth regardless of the partition or charts chosen. Moreover, this map coincide with the identity map $\text{id}_{\Gamma}:{\Gamma_{AC}(\eta)}\to{\Gamma_{AC}(\eta)}$, $\sigma\mapsto\sigma$.
+is smooth regardless of the partition or charts chosen.
 Remark 2.32
@@
 $AC_{L^{p}}([t_{i-1},t_{i}],E)\to C([t_{i-1},t_{i}],E)$
-are continuous [24, Lemma 3.2], the inclusion map $J_{\Gamma}:\Gamma_{AC}(\gamma)\to\Gamma_{C}(\gamma)$ is also continuous.
-This implies that set
+is continuous [33, Lemma 3.2], by Proposition 2.27 the inclusion map $J_{\Gamma}:\Gamma_{AC}(\gamma)\to\Gamma_{C}(\gamma)$ is also continuous. This implies that the set
 $\mathcal{V}:=\{\sigma\in{\Gamma_{AC}(\eta)}:\sigma([a,b])\subseteq V\}$
@@
 Let $N_{1}$ and $N_{2}$ be smooth manifolds modeled on sequentially complete locally convex spaces, $p\in[1,\infty]$ and $\operatorname{pr}_{i}:N_{1}\times N_{2}\to N_{i}$ be the $i$-projection for $i\in\{1,2\}$.
-If $\eta_{1}\in AC_{L^{p}}([a,b],N_{1})$ and $\eta_{2}\in AC_{L^{p}}([a,b],N_{2})$ then the map
-$\widetilde{(\operatorname{pr}_{1},\operatorname{pr}_{2})}:\Gamma_{AC}(\eta_{1}\times\eta_{2})\to\Gamma_{AC}(\eta_{1})\times\Gamma_{AC}(\eta_{1}),\quad\sigma\mapsto(T\operatorname{pr}_{1},T\operatorname{pr}_{2})(\sigma)$
+If $\eta_{1}\in AC_{L^{p}}([a,b],N_{1})$ and $\eta_{2}\in AC_{L^{p}}([a,b],N_{2})$, then the map
+$\mathcal{P}:\Gamma_{AC}(\eta_{1},\eta_{2})\to\Gamma_{AC}(\eta_{1})\times\Gamma_{AC}(\eta_{1}),\quad\sigma\mapsto(T\operatorname{pr}_{1},T\operatorname{pr}_{2})(\sigma)$
 is a linear homeomorphism.
 Proof.
-This is consequence of Proposition 3.5, with $f=(\operatorname{pr}_{1},\operatorname{pr}_{2})$ and considering that the map $Tf=(T\operatorname{pr}_{1},T\operatorname{pr}_{2}):T(N_{1}\times N_{2})\to TN_{1}\times TN_{2}$ is a smooth diffeomorphism.
+By Proposition 2.30 the map $\mathcal{P}$ is continuous and clearly linear.
+Let $P=\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$; let $\{(\phi_{1,i},U_{1,i})\colon i\in\{1,\ldots,n\}\}$ and $\{(\phi_{2,i},U_{2,i})\colon i\in\{1,\ldots,n\}\}$ be families of charts of $N_{1}$ and $N_{2}$, respectively, that verify the definition of absolute continuity for $\eta_{1}$ and $\eta_{2}$, respectively.
+Then $\eta:=(\eta_{1},\eta_{2})\colon[a,b]\to N_{1}\times N_{2}$ is $L^{p}$-absolutely continuous and it is clear that the charts $\{(\phi_{1,i}\times\phi_{2,i},U_{1,i}\times U_{2,i})\colon i\in\{1,\ldots,n\}\}$ satisfy the condition of absolute continuity for $\eta$.
+For $j\in\{1,2\}$, consider the linear topological embedding
+$\Phi_{\eta_{j},P}\colon\Gamma_{\eta_{j}}\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{j}),\;\;\tau\mapsto(d\phi_{j,i}\circ\tau|_{[t_{i-1},t_{i}]})_{i=1}^{n}$
+where $E_{j}$ is the modeling space of $N_{j}$. Also
+$\Phi_{\eta,P}\colon\Gamma_{\eta}\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{1}\times E_{2}),\;\;\tau\mapsto((d\phi_{1,i}\times d\phi_{2,i})\circ\tau|_{[t_{i-1},t_{i}]})_{i=1}^{n}$
+is a linear topological embedding; here $T_{(x_{1},x_{2})}(N_{1}\times N_{2})$ is identified with $T_{x_{1}}N_{1}\times T_{x_{2}}N_{2}$. For $i\in\{1,\ldots,n\}$, let
+$\alpha_{i}\colon AC_{L^{p}}([t_{i-1},t_{i}],E_{1})\times AC_{L^{p}}([t_{i-1},t_{i}],E_{2})\to AC_{L^{p}}([t_{i-1},t_{i}],E_{1}\times E_{2})$
+be the map taking a pair $(f_{1},f_{2})$ of functions to the function $t\mapsto(f_{1}(t),f_{2}(t))$; we know that $\alpha_{i}$ is an isomorphism of topological vector spaces. Then also
+$\alpha\colon\left(\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{1})\right)\times\left(\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{2})\right)\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{1}\times E_{2}),$
+$((f_{i})_{i=1}^{n},(g_{i})_{i=1}^{n})\mapsto(\alpha(f_{i},g_{i}))_{i=1}^{n}$
+is an isomorphism of topological vector spaces.
+If $(f_{i})_{i=1}^{n}$ is in the image of $\Phi_{\eta_{1},P}$ and $(g_{i})_{i=1}^{n}$ is in the image of $\Phi_{\eta_{2},P}$, then $\alpha((f_{i})_{i=1}^{n},(g_{i})_{i=1}^{n})$ is in the image of $\Phi_{\eta,P}$, as the compatibility at the endpoints can be checked by considering the components in $E_{1}$ and $E_{2}$.
+We can therefore define a function
+$\Theta:=\Phi_{\eta,P}^{-1}\circ\alpha\circ(\Phi_{\eta_{1},P}\times\Phi_{\eta_{2},P})\colon\Gamma_{\eta_{1}}\times\Gamma_{\eta_{2}}\to\Gamma_{\eta},$
+which is continuous and linear.
+We readily check that $\mathcal{P}(\Theta(\sigma,\tau))=(\sigma,\tau)$ for all $\sigma\in\Gamma_{\eta_{1}}$ and $\tau\in\Gamma_{\eta_{2}}$.
+Hence $\mathcal{P}$ is surjective and thus bijective, with $\mathcal{P}^{-1}=\Theta$ a continuous map.
 ∎
@@
 For a partition $P=\{t_{0},...,t_{n}\}$ of $[a,b]$, the map
-$\rho:{\Gamma_{AC}(\eta)}\to\prod_{i=1}^{n}\Gamma_{AC}(\eta|_{[t_{i-1},t_{i}]}),\quad\sigma\mapsto(\sigma|_{[t_{i-1},t_{i}]})_{i=1}^{n}$
+$\rho:{\Gamma_{AC}(\eta)}\to\prod_{i=1}^{n}\Gamma_{AC}\left(\eta|_{[t_{i-1},t_{i}]}\right),\quad\sigma\mapsto\left(\sigma|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}$
 is a linear topological embedding with closed image.
 Proof.
-Let $j\in\{1,...,n\}$, If $P_{j}=\{t_{i-1},t_{i}\}$, since $\eta_{j}=\eta|_{[t_{j-1},t_{j}]}\in AC_{L^{p}}([t_{j-1},t_{j}],N)$, we have
-$\rho=\Big{(}\Phi_{\eta_{i},P_{i}}^{-1}\Big{)}_{i=1}^{n}\circ\Phi_{\eta,P}.$
+Let $j\in\{1,...,n\}$, If $P_{j}=\{t_{j-1},t_{j}\}$, since $\eta_{j}:=\eta|_{[t_{j-1},t_{j}]}\in AC_{L^{p}}([t_{j-1},t_{j}],N)$, we have
+$\rho=\left(\prod_{i=1}^{n}\Phi_{\eta_{i},P_{i}}^{-1}\right)\circ\Phi_{\eta,P}.$
 The image is given by the closed subspace
@@
 Thus $(\rho|^{\text{Im}(\rho)})^{-1}:\text{Im}(\rho)\mapsto{\Gamma_{AC}(\eta)}$ is well defined and
-$(\rho|^{\text{Im}(\rho)})^{-1}=\Phi_{\eta,P}^{-1}\circ\Big{(}\Phi_{\eta_{i},P_{i}}\Big{)}_{i=1}^{n}.$
+$\left(\rho|^{\text{Im}(\rho)}\right)^{-1}=\Phi_{\eta,P}^{-1}\circ\left(\prod_{i=1}^{n}\Phi_{\eta_{i},P_{i}}\right).$
 Proposition 2.35
 Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ and $p\in[1,\infty]$.
-If $g:[c,d]\to[a,b]$ a the map as in Lemma 2.14, then $\eta\circ g\in AC_{L^{p}}([c,d],N)$ for each $\eta\in AC_{L^{p}}([a,b],N)$.
+If $g:[c,d]\to[a,b]$ is the map as in Lemma 2.15, then $\eta\circ g\in AC_{L^{p}}([c,d],N)$ for each $\eta\in AC_{L^{p}}([a,b],N)$.
 Moreover, if $\eta\in AC_{L^{p}}([a,b],N)$, then the map
@@
 Proof.
-Let $\eta\in AC_{L^{p}}([a,b],N)$, we will show first that $\eta\circ g\in AC_{L^{p}}([c,d],N)$.
+Let $\eta\in AC_{L^{p}}([a,b],N)$.
+We will show first that $\eta\circ g\in AC_{L^{p}}([c,d],N)$.
 Let $P=\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ charts of $N$ that verify the definition of absolute continuity for $\eta$.
@@
 which are continuous by Lemma 2.15.
-Considering the topological embeddings $\Phi_{\eta,P}$ and $\Phi_{\eta\circ g,Q}$ (as Proposition 3.2) with the same family of charts, if $(\tau_{i})_{i=1}^{n}\in\text{Im}(\Phi_{\eta,P})$ we have
+Considering the topological embeddings $\Phi_{\eta,P}$ and $\Phi_{\eta\circ g,Q}$ (as in Proposition 2.27) with the same family of charts, if $(\tau_{i})_{i=1}^{n}\in\text{Im}(\Phi_{\eta,P})$ we have
 $\tau_{i}\circ g(s_{i})$ $=\tau_{i}(t_{i})$
@@
 $=d\varphi_{i}\circ(T\varphi_{i+1})^{-1}\Big{(}\varphi_{i+1}\circ\eta\circ g(s_{i}),\tau_{i+1}\circ g(s_{i})\Big{)}$
-for each $i\in\{1,...,n-1\}$. If $G=(G_{1},...,G_{n})$, then $\left(G\circ\Phi_{\eta,P}\right)(\sigma)\in\text{Im}(\Phi_{\eta\circ g,Q})$ and
+for each $i\in\{1,...,n-1\}$. If $G=(G_{1}\times...\times G_{n})$, then $\left(G\circ\Phi_{\eta,P}\right)(\sigma)\in\text{Im}(\Phi_{\eta\circ g,Q})$ and
 $L_{g}=\Phi_{\eta\circ g,Q}^{-1}\circ G\circ\Phi_{\eta,Q}.$
-Hence $L_{g}$ is continious and clearly linear. ∎
+Hence $L_{g}$ is continuous and clearly linear. ∎
 Proposition 2.36
@@
 $\epsilon:{\Gamma_{AC}(\eta)}\times[a,b]\to TN,\quad(\sigma,t)\mapsto\sigma(t)$
-is continuous and linear in the first argument. Moreover, for each $t\in[a,b]$ the point evaluation map
+is continuous and linear in the first argument. Moreover, for each $t\in[a,b]$, the point evaluation map
 $\epsilon_{t}:{\Gamma_{AC}(\eta)}\to TN,\quad\sigma\mapsto\sigma(t)$
@@
 is continuous and the evaluation map $\tilde{\epsilon}_{t}:\Gamma_{C}(\eta)\to TN$, $\sigma\mapsto\sigma(t)$ is smooth for each $t\in[a,b]$ (see [3]).
-Then $\epsilon=\tilde{\epsilon}\circ(J_{\Gamma},\text{Id}_{\mathbb{R}})$ and $\epsilon_{t}=\tilde{\epsilon}_{t}\circ J_{\Gamma}$, where $J_{\Gamma}:{\Gamma_{AC}(\eta)}\to\Gamma_{C}(\eta)$ is the inclusion map, which is continuous linear by Remark 3.7.
+Then $\epsilon=\tilde{\epsilon}\circ(J_{\Gamma}\times\text{Id}_{\mathbb{R}})$ and $\epsilon_{t}=\tilde{\epsilon}_{t}\circ J_{\Gamma}$, where $J_{\Gamma}:{\Gamma_{AC}(\eta)}\to\Gamma_{C}(\eta)$ is the inclusion map, which is continuous linear by Remark 2.32.
 ∎
@@
 Remark 3.2
-If a smooth manifold $N$ admits a local addition, then also its tangent manifold $TN$ admits a local addition [2, Lemma A.11].
-Moreover, each manifold which admits local addition also admits a normalized local addition [2, Lemma A.14].
+Every paracompact finite-dimensional smooth manifold admits a local addition.
+If a smooth manifold $N$ admits a local addition, then also its tangent manifold $TN$ admits a local addition [3, Lemma A.11].
+Moreover, each manifold which admits a local addition also admits a normalized local addition [3, Lemma A.14].
 From now we will assume that each local addition is normalized.
@@
 Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. We define the sets
-$\mathcal{V}_{\eta}:=\{\sigma\in{\Gamma_{AC}(\eta)}:\sigma([a,b])\subseteq\Omega\}.$
+$\mathcal{V}_{\eta}:=\{\sigma\in{\Gamma_{AC}(\eta)}:\sigma([a,b])\subseteq\Omega\}$
 which is open in ${\Gamma_{AC}(\eta)}$ by Remark 2.32 and
@@
 $\Psi_{\eta}^{-1}:\mathcal{U}_{\eta}\to\mathcal{V}_{\eta},\quad\gamma\mapsto\theta_{N}^{-1}\circ(\eta,\gamma).$
-The following Lemma is just an application of [3, Lemma 10.1] to our particular case.
+The following lemma is just an application of [4, Lemma 10.1] to our particular case.
 Lemma 3.4
 Let $E$ and $F$ sequentially complete locally convex spaces, $U\subseteq E$ open and $f:U\to F$ a map.
-If $F_{0}\subseteq F$ is a closed subset and $f(U)\subseteq F_{0}$, then $f:U\to F$ is smooth if and only if $f|^{F_{0}}:U\to F_{0}$ is smooth.
+If $F_{0}\subseteq F$ is a closed vector subspace and $f(U)\subseteq F_{0}$, then $f:U\to F$ is smooth if and only if $f|^{F_{0}}:U\to F_{0}$ is smooth.
 Theorem 3.5
@@
 Proof.
-We endow $AC_{L^{p}}([a,b],N)$ with the final topology with respect to the family of maps $\Psi_{\eta}\colon{\mathcal{V}}_{\eta}\to{\mathcal{U}}_{\eta}$, for each $\eta\in AC_{L^{p}}([a,b],N)$.
-If we define the maps $\Psi_{\eta}^{C}\colon{\mathcal{V}}_{\eta}^{C}\to{\mathcal{U}}_{\eta}^{C}$ on the space of continuous functions $C([a,b],N)$ for each $\eta\in C([a,b],N)$ as Remark 4.3, then the final topology on $C([a,b],N)$ coincide with its topology.
+We endow $AC_{L^{p}}([a,b],N)$ with the final topology with respect to the family of maps
+$\Psi_{\eta}\colon{\mathcal{V}}_{\eta}\to{\mathcal{U}}_{\eta},\quad\text{for each }\eta\in AC_{L^{p}}([a,b],N).$
+If we define the maps $\Psi_{\eta}^{C}\colon{\mathcal{V}}_{\eta}^{C}\to{\mathcal{U}}_{\eta}^{C}$ on the space of continuous functions $C([a,b],N)$ for each $\eta\in C([a,b],N)$ as in Remark 3.3, then the final topology on $C([a,b],N)$ with respect to this maps coincides with its compact-open topology.
 Hence the inclusion map
@@
 is continuous. Moreover, for each $\eta\in AC_{L^{p}}([a,b],N)$ the set
-$\mathcal{U}_{J(\eta)}^{C}:=\{\gamma\in C([a,b],N):(J(\eta),\gamma)([a,b])\subseteq\Omega^{\prime}\}$
-is open in $C([a,b],N)$, thus
+$\mathcal{U}_{\eta}^{C}:=\{\gamma\in C([a,b],N):(\eta,\gamma)([a,b])\subseteq\Omega^{\prime}\}$
+is open in $C([a,b],N)$, whence
 $\mathcal{U}_{\eta}=\mathcal{U}_{\eta}^{C}\cap AC_{L^{p}}([a,b],N)$
 is open in $AC_{L^{p}}([a,b],N)$.
-The goal is make the family $\{(\mathcal{U}_{\eta},\Psi_{\eta}^{-1}):\eta\in AC_{L^{p}}([a,b],N)\}$ an atlas for $AC_{L^{p}}([a,b],N)$ for a smooth manifold structure on $AC_{Lp}([a,b],N)$.
-We need to show that the charts are compatibles, i.e., the smoothness of the map
-$\Lambda_{\xi,\eta}:=\Psi_{\xi}^{-1}\circ\Psi_{\eta}:\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})\subseteq{\Gamma_{AC}(\eta)}\to\Gamma_{AC}(\xi),\quad\sigma\mapsto\theta_{N}^{-1}\circ(\xi,\Sigma\circ\sigma).$ (2)
-Let $R=\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$, let $\{(U_{\varphi_{i}},\varphi_{i}):i\in\{1,...n\}\}$ and $\{(U_{\phi_{i}},\phi_{i}):i\in\{1,...n\}\}$ be charts that verify the definition of absolutely continuous for $\eta$ and $\xi$ respectively.
-For $\sigma:[a,b]\to TN$, we denote $\sigma_{i}:=\sigma|_{[t_{i-1},t_{i}]}$.
+The goal is to make the family $\{(\mathcal{U}_{\eta},\Psi_{\eta}^{-1}):\eta\in AC_{L^{p}}([a,b],N)\}$ an atlas for $AC_{L^{p}}([a,b],N)$ for a smooth manifold structure on $AC_{Lp}([a,b],N)$.
+We need to show that the charts are compatible, i.e., the smoothness of the map
+$\Lambda_{\xi,\eta}:=\Psi_{\xi}^{-1}\circ\Psi_{\eta}:\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})\subseteq{\Gamma_{AC}(\eta)}\to\Gamma_{AC}(\xi),\quad\sigma\mapsto\theta_{N}^{-1}\circ(\xi,\Sigma\circ\sigma)$
+for each $\eta,\xi\in AC_{L^{p}}([a,b],N)$ such that the open set
+$\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})=\left(\Psi_{\eta}^{C}\right)^{-1}(\mathcal{U}_{\eta}^{C}\cap\mathcal{U}_{\xi}^{C})\cap{\Gamma_{AC}(\eta)}$
+is not empty.
+Let $R=\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$, let $\{(U_{\varphi_{i}},\varphi_{i}):i\in\{1,...,n\}\}$ and $\{(U_{\phi_{i}},\phi_{i}):i\in\{1,...,n\}\}$ be charts that verify the definition of absolute continuity for $\eta$ and $\xi$, respectively.
+For each $\sigma\in AC_{L^{p}}([a,b],TN)$, we denote $\sigma_{i}:=\sigma|_{[t_{i-1},t_{i}]}$ and
+$\Lambda_{\xi,\eta}(\sigma_{i}):=\Lambda_{\xi,\eta}(\sigma)\big{|}_{[t_{i-1},t_{i}]}.$
 We will study the smoothness of the composition
-$\Phi_{\xi,R}\circ\Lambda_{\xi,\eta}:\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})\to\text{Im}(\Phi_{\xi,R})\subset\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\sigma\mapsto\left(d\phi_{i}\circ\Lambda_{\xi,\eta}(\sigma_{i})\right)_{i=1}^{n}$
-that by Lemma 4.4, is equivalent to the smoothness of $\Lambda_{\xi,\eta}$. For each $i\in\{1,...,n\}$ we have
-$d\tilde{\varphi}_{i}\circ\Lambda_{\xi,\eta}(\sigma_{i})$ $=d\phi_{i}\circ\theta_{N}^{-1}\circ(\xi,\Sigma\circ\sigma_{i})$
-$=d\phi_{i}\circ\theta_{N}^{-1}\circ(\phi_{i}^{-1}\left(\phi_{i}\circ\xi\right),\Sigma\circ T\varphi_{i}^{-1}\circ T\varphi_{i}\circ\sigma_{i})$
-$=d\phi_{i}\circ\theta_{N}^{-1}\circ(\phi_{i}^{-1}\left(\phi_{i}\circ\xi\right),\Sigma\circ T\varphi_{i}^{-1}\left(\varphi_{i}\circ\eta,d\varphi_{i}\circ\sigma_{i}\right)).$
-Lets $H_{i}:\phi_{i}(U_{\phi_{i}})\times\varphi_{i}(U_{i})\times E\to E$ defined by
-$H_{i}(x,y,z)=d\phi_{i}\circ\theta_{N}^{-1}\circ(\phi_{i}^{-1}\left(x\right),\Sigma\circ T\varphi_{i}^{-1}(y,z)).$ (3)
-Since the function $H_{i}$ is smooth (because is composition of smooth functions) the map
-$AC([t_{i-1},t_{i}],H_{i}):AC_{L^{p}}([t_{i-1},t_{i}],E\times E\times E)\to AC_{L^{p}}([t_{i-1},t_{i}],E)\quad\alpha\mapsto H_{i}\circ\alpha$
-is smooth by Lemma 2.16.
-Doing the identification of products of $AC_{L^{p}}$ spaces (Remark 2.17), if we fix the functions $\phi_{i}\circ\xi\in AC_{L^{p}}([t_{i-1},t_{i}],E)$ and $\varphi_{i}\circ\eta\in AC_{L^{p}}([t_{i-1},t_{i}],E)$, we see that the map
-$\Theta_{i}:AC_{L^{p}}([t_{i-1},t_{i}],TU_{i})\to AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\tau\mapsto H_{i}\circ(\tilde{\varphi}_{i}\circ\xi,\varphi_{i}\circ\eta,d\varphi_{i}\circ\tau)$
+$\Phi_{\xi,R}\circ\Lambda_{\xi,\eta}:\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})$ $\to\text{Im}(\Phi_{\xi,R})\subset\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E)$
+$\sigma$ $\mapsto\left(d\phi_{i}\circ\Lambda_{\xi,\eta}(\sigma_{i})\right)_{i=1}^{n}$
+that by Lemma 3.4, is equivalent to the smoothness of $\Lambda_{\xi,\eta}$. For each $i\in\{1,...,n\}$, we denote $\eta_{i}:=\eta|_{[t_{i-1},t_{i}]}$ and $\xi_{i}:=\xi|_{[t_{i-1},t_{i}]}$ and we have
+$d{\phi}_{i}\circ\Lambda_{\xi,\eta}|_{[t_{i-1},t_{i}]}(\sigma_{i})$ $=d\phi_{i}\circ\theta_{N}^{-1}\circ(\xi_{i},\Sigma\circ\sigma_{i})$
+$=d\phi_{i}\circ\theta_{N}^{-1}\circ(\phi_{i}^{-1}\left(\phi_{i}\circ\xi_{i}\right),\Sigma\circ T\varphi_{i}^{-1}\circ T\varphi_{i}\circ\sigma_{i})$
+$=d\phi_{i}\circ\theta_{N}^{-1}\circ(\phi_{i}^{-1}\left(\phi_{i}\circ\xi_{i}\right),\Sigma\circ T\varphi_{i}^{-1}\left(\varphi_{i}\circ\eta_{i},d\varphi_{i}\circ\sigma_{i}\right)).$
+Because all of the functions involved are continuous and have an open domain, also the composition
+$H_{i}(x,y,z):=d\phi_{i}\circ\theta_{N}^{-1}\circ(\phi_{i}^{-1}\left(x\right),\Sigma\circ T\varphi_{i}^{-1}(y,z))$ (3.1)
+has an open domain $\mathcal{O}_{i}$. Hence the map $H_{i}:\mathcal{O}_{i}\to E$ is smooth. By Lemma 2.17, the map
+$AC_{L^{p}}([t_{i-1},t_{i}],H_{i}):AC_{L^{p}}([t_{i-1},t_{i}],\mathcal{O}_{i})\to AC_{L^{p}}([t_{i-1},t_{i}],E)\quad\alpha\mapsto H_{i}\circ\alpha$
+is smooth. Doing the identification of products of $AC_{L^{p}}$ spaces (Remark 2.18), if we fix the functions $\phi_{i}\circ\xi_{i}$ and $\varphi_{i}\circ\eta_{i}$, we have the continuous linear map
+$AC_{L^{p}}([t_{i-1},t_{i}],E)\to AC_{L^{p}}([t_{i-1},t_{i}],E\times E\times E),\quad\tau\mapsto(\phi_{i}\circ\xi_{i},\varphi_{i}\circ\eta_{i},\tau).$
+We write $W_{i}$ for the preimage of $AC_{L^{p}}([t_{i-1},t_{i}],\mathcal{O}_{i})$ under this map. Then the map
+$\Theta_{i}:W_{i}\to AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\tau\mapsto H_{i}\circ(\phi_{i}\circ\xi,\varphi_{i}\circ\eta,d\varphi_{i}\circ\tau)$
 is also smooth. Since the maps $h_{i}:{\Gamma_{AC}(\eta)}\to AC_{L^{p}}([t_{i-1},t_{i}],E),\sigma\mapsto d\phi_{i}\circ\sigma_{i}$ are continuous by definition of the topology, rewriting we have
@@
 for each $\sigma\in\Psi_{\eta}^{-1}(\mathcal{U}_{\eta}\cap\mathcal{U}_{\xi})$, hence $\Lambda_{\xi,\eta}$ is smooth. ∎
-Proceding in the same way, using the fact that composition of $\mathbb{K}$-analytics maps is $\mathbb{K}$-analytic and using the analytic version of Lemma 4.4 (see [9]), we can obtain the analogous case.
+Proceding in the same way, using the fact that compositions of $\mathbb{K}$-analytic maps are $\mathbb{K}$-analytic and using the analytic version of Lemma 3.4 (see [13]), we obtain the analogous case.
 Corollary 3.6
@@
 Proposition 3.7
-Let $N$ be smooth manifolds modeled on sequentially complete locally convex spaces which admits a local addition and $1\leq q\leq p\leq\infty$. Then
+Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space which admits a local addition and $1\leq q\leq p\leq\infty$. Then
 $AC_{L^{\infty}}([a,b],N)\subseteq AC_{L^{p}}([a,b],N)\subseteq AC_{L^{q}}([a,b],N)\subseteq AC_{L^{1}}([a,b],N)$
@@
 Let $\eta\in AC_{L^{p}}([a,b],N)$.
-Let $\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ be charts of $N$ such that verify the definition of absolutely continuous for $\eta$.
-Then, by Proposition 1.2, the smoothness of each inclusion in $\eta$ is equivalent to the smoothness of each inclusion map
-$\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],N)\to\prod_{i=1}^{n}AC_{L^{q}}([t_{i-1},t_{i}],N),\quad(\eta)_{i=1}^{n}\mapsto(\eta)_{i=1}^{n}.$
-Since each inclusion $AC_{L^{p}}([t_{i-1},t_{i}],N)\to AC_{L^{q}}([t_{i-1},t_{i}],N)$ is continuous linear [24, Remark 3.2], the inclusions are smooth. ∎
+Let $\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ be charts of $N$ that verify the definition of absolute continuity for $\eta$.
+By [33, Remark 3.2], we know that each inclusion map $AC_{L^{p}}([t_{i-1},t_{i}],E)\to AC_{L^{q}}([t_{i-1},t_{i}],E)$ is continuous linear, hence $AC_{L^{p}}([a,b],N)\subseteq AC_{L^{q}}([a,b],N)$.
+By Proposition 2.27, the inclusion is smooth since each map
+$\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E)\to\prod_{i=1}^{n}AC_{L^{q}}([t_{i-1},t_{i}],E),\quad(\eta_{i})_{i=1}^{n}\mapsto(\eta_{i})_{i=1}^{n}.$
+is continuous linear. ∎
 Proposition 3.8
-Let $M$ and $N$ be smooth manifolds modeled on sequentially complete locally convex spaces $E$ which admits a local addition, $p\in[1,\infty]$ and $k\in\mathbb{N}\cup\{0,\infty\}$.
+Let $M$ and $N$ be smooth manifolds modeled on sequentially complete locally convex spaces which admits a local addition, $p\in[1,\infty]$ and $k\in\mathbb{N}\cup\{0,\infty\}$.
 If $f:M\to N$ is a $C^{k+2}$-map, then the map
@@
 Proof.
+Let $E_{M}$ and $E_{N}$ be the modeleding space of $M$ and $N$, respectively.
 Let $(\Omega_{M},\Sigma_{M})$ and $(\Omega_{N},\Sigma_{N})$ be local additions on $M$ and $N$ respectively.
-The map make sense by Lemma 2.23.
-Let $\eta\in AC_{L^{p}}([a,b],N)$, $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ and $(\mathcal{U}_{f\circ\eta},\Psi_{f\circ\eta}^{-1})$ be charts in $\eta\in AC_{L^{p}}([a,b],N)$ and $f\circ\eta\in AC_{L^{p}}([a,b],N)$ respectively, if $\theta_{N}=(\pi_{TN},\Sigma_{N})$ then we define
-$F(\sigma):=\Psi_{f\circ\eta}^{-1}\circ AC_{L^{p}}([a,b],f)\circ\Psi_{\eta}=\theta_{N}^{-1}\circ\Big{(}(f\circ\eta),(f\circ\Sigma_{M})\circ\sigma\Big{)}$
+The map makes sense by Lemma 2.25.
+Let $\eta\in AC_{L^{p}}([a,b],M)$, $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ and $(\mathcal{U}_{f\circ\eta},\Psi_{f\circ\eta}^{-1})$ be charts around $\eta\in AC_{L^{p}}([a,b],M)$ and $f\circ\eta\in AC_{L^{p}}([a,b],N)$, respectively.
+We see that the set
+$\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}([a,b],f)^{-1}(\mathcal{U}_{f\circ\eta})\right)={\Gamma_{AC}(\eta)}\cap\left(\Psi_{\eta}^{C}\right)^{-1}\left(\mathcal{U}_{\eta}^{C}\cap AC_{L^{p}}([a,b],f)^{-1}(\mathcal{U}_{f\circ\eta}^{C})\right)$
+is open in ${\Gamma_{AC}(\eta)}$. If $\theta_{N}=(\pi_{TN},\Sigma_{N})$, then we define
+$F(\sigma):=\Psi_{f\circ\eta}^{-1}\circ AC_{L^{p}}([a,b],f)\circ\Psi_{\eta}(\sigma)=\theta_{N}^{-1}\circ\Big{(}(f\circ\eta),(f\circ\Sigma_{M})\circ\sigma\Big{)}$
 for all $\sigma\in\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}([a,b],f)^{-1}(\mathcal{U}_{f\circ\eta})\right)$.
-Proceeding as the proof of Theorem 3.1, choosing the corresponding partition $P=\{t_{0},...,t_{n}\}$ of $[a,b]$ and the families of charts $\{(U_{\varphi_{i}},\varphi_{i}):i\in\{1,...,n\}\}$ and $\{(U_{\phi_{i}},{\phi}_{i}):i\in\{1,...,n\}\}$ that verify the definition of absolute continuity for $\eta$ and $f\circ\eta$ respectively, we denote $\sigma_{i}=\sigma|_{[t_{i-1},t_{i}]}$.
+Proceeding as the proof of Theorem 3.5, choosing the corresponding partition $P=\{t_{0},...,t_{n}\}$ of $[a,b]$ and the families of charts $\{(U_{\varphi_{i}},\varphi_{i}):i\in\{1,...,n\}\}$ and $\{(U_{\phi_{i}},{\phi}_{i}):i\in\{1,...,n\}\}$ that verify the definition of absolute continuity for $\eta$ and $f\circ\eta$ respectively, we denote $\sigma_{i}:=\sigma|_{[t_{i-1},t_{i}]}$, $\eta_{i}:=\eta|_{[t_{i-1},t_{i}]}$ and
+$F(\sigma)|_{[t_{i-1},t_{i}]}:=F(\sigma_{i}).$
 We will study the continuity of the map
 $\Phi_{f\circ\eta}\circ F:\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}([a,b],f)^{-1}(\mathcal{U}_{f\circ\eta})\right)\to\text{Im}(\Phi_{f\circ\eta}),\quad\sigma\mapsto\left(d\phi_{i}\circ F(\sigma_{i})\right)_{i=1}^{n}$
-where $\Phi_{f\circ\eta}$ is the topological embedding as Proposition 3.2.
+where $\Phi_{f\circ\eta}$ is the topological embedding as in Proposition 2.27.
 For each $i\in\{1,...,n\}$ and $\sigma\in\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}([a,b],f)^{-1}(\mathcal{U}_{f\circ\eta})\right)$ we have
-$d\phi_{i}\circ F(\sigma)$ $=d\phi_{i}\circ\theta_{N}^{-1}\circ\Big{(}f\circ\eta,f\circ\Sigma_{M}\circ\sigma\Big{)}$
-$=d\phi_{i}\circ\theta_{N}^{-1}\circ\Big{(}\phi_{i}^{-1}\circ\phi_{i}\circ f\circ\eta,f\circ\Sigma_{M}\circ T\varphi_{i}^{-1}\left(\varphi_{i}\circ\eta,d\varphi_{i}\circ\sigma_{i}\right)).$
-We define the map $H_{i}:\phi_{i}(U_{\phi_{i}})\times\varphi_{i}(U_{\varphi_{i}})\times E_{M}\to E_{N}$ given by
-$H_{i}(x,y,z)=d\phi_{i}\circ\theta_{N}^{-1}\circ\Big{(}\phi_{i}^{-1}(x),f\circ\Sigma_{M}\circ T\varphi_{i}^{-1}(y,z)\Big{)}$
-which is a $C^{k+2}$-map. Since $\phi_{i}\circ f\circ\eta\in AC_{Lp}([t_{i-1},t_{i}],E_{N})$ and $\varphi_{i}\circ\eta\in AC_{Lp}([t_{i-1},t_{i}],E_{M})$, then the map
-$\Theta_{i}:AC_{L^{p}}([t_{i-1},t_{i}],E_{M})\to AC_{L^{p}}([t_{i-1},t_{i}],E_{N})\quad\tau\mapsto H_{i}\circ(\phi_{i}\circ f\circ\eta,\varphi_{i}\circ\eta,\tau)$
-is $C^{k}$.
-Since the maps $h_{i}:{\Gamma_{AC}(\eta)}\to AC_{L^{p}}([t_{i-1},t_{i}],E_{M})$, $d\sigma\mapsto\varphi_{i}\circ\sigma_{[t_{i-1},t_{i}]}$ are continuous linear by definition, rewriting we have
+$d\phi_{i}\circ F(\sigma)|_{[t_{i-1},t_{i}]}$ $=d\phi_{i}\circ\theta_{N}^{-1}\circ\Big{(}f\circ\eta_{i},f\circ\Sigma_{M}\circ\sigma_{i}\Big{)}$
+$=d\phi_{i}\circ\theta_{N}^{-1}\circ\Big{(}\phi_{i}^{-1}\circ\phi_{i}\circ f\circ\eta_{i},f\circ\Sigma_{M}\circ T\varphi_{i}^{-1}\left(\varphi_{i}\circ\eta_{i},d\varphi_{i}\circ\sigma_{i}\right)\Big{)}.$
+Because all of the functions involved are continuous and have an open domain, also the composition
+$H_{i}(x,y,z):=d\phi_{i}\circ\theta_{N}^{-1}\circ\Big{(}\phi_{i}^{-1}(x),f\circ\Sigma_{M}\circ T\varphi_{i}^{-1}(y,z)\Big{)}$ (3.2)
+has an open domain $\mathcal{O}_{i}$ in $E_{N}\times E_{M}\times E_{M}$.
+Hence the map $H_{i}:\mathcal{O}_{i}\to E_{N}$ is smooth.
+Fixing the absolutely continuous functions $\phi_{i}\circ f\circ\eta_{i}$ and $\varphi_{i}\circ\eta_{i}$, we consider the continuous linear map
+$AC_{L^{p}}([t_{i-1},i],E_{M})\to AC_{L^{p}}([t_{i-1},i],E_{N}\times E_{M}\times E_{M}),\quad\tau\mapsto(\phi_{i}\circ f\circ\eta,\varphi_{i}\circ\eta,\tau)$
+and we write $W_{i}$ for the preimage of $AC_{L^{p}}([a,b],\mathcal{O}_{i})$ under this map. Then the map
+$\Theta_{i}:W_{i}\to AC_{L^{p}}([t_{i-1},t_{i}],E_{N})\quad\tau\mapsto H_{i}\circ(\phi_{i}\circ f\circ\eta_{i},\varphi_{i}\circ\eta_{i},\tau)$
+is $C^{k}$. Since the maps $h_{i}:{\Gamma_{AC}(\eta)}\to AC_{L^{p}}([t_{i-1},t_{i}],E_{M})$, $\sigma\mapsto d\varphi_{i}\circ\sigma_{i}$ are continuous linear by definition, rewriting we have
 $\Phi_{f\circ\eta}\circ F(\sigma)=\left(\Theta_{i}\circ h_{i}(\sigma)\right)_{i=1}^{n}$
@@
 Corollary 3.9
-If $f:M\to N$ is a $\mathbb{K}$-analytic map between $\mathbb{K}$-analytic manifolds modeled on sequentially complete locally convex spaces which admits a $\mathbb{K}$-analytic local addition and $p\in[1,\infty]$.
+Let $f:M\to N$ be a $\mathbb{K}$-analytic map between $\mathbb{K}$-analytic manifolds modeled on sequentially complete locally convex spaces which admit $\mathbb{K}$-analytic local additions and $p\in[1,\infty]$.
 Then the map
@@
 Remark 3.10
-The manifold structures for $AC_{L^{p}}([a,b],N)$ given by different local additions are diffeomorphic.
+The manifold structures for $AC_{L^{p}}([a,b],N)$ given by different local additions coincide.
 Indeed, since the identity map $\text{id}_{N}:N\to N$ is smooth regardless of the chosen local addition, the map
-$AC([a,b],\text{id}_{M}):AC_{L^{p}}([a,b],N)\to AC_{L^{p}}([a,b],N),\quad\eta\to\text{id}_{M}\circ\eta$
+$AC([a,b],\text{id}_{M}):AC_{L^{p}}([a,b],N)\to AC_{L^{p}}([a,b],N),\quad\eta\to\text{id}_{N}\circ\eta$
 is, again, smooth regardless of the chosen local addition in each space.
@@
 The inclusion map $J:AC_{L^{p}}([a,b],N)\to C([a,b],N)$ is smooth.
-Indeed, let $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ and $(\mathcal{U}_{J(\eta)}^{C},\Psi_{J(\eta)}^{-1})$ be charts in $\eta\in AC_{L^{p}}([a,b],N)$ and $J(\eta)\in C([a,b],N)$ respectively, then
-$\Psi_{J(\eta)}^{-1}\circ J\circ\Psi_{\eta}^{-1}(\sigma):\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap J^{-1}(\mathcal{U}_{j(\eta)})\right)\subseteq{\Gamma_{AC}(\eta)}\to\Gamma_{C}(\eta)$
+Indeed, let $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ and $(\mathcal{U}_{\eta}^{C},\left(\Psi_{\eta}^{C}\right)^{-1})$ be charts around $\eta\in AC_{L^{p}}([a,b],N)$ and $\eta\in C([a,b],N)$ respectively, then
+$\Psi_{\eta}^{-1}\circ J\circ\Psi_{\eta}^{C}(\sigma):\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap J^{-1}(\mathcal{U}_{\eta}^{C})\right)\subseteq{\Gamma_{AC}(\eta)}\to\Gamma_{C}(\eta)$
 is a restriction of the inclusion map ${\Gamma_{AC}(\eta)}\to\Gamma_{C}(\eta)$.
@@
 $AC_{L^{p}}([a,b],U):=\{\eta\in AC_{L^{p}}([a,b],N):\eta([a,b])\subseteq U\}.$
-coincides with the manifold structure on $AC_{L^{p}}([a,b],U)$ constructed with the local addition
-$\Sigma|_{\Omega\cap TU}:\Omega\cap TU\to U.$
+coincides with the manifold structure on $AC_{L^{p}}([a,b],U)$.
 Proposition 3.12
-Let $N_{1}$ and $N_{2}$ be smooth manifolds with local addition modeled on a sequentially complete locally convex space which admits a local addition, $p\in[1,\infty]$ and let $\operatorname{pr}_{i}:N_{1}\times N_{2}\to N_{i}$ be the i-th projection for $i\in\{1,2\}$.
+Let $N_{1}$ and $N_{2}$ be smooth manifolds with local addition modeled on a sequentially complete locally convex spaces which admit a local addition, $p\in[1,\infty]$ and let $\operatorname{pr}_{i}:N_{1}\times N_{2}\to N_{i}$ be the i-th projection for $i\in\{1,2\}$.
 Then the map
@@
 Proof.
-By previous remark, if $(\Omega_{1},\Sigma_{1})$ and $(\Omega_{1},\Sigma_{1})$ are the local addition on $N_{1}$ and $N_{2}$ respectively, then we can assume that the local additions on $N_{1}\times N_{2}$ is
+By the previous remark, if $(\Omega_{1},\Sigma_{1})$ and $(\Omega_{1},\Sigma_{1})$ are the local addition on $N_{1}$ and $N_{2}$ respectively, then we can assume that the local addition on $N_{1}\times N_{2}$ is
 $\Sigma:=\Sigma_{1}\times\Sigma_{2}:\Omega_{1}\times\Omega_{2}\to N_{1}\times N_{2}$
@@
 $\mathcal{Q}:\Gamma_{AC}(\eta)\to\Gamma_{AC}(\eta_{1})\times\Gamma_{AC}(\eta_{2}),\quad\tau\mapsto(\text{q}_{1},\text{q}_{2})\circ\tau$
-is an diffeomorphism of vector spaces, where $\text{q}_{1}$ and $\text{q}_{2}$ are the corresponding projection of the space, we have
+is an isomorphism of topological vector spaces, where $\text{q}_{1}$ and $\text{q}_{2}$ are the corresponding projection of the space, we have
 $\Psi_{\eta}^{-1}\circ\mathcal{P}^{-1}\circ(\Psi_{\eta_{1}}\times\Psi_{\eta_{2}})(\sigma_{1},\sigma_{2})$ $=(\pi_{N_{1}\times N_{2}},\Sigma)^{-1}\circ\left(\eta,\mathcal{P}^{-1}\circ(\Sigma_{1}\times\Sigma_{2})(\sigma_{1},\sigma_{2})\right)$
@@
 Proof.
-It is clear that the map $T$ is well defined and injective. Let denote $\text{Im}(T)$ the image set of the map $T$, then
+It is clear that the map $T$ is well defined and injective. Let $\text{Im}(T)$ be the image of the map $T$. Then
 $\text{Im}(T)=\{(\gamma_{i})_{i=1}^{n}\in\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],N):\gamma_{i}(t_{i})=\gamma_{i+1}(t_{i})\text{ for all }i\in\{1,...,n-1\}\}.$
 Let $\gamma:=(\gamma_{i})_{i=1}^{n}\in\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],N)$.
-For each $i\in\{1,...,n\}$ let $\Psi_{\gamma_{i}}^{-1}:\mathcal{U}_{i}\to\mathcal{V}_{i}$ be charts of $\gamma_{i}$, then the map
+For each $i\in\{1,...,n\}$ let $\Psi_{\gamma_{i}}^{-1}:\mathcal{U}_{i}\to\mathcal{V}_{i}$ be charts around $\gamma_{i}$.
+Then the map
 $\Psi_{\gamma}^{-1}:=\prod_{i=1}^{n}\Psi_{\gamma_{i}}^{-1}:\prod_{i=1}^{n}\mathcal{U}_{i}\to\prod_{i=1}^{n}\mathcal{V}_{i}$
-is a chart for $\gamma$.
-Let $\eta\in AC_{L^{p}}([a,b],N)$ and $\tilde{\eta}=T(\eta)$, then $\Psi_{\tilde{\eta}}\circ T\circ\Psi_{\eta}^{-1}$ is just restriction of the product of restriction of the smooth maps
+is a chart around $\gamma$.
+Let $\eta\in AC_{L^{p}}([a,b],N)$ and $\tilde{\eta}=T(\eta)$, then $\Psi_{\tilde{\eta}}^{-1}\circ T\circ\Psi_{\eta}$ is just restriction of the product of the restrictions of the smooth maps
 ${\Gamma_{AC}(\eta)}\to\prod_{i=1}^{n}\Gamma_{AC}(\eta_{i}),\quad\sigma\mapsto\left(\sigma|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}$
-thus $T$ is a smooth.
-Let $\tilde{T}:AC_{L^{p}}([a,b],N)\to\text{Im}(T)$, $\eta\mapsto T(\eta)$, then analogously to Proposition 3.9, we can see that $\tilde{T}^{-1}$ is smooth.
-Hence $T$ it is a diffeomorphism onto its image.
-Now we will show that the image $\text{Im}(T)$ is a submanifold.
-Let $\gamma=(\gamma_{i})_{i=1}^{n}\in\text{Im}(T)$ with charts as before, then for each $i\in\{1,...,n-1\}$ and $\xi=\left(\xi_{i}\right)_{j=1}^{n}\in\text{Im}(T)\cap\prod_{j=1}^{n}\mathcal{U}_{j}$ we have
+thus $T$ is a smooth. Now we will show that the image $\text{Im}(T)$ is a submanifold.
+Let $\gamma=(\gamma_{i})_{i=1}^{n}\in\text{Im}(T)$ with charts as before, then for each $i\in\{1,...,n-1\}$ and
+$\xi=\left(\xi_{i}\right)_{j=1}^{n}\in\text{Im}(T)\cap\prod_{j=1}^{n}\mathcal{U}_{j}$ we have
 $\Psi_{\gamma_{i}}^{-1}(\xi_{i})(t_{i})$ $=\theta_{N}^{-1}\circ(\gamma_{i},\xi_{i})(t_{i})$
@@
 $K:=\{\left(\sigma_{i}\right)_{i=1}^{n}\in\prod_{i=1}^{n}\Gamma_{AC}(\gamma_{i}):\sigma_{i}(t_{i})=\sigma_{i+1}(t_{i})\text{ for all }i\in\{1,...,n-1\}\}.$
-Then $\Psi_{\gamma}^{-1}|_{\mathcal{U}_{\gamma}\cap\text{Im}(T)}:\text{Im}(\Psi)\cap\prod_{i=1}^{n}\mathcal{U}_{i}\to K\cap\prod_{i=1}^{n}\mathcal{V}_{i}$ is a chart of $\text{Im}(T)$, making $\text{Im}(T)$ a smooth submanifold.
+Then $\Psi_{\gamma}^{-1}|_{\mathcal{U}_{\gamma}\cap\text{Im}(T)}:\text{Im}(\Psi)\cap\prod_{i=1}^{n}\mathcal{U}_{i}\to K\cap\prod_{i=1}^{n}\mathcal{V}_{i}$ is a chart of $\text{Im}(T)$, making $\text{Im}(T)$ a smooth submanifold and the map $\tilde{T}:AC_{L^{p}}([a,b],N)\to\text{Im}(T)$, $\eta\mapsto T(\eta)$ a diffeomorphism.
+∎
 Remark 3.14
-Let $N$ is a $\mathbb{K}$-analytic manifold modeled on a sequentially complete locally convex space which admits a $\mathbb{K}$-analytic local addition and $p\in[1,\infty]$.
+Let $N$ be a $\mathbb{K}$-analytic manifold modeled on a sequentially complete locally convex space which admits a $\mathbb{K}$-analytic local addition and $p\in[1,\infty]$.
 Since every continuous linear operator is analytic, the isomorphism
@@
 Proposition 3.15
-Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits a local addition, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. If
-$g:[c,d]\to[a,b]$ is a map as in Lemma 2.14, then the map
+Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits a local addition, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$.
+If $g:[c,d]\to[a,b]$ is a map as in Lemma 2.15, then the map
 $AC_{L^{p}}(g,N):AC_{L^{p}}([a,b],N)\to AC_{L^{p}}([c,d],N),\quad\eta\mapsto\eta\circ g$
@@
 By Proposition 2.35 we know that the map is well defined.
-Let $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ and $(\mathcal{U}_{\eta\circ g},\Psi_{\eta\circ g}^{-1})$ be charts in $\eta\in AC_{L^{p}}([a,b],N)$ and $\eta\circ g\in AC_{L^{p}}([c,d],N)$ respectively, then we have
+Let $(\mathcal{U}_{\eta},\Psi_{\eta}^{-1})$ and $(\mathcal{U}_{\eta\circ g},\Psi_{\eta\circ g}^{-1})$ be charts around $\eta\in AC_{L^{p}}([a,b],N)$ and $\eta\circ g\in AC_{L^{p}}([c,d],N)$ respectively, then we have
 $\Psi_{\eta\circ g}^{-1}\circ AC_{L^{p}}(g,M)\circ\Psi_{\eta}(\sigma)=\theta_{N}^{-1}\circ(\eta\circ g,\Sigma\circ(\sigma\circ g))$
-for all $\sigma\in\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}(g,N)^{-1}(\mathcal{U}_{\eta\circ g})\right)$.
-Let $\alpha=\eta\circ g:[c,d]\to N$ and $\tau=\sigma\circ g:[c,d]\to TN$, then both are absolutely continuous, with $\pi_{TN}\circ\tau=\alpha$, then $\tau\in\Gamma_{AC}(\alpha)$ and
+for all $\sigma\in\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap AC_{L^{p}}(g,N)^{-1}(\mathcal{U}_{\eta\circ g})\right)$. This set coincides with
+$\Psi_{\eta}^{-1}\left(\mathcal{U}_{\eta}\cap C(g,N)^{-1}(\mathcal{U}_{\eta\circ g}^{C})\right)$
+which is open given by the continuity of the map $C(g,N)$.
+Let $\alpha=\eta\circ g:[c,d]\to N$ and $\tau=\sigma\circ g:[c,d]\to TN$.
+Then both are absolutely continuous, with $\pi_{TN}\circ\tau=\alpha$, now $\tau\in\Gamma_{AC}(\alpha)$ and
 $\Psi_{\eta\circ g}^{-1}\circ AC_{L^{p}}(g,M)\circ\Psi_{\eta}(\sigma)$ $=\theta_{N}^{-1}\circ(\alpha,\Sigma\circ\tau)$
@@
 Hence, $\Psi_{\eta\circ g}^{-1}\circ AC_{L^{p}}(g,M)\circ\Psi_{\eta}$ is a restriction of the map
-${\Gamma_{AC}(\eta)}\to{\Gamma_{AC}(\eta\circ g)},\quad\sigma\to\sigma\circ g$
+${\Gamma_{AC}(\eta)}\to{\Gamma_{AC}(\eta\circ g)},\quad\sigma\mapsto\sigma\circ g$
 which is continuous linear by Proposition 2.35. ∎
@@
 is continuous and the point evaluation $(\varepsilon_{c})_{t}:C([a,b],N)\to N$, $\eta\mapsto\eta(t)$ is smooth for each $t\in[a,b]$ (see [1]).
-Since the inclusion map $J:AC_{L^{p}}([a,b],N)\to C([a,b],N)$ is smooth, we have $\varepsilon=\varepsilon_{c}\circ(J,\text{Id}_{\mathbb{R}})$ and $\varepsilon_{t}=(\varepsilon_{c})_{t}\circ J$ for each $t\in[a,b]$.
+Since the inclusion map $J:AC_{L^{p}}([a,b],N)\to C([a,b],N)$ is smooth, the assumptions follow from the observation that $\varepsilon=\varepsilon_{c}\circ(J\times\text{Id}_{[a,b]})$ and $\varepsilon_{t}=(\varepsilon_{c})_{t}\circ J$ for each $t\in[a,b]$.
 ∎
@@
 If $f:L\times M\to N$ is a $C^{2}$-map and $\gamma\in AC_{L^{p}}([a,b],L)$ is fixed, then the map
-$F:[a,b]\times M\to N,\quad F(t,m)=f(\gamma(t),m)$
+$F:[a,b]\times M\to N,\quad(t,p)\mapsto f(\gamma(t),p)$
 is continuous.
 Proof.
-For each $(t,m)\in[a,b]\times M$ we have
-$F(t,m)=AC_{L^{p}}([a,b],f)\circ(\varepsilon(\gamma,\cdot)\times\text{id}_{M})(t,m)$
-where $\varepsilon$ is the evaluation map as before. Therefore, the map $F$ is composition of continuous maps. ∎
+It follows from the fact that $F=f\circ(\gamma\times\text{id}_{M})$. ∎
 Proposition 3.19
@@
 Proof.
-Let consider the local addition $\Sigma:\Omega\to N$ and $\theta_{N}:\Omega\to\Omega^{\prime}$ as in Definition 4.1.
-Let $(U,\varphi)$ be a chart in $q\in N$ such that $\{q\}\times U\subseteq\Omega^{\prime}$ and $(\mathcal{U}_{\zeta_{q}},\Psi_{\zeta_{q}}^{-1})$ be a chart in $\zeta_{q}\in AC_{L^{p}}([a,b],N)$.
+Consider the local addition $\Sigma:\Omega\to N$ and $\theta_{N}:\Omega\to\Omega^{\prime}$ as in Definition 3.1.
+Let $(U,\varphi)$ be a chart around $q\in N$ such that $\{q\}\times U\subseteq\Omega^{\prime}$ and $(\mathcal{U}_{\zeta_{q}},\Psi_{\zeta_{q}}^{-1})$ be a chart in $\zeta_{q}\in AC_{L^{p}}([a,b],N)$.
 If $x\in\varphi\left(U\cap\zeta^{-1}(\mathcal{U}_{\zeta_{q}})\right)$, then for each $t\in[a,b]$ we have
@@
 Hence $\zeta$ is smooth. Moreover, if $t\in[a,b]$, then $\varepsilon_{t}\circ\zeta=\text{id}_{N}:N\to N$. ∎
+Following more general cases of manifolds of mappings, such as the case of $C^{\ell}$-maps (with $\ell\geq 0$) from a compact manifold (possibly with rough boundary) to a smooth manifold which admits local addition (see e.g.
+[3, 24]), we well study the tangent bundle of the manifold of absolute continuous functions.
 Remark 3.20
@@
 Then $\Theta_{N}(v)\in{\Gamma_{AC}(\eta)}$ for each $v\in T_{\eta}AC_{L^{p}}([a,b],N)$ and the map
-$\Theta_{\eta}:T_{\eta}AC_{L^{p}}([a,b],N)\to{\Gamma_{AC}(\eta)},\quad v\mapsto\Theta_{\eta}(v):=\Theta_{N}|_{T_{\eta}AC_{L^{p}}([a,b],N)}(v)$
+$\Theta_{\eta}:T_{\eta}AC_{L^{p}}([a,b],N)\to{\Gamma_{AC}(\eta)},\quad v\mapsto\Theta_{\eta}(v):=\Theta_{N}(v)$
 is an isomorphism of topological vector spaces.
@@
 Since ${\Gamma_{AC}(\eta)}$ is a vector space, we identify its tangent bundle with ${\Gamma_{AC}(\eta)}\times{\Gamma_{AC}(\eta)}$.
-Let $\Psi_{\eta}:\mathcal{V}_{\eta}\to\mathcal{U}_{\eta}$ be a chart on $\eta$ such that $\Psi_{\eta}(0)=\eta$, then
+Let $\Psi_{\eta}:\mathcal{V}_{\eta}\to\mathcal{U}_{\eta}$ be a chart around $\eta$ such that $\Psi_{\eta}(0)=\eta$, then
 $T\Psi_{\eta}:T\mathcal{V}_{\eta}\simeq\mathcal{V}_{\eta}\times{\Gamma_{AC}(\eta)}\to TAC_{L^{p}}([a,b],N)$
@@
 $\Theta_{\eta}(v):[a,b]\to TN,\quad t\mapsto\Theta_{N}(v)(t)=\sigma(t)\in T_{\eta(t)}N$
-is absolutely continuous with $\pi_{TN}\circ\Theta_{\eta}(v)=\eta$, making the map $\Theta_{\eta}$ an isomorphism of topological vector spaces. ∎
+is in $AC_{L^{p}}([a,b],TN)$ and verifies $\pi_{TN}\circ\Theta_{\eta}(v)=\eta$, making the map $\Theta_{\eta}$ an isomorphism of topological vector spaces. ∎
 Remark 3.22
 Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits local addition and $p\in[1,\infty]$.
-Since $TN$ admits local addition and the vector bundle $\pi_{TN}:TN\to N$ is smooth, the map
-$AC([a,b],\pi_{TN}):AC_{L^{p}}([a,b],TN)\to AC_{L^{p}}([a,b],N),\quad\tau\mapsto\pi_{TN}\circ\tau$
-is smooth. Moreover, if $\eta\in AC_{L^{p}}([a,b],N)$, then
-$AC([a,b],\pi_{TN})^{-1}(\{\eta\})={\Gamma_{AC}(\eta)}.$
-The following result follows the same steps as for the case of $C^{\ell}$-maps (with $\ell\geq 0$) from a compact manifold (possibly with rough boundary) to a smooth manifold which admits local addition [2, Theorem A.12].
+Since $TN$ admits local addition and the vector bundle $\pi_{TN}:TN\to N$ is smooth, by Proposition 3.8, the map
+$AC_{L^{p}}([a,b],\pi_{TN}):AC_{L^{p}}([a,b],TN)\to AC_{L^{p}}([a,b],N),\quad\tau\mapsto\pi_{TN}\circ\tau$
+is smooth. Moreover, for $\eta\in AC_{L^{p}}([a,b],N)$, we have
+$AC_{L^{p}}([a,b],\pi_{TN})^{-1}(\{\eta\})={\Gamma_{AC}(\eta)}.$
+Then $AC_{L^{p}}([a,b],\pi_{TN})$ defines a vector bundle on $AC_{L^{p}}([a,b],N)$.
+By Proposition 3.8, we can apply [3, Theorem A.12] for our particular case:
 Proposition 3.23
 Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ which admits a local addition, $p\in[1,\infty]$ and $\pi_{TN}:TN\to N$ its tangent bundle. Then the map
-$AC([a,b],\pi_{TN}):AC_{L^{p}}([a,b],TN)\to AC_{L^{p}}([a,b],N),\quad\tau\mapsto\pi_{TN}\circ\tau$
-is a smooth vector bundle with fiber ${\Gamma_{AC}(\eta)}$ over $\eta\in AC_{L^{p}}([a,b],N)$. Moreover, the map
 $\Theta_{N}:TAC_{L^{p}}([a,b],N)\to AC_{L^{p}}([a,b],TN),\quad v\mapsto\Theta_{N}(v)$
@@
 Proposition 3.24
-Let $M$ and $N$ be a smooth manifolds modeled on sequentially complete locally convex spaces which admits a local addition and $p\in[1,\infty]$. If $f:M\to N$ is a $C^{3}$-map, then the tangent map of
+Let $M$ and $N$ be a smooth manifolds modeled on sequentially complete locally convex spaces which admits a local addition and $p\in[1,\infty]$. If $f:M\to N$ is a smooth map, then the tangent map of
 $AC_{L^{p}}([a,b],f):AC_{L^{p}}([a,b],M)\to AC_{L^{p}}([a,b],N),\quad\eta\mapsto f\circ\eta$
@@
 Proof.
-By Proposition 4.8, since $f$ is $C^{3}$ we know that $AC_{L^{p}}([a,b],f)$ is $C^{1}$, thus $TAC_{L^{p}}([a,b],f)$ exists.
-Let consider the local addition $\Sigma_{M}:\Omega_{M}\to M$ and $\eta\in AC_{L^{p}}([a,b],M)$.
-If $\Psi_{\eta}:\mathcal{V}_{\eta}\to\mathcal{U}_{\eta}$ is a chart on $\eta$ such that $\Psi_{\eta}(0)=\eta$. We consider the isomorphism of vector space
+By Proposition 3.8, we know that $AC_{L^{p}}([a,b],f)$ is smooth, thus $TAC_{L^{p}}([a,b],f)$ makes sense.
+Let us consider the local addition $\Sigma_{M}:\Omega_{M}\to M$ and $\eta\in AC_{L^{p}}([a,b],M)$.
+If $\Psi_{\eta}:\mathcal{V}_{\eta}\to\mathcal{U}_{\eta}$ is a chart around $\eta$ such that $\Psi_{\eta}(0)=\eta$. We consider the isomorphism of vector space
 $T\Psi_{\eta}:\{0\}\times{\Gamma_{AC}(\eta)}\to T_{\eta}AC_{L^{p}}([a,b],M).$
 For $t\in[a,b]$ we denote the point evaluation in $M$ and $N$ as $\varepsilon_{t}^{M}$ and $\varepsilon_{t}^{N}$ respectively, then for each $\sigma\in{\Gamma_{AC}(\eta)}$ we have
-$T\varepsilon_{t}^{N}\circ TAC_{L^{p}}([a,b],f)\circ T\Psi_{\eta}(0,\sigma)$ $=T\varepsilon_{t}\circ TAC_{L^{p}}([a,b],f)\circ T\Psi_{\eta}([s\mapsto s\sigma])$
+$T\varepsilon_{t}^{N}\circ TAC_{L^{p}}([a,b],f)\circ T\Psi_{\eta}(0,\sigma)$ $=T\varepsilon_{t}^{N}\circ TAC_{L^{p}}([a,b],f)\circ T\Psi_{\eta}([s\mapsto s\sigma])$
 $=T\varepsilon_{t}^{N}\circ TAC_{L^{p}}([a,b],f)([s\mapsto\Sigma_{M}(s\sigma)])$
@@
 $=[s\mapsto f\circ\Sigma_{M}(s\sigma(t))]$
-$=Tf\circ T_{0}\Sigma_{M}|_{T_{\eta(t)}N}([s\sigma(t)])$
+$=Tf\circ T_{0}\Sigma_{M}|_{T_{\eta(t)}M}([s\sigma(t)])$
 $=Tf([s\mapsto s\sigma(t)])$
-$=AC_{L^{p}}([a,b],Tf)(\sigma(t))$
+$=Tf(\sigma(t))$
 $=AC_{L^{p}}([a,b],Tf)\circ T\varepsilon_{t}^{M}\circ T\Psi_{\eta}(0,\sigma).$
@@
 Let $p\in[1,\infty]$.
-If $G$ is a Lie group modeled in a sequentially locally convex space $E$, then we already know that the space $AC_{L^{p}}([a,b],G)$ is a Lie group with Lie algebra given by $AC_{L^{p}}([a,b],T_{e}G)$ (see [10, 24]).
+If $G$ is a Lie group modeled in a sequentially locally convex space $E$, then we already know that the space $AC_{L^{p}}([a,b],G)$ is a Lie group (see [14, 33]).
 We will give an alternative proof of this.
-Let $e\in G$ be the neutral element, let $L_{g}:G\to G$, $h\mapsto gh$ be the left translation by $g\in G$ and the action
-$G\times TG\to TG,\quad(g,v_{h})\mapsto g.v_{h}:=TL_{g}(v_{h})\in T_{gh}G.$
+Let $e\in G$ be the neutral element, let $L_{g}:G\to G$, $h\mapsto gh$ be the left translation by $g\in G$ and the left action
+$G\times TG\to TG,\quad(g,v_{h})\mapsto g.v:=TL_{g}(v).$
 If $\varphi:U\subseteq G\to V\subseteq T_{e}G$ is a chart in $e$ such that $\varphi(e)=0$, then the set
@@
 $\Sigma_{\varphi}:\Omega_{\varphi}\to G,\quad v\mapsto\pi_{TG}(v)\left(\varphi^{-1}(\pi_{TG}(v)^{-1}.v)\right)$
-defines a local addition for $G$ (see e.g. [15]); hence $AC_{L^{p}}([a,b],G)$ is a smooth manifold with charts constructed with $(\Omega_{\varphi},\Sigma_{\varphi})$.
+defines a local addition for $G$ (see e.g. [23]); hence $AC_{L^{p}}([a,b],G)$ is a smooth manifold with charts constructed with the local adition $(\Omega_{\varphi},\Sigma_{\varphi})$.
 Let $\mu_{G}:G\times G\to G$ and $\lambda_{G}:G\to G$ be the multiplication map and inversion maps on $G$ respectively, we define the multiplication map $\mu_{AC}$ and the inversion map $\lambda_{AC}$ on $AC_{L^{p}}([a,b],G)$ as
@@
 $=\eta.$
-This enable us to say that for the neutral element ${\zeta_{e}}\in AC_{L^{p}}([a,b],G)$ the chart is given by
+This enables us to say that for the neutral element ${\zeta_{e}}\in AC_{L^{p}}([a,b],G)$ we have a chart given by
 $AC_{L^{p}}([a,b],\varphi):AC_{L^{p}}([a,b],U)\to AC([a,b],V),\quad\eta\mapsto\varphi\circ\eta.$
@@
 Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space. We define the right action
-$TG\times G\to TG,\quad(v_{h},g)\mapsto v_{h}.g:=T\rho_{g}(v).$
-And its restriction
-$T_{e}G\times G\to TG,\quad(v,g)\mapsto v.g:=T\rho_{g}(v).$
-Unlike on Lie groups, on half-Lie groups this actions may not be smooth. Hence we can not construct a local addition using a convenient chart on the identity (see Example 4.25).
+$TG\times G\to TG,\quad(v,g)\mapsto v.g:=T\rho_{g}(v)$ (4.1)
+and consider its restriction
+$T_{e}G\times G\to TG,\quad(v,g)\mapsto v.g:=T\rho_{g}(v).$ (4.2)
+Unlike on Lie groups, on half-Lie groups the latter action may not be smooth.
 The following proposition is direct application of Proposition 3.8.
@@
 $\eta.g(t):=\eta(t)g,\quad\text{for all }t\in[a,b].$
-Then $\eta.g\in AC_{L^{p}}([a,b],G)$ for each $g\in G$, the map
+Then $\eta.g\in AC_{L^{p}}([a,b],G)$ for each $g\in G$ and the map
 $AC_{L^{p}}([a,b],\rho_{g}):AC_{L^{p}}([a,b],G)\to AC_{L^{p}}([a,b],G),\quad\eta\mapsto\eta.g$
@@
 would imply the smoothness of the multiplication map on $G$.
-In fact, since the point evaluation map $\varepsilon_{a}:AC_{L^{p}}([a,b],G)\to G$, $\eta\mapsto\eta(a)$ and the map $\zeta:G\to AC_{L^{p}}([a,b],G)$, $g\mapsto[t\mapsto g]$, are smooth, the multiplication map on $G$ would be smooth as it coincides with the composition
+In fact, since the point evaluation map $\varepsilon_{a}:AC_{L^{p}}([a,b],G)\to G$, $\eta\mapsto\eta(a)$ and the map $\zeta:G\to AC_{L^{p}}([a,b],G)$, $g\mapsto[t\mapsto g]$, are smooth (see [37]), the multiplication map on $G$ would be smooth as it coincides with the composition
 $\varepsilon_{a}\circ\mathcal{R}\circ(\zeta,\text{id}_{G}):G\times G\to G,\quad(h,g)\mapsto hg.$
 Definition 4.5
-Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space $E$ which admits local addition, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$.
-Let $P=\{t_{0},..,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ charts of $G$ such that verify the definition of absolutely continuous for $\eta$.
-For each $j\in\{1,...,n\}$ we denote
-$\eta_{j}:=\varphi_{j}\circ\eta_{[t_{i-1},t_{i}]}\in AC_{L^{p}}([t_{i-1},t_{i}],E).$
-Then $\eta_{j}^{\prime}\in L^{p}([t_{i-1},t_{i}],E)$. Let $\eta^{\prime}_{j}=[\gamma_{j}]$ with $\gamma_{j}\in\mathcal{L}^{p}([t_{i-1},t_{i}],E)$, we define $\gamma:[a,b]\to TG$ via
-$\gamma(t):=T(\varphi_{j})^{-1}(\eta_{j}(t),\gamma_{j}(t)),\text{ if }t\in[t_{i-1},t_{i}[\text{ with }j\in\{1,...,n\},$
-and $\gamma(b)=T(\varphi_{j})^{-1}(\eta_{j}(b),\gamma_{j}(b))$. Then $\gamma$ is measurable (see [10]) and we define
-$\dot{\eta}:=\gamma.$
+Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space $E$ which admits local addition, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],G)$.
+Let $\{t_{0},..,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ charts of $G$ such that verify the definition of absolute continuity for $\eta$.
+For each $i\in\{1,...,n\}$ we denote
+$\eta_{i}:=\varphi_{i}\circ\eta_{[t_{i-1},t_{i}]}\in AC_{L^{p}}([t_{i-1},t_{i}],E).$
+Then $\eta_{i}^{\prime}\in L^{p}([t_{i-1},t_{i}],E)$. Let $\eta^{\prime}_{i}=[\gamma_{i}]$ with $\gamma_{i}\in\mathcal{L}^{p}([t_{i-1},t_{i}],E)$. We define $\gamma:[a,b]\to TG$ via
+$\gamma(t):=T(\varphi_{i})^{-1}(\eta_{i}(t),\gamma_{i}(t)),\text{ if }t\in[t_{i-1},t_{i}[\text{ with }i\in\{1,...,n\},$
+and $\gamma(b)=T(\varphi_{n})^{-1}(\eta_{n}(b),\gamma_{n}(b))$. Then $\gamma$ is measurable (see [14]) and we define
+$\dot{\eta}:=[\gamma].$
 Definition 4.6
@@
 We say that $G$ is $L^{p}$-semiregular if for each $[\gamma]\in L^{p}([0,1],T_{e}G)$, there exists an $AC_{L^{p}}$-Carathéodory solution $\eta_{\gamma}\in AC_{L^{p}}([0,1],G)$ of the equation
-${\dot{\eta}}_{\gamma}=\gamma.\eta_{\gamma},\quad\eta_{\gamma}(0)=e$
+${\dot{y}}(t)$ $=\gamma(t).y(t),\quad t\in[0,1]$ (4.3)
+$y(0)$ $=e$ (4.4)
 such that the differential equation satisfies local uniqueness of Carathéodory solutions in the sense of [18]. In this case, we define the evolution map
-$\operatorname{Evol}:L^{p}([0,1],T_{e}G)\to AC_{L^{p}}([a,b],G),\quad[\gamma]\mapsto\operatorname{Evol}(\gamma):=\eta_{\gamma}.$
+$\operatorname{Evol}:L^{p}([0,1],T_{e}G)\to AC_{L^{p}}([a,b],G),\quad[\gamma]\mapsto\operatorname{Evol}(\gamma):=\eta_{\gamma}.$ (4.5)
 Additionally, if $G$ admits a local addition, we say that $G$ is $L^{p}$-regular if $G$ is $L^{p}$-semiregular and if the evolution map is smooth.
@@
 Definition 4.7
-Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space and $p\in[1,\infty]$.
+Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space and $p\in[1,\infty]$.
 We say that $G$ is locally $L^{p}$-semiregular if there is exists an open 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that for each $[\gamma]\in B$ there exists a $AC_{L^{p}}$-Carathéodory solution $\eta_{\gamma}\in AC_{L^{p}}([0,1],G)$ of the equation
-$\dot{\eta}_{\gamma}=\gamma.\eta_{\gamma},\quad\eta_{\gamma}(0)=e,$
+$\dot{y}(t)$ $=\gamma(t).y(t),\quad t\in[0,1]$
+$y(0)$ $=e$
 and the latter differential equation satisfies local uniqueness of solutions.
-For our purpuse, we will use the subdivision property [24, Lemma 2.17].
+For our purpose, we will use the subdivision property [33, Lemma 2.17].
 Lemma 4.8
@@
 Then $[\gamma_{n,k}]\in L^{p}([0,1],E)$. Moreover, for each $[\gamma]\in L^{p}([0,1],E)$ and continuous seminorm $q$ on $L^{p}([0,1],E)$, we have
-$\sup_{k\in\{1,...,n-1\}}q(\gamma_{n,k})\to 0,\quad\text{as }n\to\infty.$
+$\sup_{k\in\{0,...,n-1\}}q(\gamma_{n,k})\to 0,\quad\text{as }n\to\infty.$
 Lemma 4.9
-Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space which admits local addition and $p\in[1,\infty]$.
+Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space which admits local addition and $p\in[1,\infty]$.
 Then $G$ is locally $L^{p}$-semiregular if and only if $G$ is $L^{p}$-semiregular.
@@
 Let $G$ be locally $L^{p}$-semiregular.
-Then there is exists and 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that $\operatorname{Evol}(\beta):=\eta_{\beta}\in AC_{L^{p}}([a,b],G)$ exists of each $[\beta]\in B$.
-without loss of generalization we assume that
+Then there exists an 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that $\operatorname{Evol}(\beta)\in AC_{L^{p}}([a,b],G)$ exists for each $[\beta]\in B$.
+Without loss of generality we assume that
 $B=\{[\beta]\in L^{p}([0,1],T_{e}G):\lVert[\beta]\lVert_{L^{p},q}<1\}$
@@
 $\alpha_{k}(W_{\gamma})\subseteq B,\quad\text{for all }k\in\{0,1,...,n-1\}.$
-For $[\beta]\in W_{\gamma}$ and $j\in\{0,...,n-1\}$ we denote $\eta_{j}=\operatorname{Evol}(\beta_{n,j})$ and we define the function $\eta_{\beta}:[0,1]\to G$ via
+For $[\beta]\in W_{\gamma}$ and $j\in\{0,...,n-1\}$ we write $\eta_{j}=\operatorname{Evol}(\beta_{n,j})$ and we define the function $\eta_{\beta}:[0,1]\to G$ via
 $\eta_{\beta}(t):=\eta_{0}(nt),\quad\text{ if }t\in[0,1/n]$
@@
 Definition 4.10
-Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space which admits local addition and $p\in[1,\infty]$.
-We say that $G$ is locally $L^{p}$-regular if $G$ is $L^{p}$-semiregular and there is exists an 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that its restricted evolution map $\operatorname{Evol}|_{B}$ is smooth.
-The following Lemma is just an application of [23, Proposition 4.11] to our case.
+Let $G$ be a right half-Lie group modeled on a sequentially complete locally convex space which admits local addition and $p\in[1,\infty]$.
+We say that $G$ is locally $L^{p}$-regular if $G$ is $L^{p}$-semiregular and there is exists a 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that its restricted evolution map $\operatorname{Evol}|_{B}$ is smooth.
+The following lemma is just an application of [14, Proposition 5.25] or [32, Proposition 4.11] to our case.
 Lemma 4.11
-Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space which admits local addition and $p\in[1,\infty]$. We consider the evolution map with continuous values
+Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space which admit a local addition and $p\in[1,\infty]$. We consider the evolution map with continuous values
 $\operatorname{Evol}_{C}:L^{p}([0,1],T_{e}G)\to C([0,1],G),\quad[\gamma]\mapsto\operatorname{Evol}_{C}(\gamma):=\eta_{\gamma}.$
-Then, $\operatorname{Evol}_{C}$ is continuous if and only if there is exists an 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that its restricted evolution map $\operatorname{Evol}_{C}|_{B}$ is continuous.
+Then, $\operatorname{Evol}_{C}$ is continuous if and only if there exists a 00-neighborhood $B$ of $L^{p}([0,1],T_{e}G)$ such that the restricted evolution map $\operatorname{Evol}_{C}|_{B}$ is continuous.
 Proof.
@@
 is continuous.
-Following Lemma 5.9, for each $[\beta]\in W_{\gamma}$, we have that the construction of the function $\operatorname{Evol}(\beta)$ implies that $Evol_{C}$ is just the product of composition of continuous maps.
+Following Lemma 4.9, if $[\gamma]\in L^{p}([0,1],T_{e}G)$, for each $[\beta]\in W_{\gamma}$, we have that the construction $\operatorname{Evol}_{C}(\beta)$ implies that $\left(\text{Evol}_{C}\right)|_{W_{\gamma}}$ is just the product of composition of continuous maps.
 ∎
 Lemma 4.12
-Let $E_{1},E_{2}$ and $F$ be locally convex spaces, $1\leq p<\infty$ and $U\subseteq E_{2}$ an open subset.
+Let $1\leq p<\infty$, $E_{1}$ be a Fréchet space, $E_{2}$ and $F$ be locally convex spaces and $U\subseteq E_{2}$ an open subset.
 If $r\in\mathbb{N}\cup\{0,\infty\}$ and $f:E_{1}\times U\to F$ is a $C^{r}$-map such that for each $y\in U$ the map $f(\cdot,y):E_{1}\to F$ is linear, then the map
@@
 $\leq\lVert\gamma_{0}\lVert_{L^{p},\kappa}.$
-Hence $\tilde{f}(\gamma_{0},\eta_{0})\in L^{p}([0,1],F)$. Since the space $C([0,1],E_{1})$ is dense in $L^{p}([0,1],E_{1})$, there exists a $\gamma_{c}\in C([0,1],E_{1})$ such that
-$\lVert\gamma_{c}-\gamma_{0}\lVert_{L^{p},p}\leq 2\varepsilon/5.$
+Hence $\tilde{f}(\gamma_{0},\eta_{0})\in L^{p}([0,1],F)$.
+Since the space $C([0,1],E_{1})$ is dense in $L^{p}([0,1],E_{1})$ (see [14, Lemma A.4]) there exists a $\gamma_{c}\in C([0,1],E_{1})$ such that
+$\lVert\gamma_{c}-\gamma_{0}\lVert_{L^{p},\kappa}\leq\varepsilon/5.$
 Additionally, since the maps
@@
 $\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{c},\eta_{0})\lVert_{L^{p},\beta}\leq 2\varepsilon/5,\quad\text{for all }\eta\in W.$
-With this, for each $\eta\in W$ and $\gamma\in L^{p}([0,1],E_{1})$ with $\lVert\gamma-\gamma_{0}\lVert_{L^{p},\kappa}\leq 2\varepsilon/5$ we estimate
-$\lVert\tilde{f}(\gamma,\eta)-\tilde{f}(\gamma_{0},\eta_{0})\lVert_{L^{p},\beta}$ $\leq\lVert\tilde{f}(\gamma-\gamma_{c},\eta)\lVert_{L^{p},\beta}+\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{0},\eta_{0})\lVert_{L^{p},\beta}$
-$\leq\lVert\tilde{f}(\gamma-\gamma_{c},\eta)\lVert_{L^{p},\beta}+\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{c},\eta_{0})\lVert_{L^{p},\beta}+\lVert\tilde{f}(\gamma_{c}-\gamma_{0},\eta_{0})\lVert_{L^{p},\beta}$
-$\leq\lVert\gamma-\gamma_{c}\lVert_{L^{p},\kappa}+\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{c},\eta_{0})\lVert_{L^{p},\beta}+\lVert\gamma_{c}-\gamma_{0}\lVert_{L^{p},\kappa}.$
-Since $\lVert\gamma_{c}-\gamma_{0}\lVert_{L^{p},\kappa}\leq 2\varepsilon/10$ and $\lVert\gamma-\gamma_{0}\lVert_{L^{p},\kappa}\leq 2\varepsilon/10$ we have
-$\lVert\gamma-\gamma_{c}\lVert_{L^{p},\kappa}\leq 2\varepsilon/5$
-and
-$\lVert\tilde{f}(\gamma,\eta)-\tilde{f}(\gamma_{0},\eta_{0})\lVert_{L^{p},\beta}\leq 2\varepsilon/5+2\varepsilon/5+2\varepsilon/10=\varepsilon$
-Thus $\tilde{f}$ is continuous.
-By linearity in the first variable the function $\overline{f}$ has a continuous differential $d_{1}\tilde{f}:L^{p}([0,1],E_{1})\times L^{p}([0,1],E_{1})\times C([0,1],E_{2})\to L^{p}([0,1],F)$ in the first variable.
-Let consider the map $f$ as a $C^{1}$-map.
-If $x\in E_{1}$, $y,y_{1}\in E_{2}$ and $t\in\mathbb{R}^{\times}$, then we have
+With this, for each $\eta\in W$ and $\gamma\in L^{p}([0,1],E_{1})$ with $\lVert\gamma-\gamma_{0}\lVert_{L^{p},\kappa}\leq\varepsilon/5$ we estimate
+$\left\lVert\tilde{f}(\gamma,\eta)-\tilde{f}(\gamma_{0},\eta_{0})\right\lVert_{L^{p},\beta}$ $\leq\left\lVert\tilde{f}(\gamma,\eta)-\tilde{f}(\gamma_{c},\eta)\right\lVert_{L^{p},\beta}+\left\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{0},\eta_{0})\right\lVert_{L^{p},\beta}$
+$=\left\lVert\tilde{f}(\gamma-\gamma_{c},\eta)\right\lVert_{L^{p},\beta}+\left\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{c},\eta_{0})\right\lVert_{L^{p},\beta}+\left\lVert\tilde{f}(\gamma_{c},\eta_{0})-\tilde{f}(\gamma_{0},\eta_{0})\right\lVert_{L^{p},\beta}$
+$\leq\left\lVert\tilde{f}(\gamma-\gamma_{c},\eta)\right\lVert_{L^{p},\beta}+\left\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{c},\eta_{0})\right\lVert_{L^{p},\beta}+\left\lVert\tilde{f}(\gamma_{c}-\gamma_{0},\eta_{0})\right\lVert_{L^{p},\beta}$
+$\leq\Big{\lVert}\gamma-\gamma_{c}\Big{\lVert}_{L^{p},\kappa}+\left\lVert\tilde{f}(\gamma_{c},\eta)-\tilde{f}(\gamma_{c},\eta_{0})\right\lVert_{L^{p},\beta}+\Big{\lVert}\gamma_{c}-\gamma_{0}\Big{\lVert}_{L^{p},\kappa}.$
+Since $\lVert\gamma_{c}-\gamma_{0}\lVert_{L^{p},\kappa}\leq\varepsilon/5$ and $\lVert\gamma-\gamma_{0}\lVert_{L^{p},\kappa}\leq\varepsilon/5$ we have
+$\lVert\gamma-\gamma_{c}\lVert_{L^{p},\kappa}\leq 2\varepsilon/5.$
+Thus
+$\lVert\tilde{f}(\gamma,\eta)-\tilde{f}(\gamma_{0},\eta_{0})\lVert_{L^{p},\beta}\leq 2\varepsilon/5+2\varepsilon/5+\varepsilon/5=\varepsilon.$
+Thus $\tilde{f}$ is continuous. By linearity in the first variable, the map $\overline{f}$ has a continuous differential in the first variable
+$d_{1}\tilde{f}:L^{p}([0,1],E_{1})\times L^{p}([0,1],E_{1})\times C([0,1],E_{2})\to L^{p}([0,1],F).$
+Let consider the map $f$ as a $C^{1}$-map. If $x\in E_{1}$, $y,y_{1}\in E_{2}$ and $t\in\mathbb{R}^{\times}$, then we have
 […]
@@
 $\widetilde{d_{2}f}:L^{p}([0,1],E_{1})\times C([0,1],U)\times C([0,1],E_{2})\to L^{p}([0,1],F),\quad(\gamma,\eta,\eta_{1})\mapsto d_{2}f(\gamma,\eta,\eta_{1})$
-is continuous. For $t\in[0,1]$ we denote
+is continuous. For $t\in[0,1]$ fixed, we denote
 $g_{t}:C([0,1],F)\to F,\quad\gamma\mapsto\gamma(t).$
@@
 $d_{2}\tilde{f}(\gamma,\eta,\eta_{1})=\widetilde{d_{2}f}(\gamma,\eta,\eta_{1}).$
-Hence the continuity of $\widetilde{d_{2}f}$ implies that the map $\overline{f}$ is $C^{1}$. Proceeding by induction, if $f$ is a $C^{k}$-map, since $\tilde{f}$ is linear in the first variable we have
+Hence the continuity of $\widetilde{d_{2}f}$ implies that the map $\widetilde{f}$ is $C^{1}$.
+Proceeding by induction, if $f$ is a $C^{k}$-map, since $\tilde{f}$ is linear in the first variable we have
 $d\tilde{f}(\gamma,\eta,\gamma_{1},\eta_{1})$ $=d_{1}\tilde{f}(\gamma,\gamma_{1},\eta)+d_{2}\tilde{f}(\gamma,\eta,\eta_{1})$
@@
 By the induction hypothesis $\tilde{f}$ and $d_{2}f$ are $C^{r-1}$, hence $\widetilde{d_{2}f}$ is $C^{r-1}$ with $\widetilde{d_{2}f}=d_{2}\tilde{f}$, thus $\tilde{f}$ is $C^{r}$. ∎
-For the case $L_{rc}^{\infty}$ we recall [10, Proposition 2.3].
+For the case $L_{rc}^{\infty}$ we recall [14, Proposition 2.3].
 Lemma 4.13
@@
 Theorem 4.14
-Let $G$ be a right half-Lie group modeled in a Fréchet space $E$ which admits local addition and $1\leq p<\infty$. Let $G$ be $L^{p}$-semiregular with continuous evolution map
-$\operatorname{Evol}_{C}:L^{p}([0,1],T_{e}G),\to C([0,1],G),\quad\gamma\mapsto\operatorname{Evol}_{C}(\gamma).$
+Let $G$ be a right half-Lie group modeled in a Fréchet space $E$ which admits a local addition and $1\leq p<\infty$. Let $G$ be $L^{p}$-semiregular with continuous evolution map
+$\operatorname{Evol}_{C}:L^{p}([0,1],T_{e}G),\to C([0,1],G)\quad\gamma\mapsto\operatorname{Evol}_{C}(\gamma).$
 If the restriction of the right action
@@
 $\operatorname{Evol}:L^{p}([0,1],T_{e}G),\to AC_{L^{p}}([0,1],G),\quad\gamma\mapsto\operatorname{Evol}(\gamma)$
-is continuous.
-If we replace $L^{p}$ with $L_{rc}^{\infty}$ the result remains valid.
+is continuous. If $G$ is a right half-Lie group modeled in an integral complete locally convex space $E$, then if we replace $L^{p}$ with $L_{rc}^{\infty}$ the result remains valid.
 Proof.
@@
 $f_{i}:T_{e}G\times V_{i}\to E,\quad(v,x)\mapsto d\varphi_{i}\circ\tau(v,\varphi_{i}^{-1}(x)).$
-Then each $f_{i}$ is continuous and for each $x\in V_{i}$ the function $f_{i}(\cdot,x):T_{e}G\to E$, $v\mapsto f_{i}(v,x)$ is linear. Hence by Lemma 5.12 the map
+Then each $f_{i}$ is continuous and for each $x\in V_{i}$ the function
+$f_{i}(\cdot,x):T_{e}G\to E,\quad v\mapsto f_{i}(v,x)$
+is linear. Hence by Lemma 4.12 the map
 $\tilde{f}_{i}:L^{p}([t_{i-1},t_{i}],T_{e}G)\times C([t_{i-1},t_{i}],V_{i})\to L^{p}([t_{i-1},t_{i}],E),\quad([\beta],\eta)\mapsto f_{i}\circ([\beta],\eta)$
 is continuous. We denote
-$F:\prod_{i=1}^{n}L^{p}([t_{i-1},t_{i}],T_{e}G)\times C([t_{i-1},t_{i}],V_{i})\to\prod_{i=1}^{n}L^{p}([t_{i-1},t_{i}],E),\quad\left([\beta_{i}],\eta_{i}\right)_{i=1}^{n}\mapsto\left(f_{i}\circ([\beta_{i}],\eta_{i}\right)_{i=1}^{n}.$
+$F:\prod_{i=1}^{n}L^{p}([t_{i-1},t_{i}],T_{e}G)\times C([t_{i-1},t_{i}],V_{i})\to\prod_{i=1}^{n}L^{p}([t_{i-1},t_{i}],E),\quad\left([\beta_{i}],\eta_{i}\right)_{i=1}^{n}\mapsto\left(f_{i}\circ([\beta_{i}],\eta_{i})\right)_{i=1}^{n}.$
 Then $F$ is continuous, where for each $i\in\{1,...,n\}$ we have
@@
 $\Phi_{4}:=\left(\Psi|^{\text{Im}(\Psi)}\right)^{-1}\circ\Phi_{3}:\mathcal{Q}_{[\gamma]}\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],V_{i}),\quad[\beta]\mapsto\left(\varphi_{i}\circ\operatorname{Evol}(\beta)|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}.$
-Let denote the homeomorphism onto its image (see Proposition 3.9)
+Let denote map
 $\Phi_{5}:AC_{L^{p}}([0,1],G)\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],G),\quad\eta\mapsto(\eta|_{[t_{i-1},t_{i}]})_{i=1}^{n}.$
-and the homeomorphism
+which is a homeomorphism onto its image the homeomorphism onto its image (see [37, Proposition 4.13]) and the homeomorphism
 $\Phi_{6}:\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],V_{i})\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],U_{i}),\quad\tau_{i}\mapsto(\varphi_{i}^{-1}\circ\tau_{i})_{i=1}^{n}.$
@@
 Let $n,m,r\in\mathbb{N}$.
-We set the Fréchet space $C(\mathbb{R}^{n},\mathbb{R}^{m})$ of continuous functions $\phi:\mathbb{R}^{n}\to\mathbb{R}^{m}$, whose topology is generated by the famility of seminorms
+We consider the Fréchet space $C(\mathbb{R}^{n},\mathbb{R}^{m})$ of continuous functions $\phi:\mathbb{R}^{n}\to\mathbb{R}^{m}$, whose topology is generated by the famility of seminorms
 $\lVert\cdot\lVert_{L}:C(\mathbb{R}^{n},\mathbb{R}^{m})\to[0,\infty),\quad\phi\mapsto\sup_{x\in L}\lvert\phi(x)\lvert$
@@
 for $\alpha\in\mathbb{N}_{0}^{n}$ with $\lvert\alpha\lvert\leq r$. Let $K\subseteq\mathbb{R}^{n}$ be a non empty compact subset, we define the Banach space of $C^{r}$-maps supported in $K$ as
-$C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{m}):=\{\phi\in C^{r}(\mathbb{R}^{n},\mathbb{R}^{m}):\phi|_{\mathbb{R}^{n}\setminus K}=0\}$
+$C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{m}):=\{\phi\in C^{r}(\mathbb{R}^{n},\mathbb{R}^{m}):\phi|_{\mathbb{R}^{n}\setminus K}=0\}$ (5.1)
 endowed with the induced topology.
 We denote derivative map
-$D:C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))\to C_{K}^{r-1}(\mathbb{R}^{n},\mathbb{R}^{n\times n}),\quad f\mapsto Df:=f^{\prime}$
+$D:C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\to C_{K}^{r-1}(\mathbb{R}^{n},\mathbb{R}^{n\times n}),\quad f\mapsto Df:=f^{\prime}$ (5.2)
 where $f^{\prime}(x)$ is the Jacobian matrix of $f$.
@@
 The set $\text{Diff}^{r}(\mathbb{R}^{n})$ is a group under the composition and we define the subgroup $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$ of $C^{r}$-diffeomorphisms which are supported in a compact set $K\subset\mathbb{R}^{n}$ as
-$\text{Diff}_{K}^{r}(\mathbb{R}^{n}):=\{\phi\in\text{Diff}^{r}(\mathbb{R}^{n}):\phi-\text{id}_{\mathbb{R}^{n}}\in C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\}.$
+$\text{Diff}_{K}^{r}(\mathbb{R}^{n}):=\{\phi\in\text{Diff}^{r}(\mathbb{R}^{n}):\phi-\text{id}_{\mathbb{R}^{n}}\in C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\}.$ (5.3)
 Let $\mathcal{V}_{K}:=\{\phi-\text{id}_{\mathbb{R}^{n}}\in C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}):\phi\in\text{Diff}_{K}^{r}(\mathbb{R}^{n})\}$, then $\mathcal{V}_{K}$ is open in $C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$ and the map
-$\Phi:\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})\to\mathcal{V}_{K},\quad\phi\mapsto\phi-\text{id}_{\mathbb{R}^{n}}$
-is a global chart for $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$, turning it into a right half-Lie group modeled on the Banach space $C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$ (See [12, Proposition 14.6]).
+$\Phi:\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})\to\mathcal{V}_{K},\quad\phi\mapsto\phi-\text{id}_{\mathbb{R}^{n}}$ (5.4)
+is a global chart for $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$, turning it into a right half-Lie group modeled on the Banach space $C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$ (See [16, Proposition 14.6]).
 On the set $\mathcal{V}_{K}$ we define a group multiplication by
@@
 Identifying $T_{0}\mathcal{V}_{K}$ with $\{0\}\times C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$, the restricted right action is given by
-($\left(\{0\}\times C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\right)\times\mathcal{V}_{K}\to T\mathcal{V}_{K},\quad((0,\varphi),\psi)\mapsto d\rho_{\psi}(0,\varphi)=\varphi\circ(id_{\mathbb{R}^{n}}+\psi)$
+($\left(\{0\}\times C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\right)\times\mathcal{V}_{K}\to T\mathcal{V}_{K},\quad((0,\varphi),\psi)\mapsto d\rho_{\psi}(0,\varphi)=\varphi\circ(id_{\mathbb{R}^{n}}+\psi)$ (5.5)
 Hence the right action is continuous (see e.g. [15]).
@@
 […]
-Since the evaluation map $\varepsilon:C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ is $C^{r,\infty}$, by Lemma 5.12 the map
+Since the evaluation map $\varepsilon:C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ is $C^{\infty,r}$ in the sense of [1] and hence $C^{r}$. By Lemma 4.12 the map
 $\Phi_{n}:L^{p}([0,1],C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n}))\times C([0,1],\mathbb{R}^{n})\to L^{p}([0,1],\mathbb{R}^{n}),\quad\Phi_{n}([\gamma],\zeta):=\varepsilon\circ([\gamma],\zeta)$
@@
 […]
-for $t\in[0,1]$. Then $T$ is $C^{\infty,\infty,r}$.
+for $t\in[0,1]$. Then $T$ is $C^{\infty,\infty,r}$ in the sense of [2].
 Let $[\gamma]\in B_{L}$ and $x\in\mathbb{R}^{n}$ be fixed, for $\zeta_{1},\zeta_{2}\in C([0,1],\mathbb{R}^{n})$, by the Mean Value Theorem we have
@@
 $\leq L\lVert\zeta_{2}-\zeta_{1}\lVert_{\infty}.$
-Hence $T([\gamma],x,\cdot)$ is a $C^{r}$-map and a $L$-contraction. By Banach’s Fixed Point Theorem, there exists $\zeta_{[\gamma],x}\in C([0,1],\mathbb{R}^{n})$ such that
+Hence $T([\gamma],x,\cdot)$ is a $C^{r}$-map and an $L$-contraction. By Banach’s Fixed Point Theorem, there exists $\zeta_{[\gamma],x}\in C([0,1],\mathbb{R}^{n})$ such that
 $T([\gamma],x,\zeta_{[\gamma],x})=\zeta_{[\gamma],x}.$
-By [10, Lemma 6.2] the map
+By [14, Lemma 6.2] the map
 $F_{1}:B_{L}\times\mathbb{R}^{n}\to C([0,1],\mathbb{R}^{n}),\quad([\gamma],x)\mapsto\zeta_{[\gamma],x}$
@@
 We define
-$S:B\to C([0,1],C^{r}(\mathbb{R}^{n},\mathbb{R}^{n})),\quad[\gamma]\mapsto F_{4}([\gamma])-\text{I}.$
+$S:B_{L}\to C([0,1],C^{r}(\mathbb{R}^{n},\mathbb{R}^{n})),\quad[\gamma]\mapsto F_{4}([\gamma])-\text{I}.$
 Then for each $t\in[0,1]$ we have $S([\gamma])(t)\in C_{K}^{r}(\mathbb{R}^{n},\mathbb{R}^{n})$.
@@
 $\Psi_{\phi}^{-1}:\mathcal{V}_{\phi}\to\mathcal{U}_{\phi},\quad X\mapsto\exp\circ X,$
-turning $\text{Diff}^{\text{ }r}(M)$ in a Banach manifold.
+turning $\text{Diff}^{\text{ }r}(M)$ into a Banach manifold.
 Moreover, under the composition the manifold $\text{Diff}^{\text{ }r}(M)$ becomes a right half-Lie group (see e.g.
@@
 The restricted right action
-$\mathcal{X}^{r}(M)\times\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})\to T\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n}),\quad(X,\phi)\mapsto X.\phi=X\circ\phi$
+$\mathcal{X}^{r}(M)\times\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})\to T\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n}),\quad(X,\phi)\mapsto X.\phi=X\circ\phi$ (6.1)
 is continuous (see e.g. [15]).
 Let $[\gamma]\in L^{p}([0,1],\mathcal{X}^{r}(M))$, we will study the equation
-$\dot{\eta}=\gamma.\eta,\quad\eta(0)=\text{id}_{M}.$
+$\dot{\eta}(t)$ $=\gamma(t).\eta(t),\quad t\in[0,1]$ (6.2)
+$\eta(0)$ $=\text{id}_{M}.$ (6.3)
 Definition 6.1
 Let $(E,\lVert\cdot\lVert_{E})$ be a normed space and $U\subseteq E$ be a subset.
-We say that a function $f:[a,b]\times U\to E$ satisfies a $L^{1}$-Lipschitz condition if there exists a measurable function $g\in L^{1}([a,b],\mathbb{R})$ such that
-$\text{Lip}(f(t,\cdot))\leq g(t),\quad t\in[a,b]$
-Where $\text{Lip}(f(t,\cdot))\in[0,\infty]$ denote the infimum of all Lipschitz constant for $f(t,\cdot):U\to E$.
+We say that a function $f:[a,b]\times U\to E$ satisfies an $L^{1}$-Lipschitz condition if there exists a measurable function $g\in L^{1}([a,b],\mathbb{R})$ such that
+$\text{Lip}(f(t,\cdot))\leq g(t),\quad t\in[a,b],$
+where $\text{Lip}(f(t,\cdot))\in[0,\infty]$ denote the infimum of all Lipschitz constants for $f(t,\cdot):U\to E$.
 Definition 6.2
@@
 $y^{\prime}=f(t,y)$
-satisfying $\tau(t_{0})=\eta(t_{0})$ for some $t_{0}\in[a,b]$, then $\tau=\eta$ [10, Proposition 10.5]. If the initial value problem
+satisfying $\tau(t_{0})=\eta(t_{0})$ for some $t_{0}\in[a,b]$, then $\tau=\eta$ [14, Proposition 10.5]. If the initial value problem
 $y^{\prime}(t)$ $=f(t,y(t)),\quad t\in[a,b]$
@@
 for all $t\in[0,1]$ and $q\in M$.
-The following result follows the same steps as the proof of the $L^{1}$-regularity of $\text{Diff}_{c}^{\infty}(M)$ proved in [10, Section 11].
+The following result follows the same steps as the proof of the $L^{1}$-regularity of $\text{Diff}_{c}^{\infty}(M)$ proved in [14, Section 11].
 We recommend to the reader to read the reference to see some steps in detail.
@@
 For $X\in\Gamma_{C^{r}}(\phi)$ we write
-$X_{i}:=d\psi_{i}\circ X\circ\kappa_{i}^{-1}:B_{5}(0)\to\mathbb{R}^{n}.$
+$X_{i}:=d\psi_{i}\circ X\circ\kappa_{i}^{-1}:B_{5}(0)\to\mathbb{R}^{m}.$
 Then, for each $\ell\in[1,5]$ the map
-$\rho_{\ell}:\Gamma_{C^{r}}(\phi)\to\prod_{i=1}^{n}C^{r}(B_{\ell}(0),\mathbb{R}^{n}),\quad X\mapsto\rho_{\ell}(X):=(X_{i}|_{B_{\ell}(0)})_{i=1}^{n}$
+$\rho_{\ell}:\Gamma_{C^{r}}(\phi)\to\prod_{i=1}^{n}C^{r}(B_{\ell}(0),\mathbb{R}^{m}),\quad X\mapsto\rho_{\ell}(X):=(X_{i}|_{B_{\ell}(0)})_{i=1}^{n}$
 is a linear topological embedding with closed image (see [3]).
@@
 Doing the corresponding identifications of product spaces, we have the linear topological embeddings with closed image
-$R_{5}:=L^{p}([0,1],\rho_{5}):L^{p}([0,1],\mathcal{X}^{r}(M))\to\prod_{i=1}^{n}L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{n})),\quad\gamma\mapsto\rho_{5}\circ\gamma$
+$R_{5}:=L^{p}([0,1],\rho_{5}):L^{p}([0,1],\mathcal{X}^{r}(M))\to\prod_{i=1}^{n}L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{m})),\quad\gamma\mapsto\rho_{5}\circ\gamma$
 and
-$R_{1}:=C([0,1],\rho_{1}):C([0,1],\mathcal{X}^{r}(M))\to\prod_{i=1}^{n}C([0,1],C^{r}(B_{1}(0),\mathbb{R}^{n})),\quad\gamma\mapsto\rho_{1}\circ\gamma.$
-Before study the equation for $[\gamma]\in L^{p}([0,1],\mathcal{X}^{r}(M))$, the embedding $R_{1}$ allow us to study it first for
-$[\gamma]\in L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{n})).$
-We consider the continuous seminorm $q:C^{r}(B_{5}(0),\mathbb{R}^{n})\to[0,\infty)$ by
+$R_{1}:=C([0,1],\rho_{1}):C([0,1],\mathcal{X}^{r}(M))\to\prod_{i=1}^{n}C([0,1],C^{r}(B_{1}(0),\mathbb{R}^{m})),\quad\gamma\mapsto\rho_{1}\circ\gamma.$
+Before we study the equation for $[\gamma]\in L^{p}([0,1],\mathcal{X}^{r}(M))$, the topological embedding $R_{1}$ allows us to study it first for
+$[\gamma]\in L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{m})).$
+We consider the continuous seminorm $q:C^{r}(B_{5}(0),\mathbb{R}^{m})\to[0,\infty)$ given by
 $q(\phi):=\sup_{x\in\overline{B}_{4}(0)}\left(\lVert\phi^{\prime}(x)\lVert_{op}+\lVert\phi(x)\lVert_{\infty}\right),\quad\text{for all }\phi\in C^{r}(B_{5}(0),\mathbb{R}^{n}).$
 For $10<L<$ 1 we set the open ball
-$B_{L}:=\{\gamma\in L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{n})):\lVert\gamma\lVert_{L^{p},q}<L\}.$
-Just like the case for $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$, we need a $AC_{L^{p}}$-Carathéodory solution $\zeta:[0,1]\to B_{4}(0)$ that verifies the integral equation
+$B_{L}:=\{\gamma\in L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{m})):\lVert\gamma\lVert_{L^{p},q}<L\}.$
+Just like the case for $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$, we need an $AC_{L^{p}}$-Carathéodory solution $\zeta:[0,1]\to B_{4}(0)$ that verifies the integral equation
 […]
@@
 $F:B_{L}\times B_{3}(0)\to C([0,1],B_{4}(0)),\quad([\gamma],x)\mapsto\zeta_{[\gamma],x}$
-is $C^{r}$ (see [10, Lemma 6.2]). By the exponential law, we have
+is $C^{r}$ (see [14, Lemma 6.2]). By the exponential law, we have
   • a)
@@
 Let consider the open set
-$\lfloor\overline{B_{2}(0)},B_{4}(0)\rfloor_{r}:=\{\varphi\in C^{r}(B_{3}(0),\mathbb{R}^{n}):\varphi(\overline{B_{2}(0)})\subseteq B_{4}(0)\}.$
+$\lfloor\overline{B_{2}(0)},B_{4}(0)\rfloor_{r}:=\{\varphi\in C^{r}(B_{3}(0),\mathbb{R}^{m}):\varphi(\overline{B_{2}(0)})\subseteq B_{4}(0)\}.$
 Then the map
-$S:C^{r}(B_{5}(0),\mathbb{R}^{n})\times\lfloor\overline{B_{2}(0)},B_{4}(0)\rfloor_{r}\to C^{r}(B_{2}(0),\mathbb{R}^{n}),\quad(\psi,\varphi)\mapsto\psi\circ\varphi$
+$S:C^{r}(B_{5}(0),\mathbb{R}^{n})\times\lfloor\overline{B_{2}(0)},B_{4}(0)\rfloor_{r}\to C^{r}(B_{2}(0),\mathbb{R}^{m}),\quad(\psi,\varphi)\mapsto\psi\circ\varphi$
 is continuous. Moreover, since $S$ is linear in the first variable, by Lemma 4.12, the map
-$\tilde{S}:L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{n}))\times C([0,1],\lfloor\overline{B_{2}(0)},B_{4}(0)\rfloor_{r})\to L^{p}([0,1],C^{r}(B_{2}(0),\mathbb{R}^{n})),\quad(\alpha,\beta)\mapsto S\circ(\alpha,\beta)$
+$\tilde{S}:L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{n}))\times C([0,1],\lfloor\overline{B_{2}(0)},B_{4}(0)\rfloor_{r})$ $\to L^{p}([0,1],C^{r}(B_{2}(0),\mathbb{R}^{m}))$
+$(\alpha,\beta)$ $\mapsto S\circ(\alpha,\beta)$
 is continuous. Hence, for $[\gamma]\in B_{L}$ we have
-$S\circ(\gamma,H([\gamma]))=\gamma\left(H(\gamma)\right)=\gamma.H([\gamma])\in L^{p}([0,1],C^{r}(B_{2}(0),\mathbb{R}^{n})).$
-This allow us to define the $L^{p}$-absolute continuous function $\tau:[0,1]\to C^{r}(B_{2}(0),\mathbb{R}^{n})$ given by
+$S\circ(\gamma,H([\gamma]))=\gamma\left(H(\gamma)\right)=\gamma.H([\gamma])\in L^{p}([0,1],C^{r}(B_{2}(0),\mathbb{R}^{m})).$
+This allow us to define the $L^{p}$-absolutely continuous function $\tau:[0,1]\to C^{r}(B_{2}(0),\mathbb{R}^{m})$ given by
 […]
-If we consider the point evaluation map $\varepsilon_{x}:C^{r}(B_{2}(0),\mathbb{R}^{n})\to\mathbb{R}^{n}$, $\varphi\mapsto\varphi(x)$, then for each $x\in B_{2}(0)$ we have
+If we consider the point evaluation map $\varepsilon_{x}:C^{r}(B_{2}(0),\mathbb{R}^{m})\to\mathbb{R}^{m}$, $\varphi\mapsto\varphi(x)$, then for each $x\in B_{2}(0)$ we have
 $\varepsilon_{x}\left(\tau(t)\right)=\varepsilon_{x}\left(H([\gamma])(t)\right),\quad\text{for all }t\in[0,1].$
-Since the family of maps $(\varepsilon_{x})_{x\in B_{2}(0)}$ separate points in $C^{r}(B_{2}(0),\mathbb{R}^{n})$, we have $\tau=H([\gamma])$.
+Since the family of maps $(\varepsilon_{x})_{x\in B_{2}(0)}$ separate points in $C^{r}(B_{2}(0),\mathbb{R}^{m})$, we have $\tau=H([\gamma])$.
 Hence $H[\gamma]$ is absolutely continuous and we can consider the map
@@
 Let $[\gamma]\in\mathcal{B}_{L}$, with $[\gamma_{i}]\in B_{L}$ for each $i\in\{1,...,n\}$.
-If $p\in M$, there is a chart $\kappa_{j}:U_{j}\to B_{5}(0)$ such that $p\in\kappa_{j}^{-1}(B_{3}(0))$ for some $j\in\{1,...,n\}$.
-Let $\zeta_{[\gamma],\kappa_{j}(p)}\in AC_{L^{P}}([0,1],C^{r}(B_{2}(0),B_{4}(0))$ be the solution of the integral equation
+If $m\in M$, there is a chart $\kappa_{j}:U_{j}\to B_{5}(0)$ such that $m\in\kappa_{j}^{-1}(B_{3}(0))$ for some $j\in\{1,...,n\}$.
+Let $\zeta_{[\gamma],\kappa_{j}(m)}\in AC_{L^{P}}([0,1],C^{r}(B_{2}(0),B_{4}(0))$ be the solution of the integral equation
 […]
-Then $\zeta_{[\gamma],\kappa_{j}(p)}$ is a $AC_{L^{p}}$-Carathéodory solution to
+Then $\zeta_{[\gamma],\kappa_{j}(m)}$ is a $AC_{L^{p}}$-Carathéodory solution to
 $x^{\prime}(t)$ $=\gamma_{j}(t)\left(x(t)\right),\quad t\in[0,1]$
-$x(0)$ $=\kappa_{j}(p).$
+$x(0)$ $=\kappa_{j}(m).$
 Hence the function
-$\eta_{[\gamma],p}:[0,1]\to M,\quad\eta_{[\gamma],p}(t):=\kappa_{j}^{-1}\circ\zeta_{[\gamma],\kappa_{j}(p)}(t)$
+$\eta_{[\gamma],m}:[0,1]\to M,\quad\eta_{[\gamma],m}(t):=\kappa_{j}^{-1}\circ\zeta_{[\gamma],\kappa_{j}(m)}(t)$
 is a $AC_{L^{p}}$-Carathéodory solution of the equation
 $y^{\prime}(t)$ $=\gamma(t)\left(y(t)\right),\quad t\in[0,1]$
-$y(0)$ $=p.$
-Since $\tilde{\gamma}$ satisfies a local $L^{1}$-Lipschitz condition, by Remark 7.3 the solution of this equation is unique and $\eta_{[\gamma],p}$ is well defined.
+$y(0)$ $=m.$
+Since $\tilde{\gamma}$ satisfies a local $L^{1}$-Lipschitz condition, by Remark 6.3 the solution of this equation is unique and $\eta_{[\gamma],m}$ is well defined.
 Moreover, $\tilde{\gamma}$ admits a global flow for initial time $t_{0}=0$, given by
-$F_{t,0}^{\overline{\gamma}}(p)=\eta_{[\gamma],p}(t)$
-for all $t\in[0,1]$, $p\in M$.
+$F_{t,0}^{\overline{\gamma}}(m)=\eta_{[\gamma],m}(t)$
+for all $t\in[0,1]$, $m\in M$.
 Following point 11.16 of [14], for each $i\in\{1,...,n\}$ we can construct an exponential map
@@
 $\kappa_{i}^{-1}\circ\text{exp}_{i}=\text{exp}\circ T\kappa_{i}^{-1}|_{\mathcal{D}_{i}}.$
-Moreover, for an open subset $\mathcal{O}_{i}\subseteq\mathcal{D}_{i}$ we consider the map
-$\psi_{i}:=(\text{pr}_{1},\text{exp}_{i})|_{\mathcal{O}_{i}}:\mathcal{O}_{i}\subseteq TB_{5}(0)\to\psi_{i}(\mathcal{O}_{i})\subseteq B_{5}(0)\times B_{4}(0).$
-Assuming that
+There is an open set $\mathcal{O}_{i}\subseteq\mathcal{D}_{i}$ containing $\overline{B_{4}(0)}\times\{0\}$ such that $(\text{pr}_{1},\exp_{i})(Q_{i})$ is open in $B_{5}(0)\times B_{5}(0)$ and the map
+$\psi_{i}:=(\text{pr}_{1},\text{exp}_{i})|_{\mathcal{O}_{i}}:\mathcal{O}_{i}\subseteq TB_{5}(0)\to\psi_{i}(\mathcal{O}_{i})\subseteq B_{5}(0)\times B_{4}(0)$
+is a $C^{\infty}$-diffeomorphism onto its image. Assuming that
 $T\kappa_{i}^{-1}(\mathcal{O}_{i})\subseteq\mathcal{W},$
@@
 $\{(x,x):x\in\overline{B_{5}(0)}\}\subseteq(\text{pr}_{1},\exp_{i})(\mathcal{O}_{i}),$
-there is exists $s_{i}\in]0,1]$ such that
+there exists $s_{i}\in]0,1]$ such that
 $\bigcup_{x\in B_{4}(0)}\{x\}\times B_{s_{i}}(x)\subseteq(\text{pr}_{1},\exp_{i})(\mathcal{O}_{i})$
-is a smooth diffeomorphism and there exists a $s_{i}\in]0,1[$, such that $\psi_{i}^{-1}$ restrict to a smooth diffeomorphism of the form
+is a smooth diffeomorphism and there exists an $s_{i}\in]0,1[$, such that $\psi_{i}^{-1}$ restrict to a smooth diffeomorphism of the form
 $(\text{Id}_{B_{4}(0)},\theta_{i}):\bigcup_{x\in B_{4}(0)}\{x\}\times B_{s_{i}}(x)\subseteq B_{5}(0)\times B_{5}(0)\to\psi_{i}^{-1}\left(\bigcup_{x\in B_{4}(0)}\{x\}\times B_{s_{i}}(x)\right)\subseteq\mathcal{O}_{i}$
 with open image in $\mathcal{O}_{i}$. For each $x\in B_{4}(0)$ we define the set
-$\mathcal{O}_{i,x}:=\{y\in\mathbb{R}^{n}:(x,y)\in\mathcal{O}_{i}\}.$
+$\mathcal{O}_{i,x}:=\{y\in\mathbb{R}^{m}:(x,y)\in\mathcal{O}_{i}\}.$
 Hence, we have
@@
 $Z_{i,x}:=\left\{\varphi\in C^{r}(B_{2}(0),\mathbb{R}^{n}):(\forall x\in\overline{B_{1}(0)})\text{ }\varphi(x)\in B_{s_{i}}(x)\right\}$
-is open in $C^{r}(B_{2}(0),\mathbb{R}^{n})$ and the map
-$(\theta_{i})_{*}:Z_{i,x}\subseteq C^{r}(B_{2}(0),\mathbb{R}^{n})\to C^{r}(B_{1}(0),\mathbb{R}^{n}),\quad\varphi\mapsto(\theta_{i})_{*}(\varphi):=\theta_{i}\circ(\text{id}_{B_{1}(0)},\varphi)$
-is smooth since $\theta$ is smooth [12, Proposition 4.23].
+is open in $C^{r}(B_{2}(0),\mathbb{R}^{m})$ and the map
+$(\theta_{i})_{*}:Z_{i,x}\subseteq C^{r}(B_{2}(0),\mathbb{R}^{m})\to C^{r}(B_{1}(0),\mathbb{R}^{m}),\quad\varphi\mapsto(\theta_{i})_{*}(\varphi):=\theta_{i}\circ(\text{id}_{B_{1}(0)},\varphi)$
+is smooth since $\theta$ is smooth [16, Proposition 4.23].
 Let $\Psi:\mathcal{U}\to\mathcal{V}$ be a chart on $\text{id}_{M}\in\text{Diff}^{r}(M)$, then since $\rho_{1}$ is a topological embedding we can assume that
 $\mathcal{V}:=\rho_{1}^{-1}\left(\prod_{i=1}^{n}\mathcal{V}_{i}\right)$
-for suitable open 00-neighborhoods $\mathcal{V}_{i}\subseteq C^{r}(B_{1}(0),\mathbb{R}^{n})$.
+for suitable open 00-neighborhoods $\mathcal{V}_{i}\subseteq C^{r}(B_{1}(0),\mathbb{R}^{m})$.
 Since $(\theta_{i})_{*}(\text{id}_{B_{2}(0)})=0$, by continuity of $(\theta_{i})_{*}$ there exists open $\text{id}_{B_{2}(0)}$-neighborhoods $\mathcal{Y}_{i}\subseteq Z_{i,x}$ such that
@@
 If
-$\mathcal{P}:=(R_{\text{id}_{M},t})^{-1}\left(\prod_{i=1}^{n}\mathcal{P}_{i}\right)$
+$\mathcal{P}:=R_{5}^{-1}\left(\prod_{i=1}^{n}\mathcal{P}_{i}\right)$
 then $\mathcal{P}$ is an open 00-neighborhood in $L^{p}([0,1],\mathcal{X}^{r}(M))$ with $\mathcal{P}\subseteq\mathcal{B}_{L}$.
 This allow us to do the composition
-$(\theta_{i})_{*}\circ H:B_{L}\subseteq L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{n})\to C([0,1],C^{r}(B_{1}(0),\mathbb{R}^{n})$
+$(\theta_{i})_{*}\circ H:B_{L}\subseteq L^{p}([0,1],C^{r}(B_{5}(0),\mathbb{R}^{m})\to C([0,1],C^{r}(B_{1}(0),\mathbb{R}^{m})$
 Following the same steps of point $1.18$ and $1.19$ of [14], we have that for each $[\gamma]\in\mathcal{P}$, there exists an unique $\theta_{[\gamma]}\in C([0,1],\mathcal{X}^{r}(M))$ such that
@@
 $\Psi^{-1}(\theta_{[\gamma]}(t))=\exp\circ\theta_{[\gamma]}(t)\in\text{Diff}^{\text{ }r}(M),\quad\text{ for all }t\in[0,1].$
-Let $p\in M$ with $x:=\kappa_{i}(p)$, then we have
-$\exp\circ\theta_{[\gamma]}(t)(p)$ $=\left(\kappa_{i}^{-1}\circ\exp_{i}\circ T\kappa_{i}\right)\circ\theta_{[\gamma]}(t)(\kappa_{i}^{-1}(x))$
+Let $m\in M$ with $x:=\kappa_{i}(m)$, then we have
+$\exp\circ\theta_{[\gamma]}(t)(m)$ $=\left(\kappa_{i}^{-1}\circ\exp_{i}\circ T\kappa_{i}\right)\circ\theta_{[\gamma]}(t)(\kappa_{i}^{-1}(x))$
 $=\kappa_{i}^{-1}\circ\exp_{i}\circ(\theta_{i})_{*}(H([\gamma_{i}]))(t)(x)$
@@
 $=\kappa_{i}^{-1}\circ\zeta_{[\gamma_{i}],x}(t)$
-$=F_{t,0}^{\tilde{\gamma}}(p).$
+$=F_{t,0}^{\tilde{\gamma}}(m).$
 Hence, we define
@@
 Since the exponential map $\exp$ is smooth, we need to show that $\theta_{[\gamma]}$ is absolutely continuous, but this is equivalent to show that
-$[0,1]\to\prod_{i=1}^{n}C^{r}(B_{1}(0),\mathbb{R}^{n}),\quad t\mapsto\rho_{1}\circ\theta_{[\gamma]}(t)$
-But this maps coincide with the map
-$[0,1]\to\prod_{i=1}^{n}C^{r}(B_{1}(0),\mathbb{R}^{n}),\quad t\mapsto\left((\theta_{i})_{*}\left(H([\gamma_{i}])\right)\right)_{i=1}^{n}(t).$
-is absolutely continuous.
-Since $(\theta_{i})_{*}$ is smooth and $H([\gamma_{i}])\in AC_{L^{p}}([0,1],C^{r}(B_{2}(0),B_{4}(0)))$, by Lemma 2.16, the composition and hence the product are absolutely continuous.
+$[0,1]\to\prod_{i=1}^{n}C^{r}(B_{1}(0),\mathbb{R}^{m}),\quad t\mapsto\rho_{1}\circ\theta_{[\gamma]}(t).$
+But this map coincides with the map
+$[0,1]\to\prod_{i=1}^{n}C^{r}(B_{1}(0),\mathbb{R}^{m}),\quad t\mapsto\left((\theta_{i})_{*}\left(H([\gamma_{i}])\right)\right)_{i=1}^{n}(t)$
+which is absolutely continuous.
+Since $(\theta_{i})_{*}$ is smooth and $H([\gamma_{i}])\in AC_{L^{p}}([0,1],C^{r}(B_{2}(0),B_{4}(0)))$, by Proposition 3.8, the composition and hence the product are absolutely continuous.
 Then we have that
@@
 $\eta_{[\gamma]}\in AC_{L^{p}}([0,1],\text{Diff}^{\text{ }r}(M)).$
-In consequence, Lemma 5.9 and Remark 7.3, the right half-Lie group $\text{Diff}^{\text{ }r}(M)$ is $L^{p}$-semiregular. Since the map
+In consequence, by Lemma 4.9 and Remark 6.3, the right half-Lie group $\text{Diff}^{\text{ }r}(M)$ is $L^{p}$-semiregular. Since the map
 $B_{L}\to\prod_{i=1}^{n}C([0,1],C^{r}(B_{1}(0),\mathbb{R}^{n})),\quad[\gamma_{i}]\mapsto\left((\theta_{i})_{*}\left(H([\gamma_{i}])\right)\right)_{i=1}^{n}$
@@
 is continuous. ∎
+Now we will focus in the $L^{p}$-semiregularity of the case $\text{Diff}^{r}(M)$ with $M$ a compact smooth manifold with boundary.
+Definition 6.6
+Let $M$ and $N$ be smooth manifolds with boundaries and suppose ${f:\partial M\to\partial N}$ is a diffeomorphism, We define adjunction space $M\cup_{f}N$ as the set formed identifying each $p\in\partial M$ with $f(p)\in\partial N$.
+We recall [26, Theorem 9.29].
+Theorem 6.7
+Let $M$ and $N$ be a smooth manifolds with boundaries and ${f:\partial M\to\partial N}$ be a diffeomorphism.
+Then the adjunction space $M\cup_{f}N$ is a topological manifold without boundary which has a smooth manifold structure such that there are regular domains $M^{\prime},N^{\prime}\subseteq M\cup_{f}N$ diffeomorphic to $M$ and $N$, respectively, such that $M^{\prime}\cup N^{\prime}=M\cup_{f}N$ and ${M^{\prime}\cap N^{\prime}=\partial M^{\prime}=\partial N^{\prime}}$.
+Moreover, $M$ and $N$ are both compact if and only if $M\cup_{f}N$ is compact.
+Definition 6.8
+Let $M$ be a smooth manifold with boundary. If $M^{\prime}$ denotes a copy of $M$, we define the double of $M$ as the smooth manifold without boundary
+$DM=M\sqcup_{\text{id}_{\partial}}M^{\prime}$ (6.4)
+where $\text{id}_{\partial}:\partial M\to\partial M^{\prime}$ is the identity map.
+For $p\in M$, we denote by $(p,0)$ and $(p,1)$ the corresponding element in $M$ and $M^{\prime}$, respectively, and if $p\in\partial M$, then $(p,0)\sim(p,1)$. By Theorem 6.7, the map
+$\rho:M\to DM,\quad p\mapsto[(p,0)]$
+is an embedding onto a regular domain of $DM$ which we identify with $M$.
+Definition 6.9
+Let $r\in\mathbb{N}\cup\{0,\infty\}$ and $M$ be a smooth manifold with boundary. We define the vector space of $C^{r}$-stratified vector field on $M$ as
+$\mathcal{X}_{str}^{r}(M):=\{X\in\mathcal{X}^{r}(M):X(\partial M)\subseteq T\partial M\}.$
+We recall [17, Colorally 1.8].
+Proposition 6.10
+For each $k\in\mathbb{N}\cup\{0,\infty\}$, $n\in\mathbb{N}$, $m\in\{0,...,d\}$ and locally convex space $F$, the restriction map
+$\mathcal{E}:C^{k}(\mathbb{R}^{d},F)\to C^{k}([0,\infty)^{m}\times\mathbb{R}^{d-m},F)$
+has a continuous linear right inverse. Moreover, the restriction map
+$C^{k}(\mathbb{R}^{d},F)\to C^{k}([0,1]^{d},F)$
+has continuous linear right inverse.
+Proposition 6.11
+Let $r\in\mathbb{N}\cup\{0,\infty\}$, $m\in\mathbb{N}$ and $M$ be an $m$-dimensional compact smooth manifold with boundary. Then there exists an extension map
+$\alpha:\mathcal{X}_{str}^{r}(M)\to\mathcal{X}^{r}(DM),\quad X\mapsto\alpha(X)$
+which is continuous linear.
+Proof.
+By compactness of $M$, we find charts $\varphi_{i}:U_{i}\to V_{i}$ of $M$ around points $p_{i}\in\partial M$ such that $(U_{i})_{i=1}^{k}$ is a finite open cover of $\partial M$ which extends to charts $\widetilde{\varphi_{i}}:\widetilde{U_{i}}\to\widetilde{V_{i}}$ around $p_{i}$.
+For $X\in\mathcal{X}_{str}^{r}(M)$, we write
+$Y_{i}:=d\varphi_{i}\circ X\circ\varphi_{i}^{-1}:V_{i}\subseteq[0,\infty)\times\mathbb{R}^{m-1}\to\mathbb{R}^{m}.$
+Without loss of generality, we assume that $V_{i}=[0,\infty)\times\mathbb{R}^{m-1}$ and we consider the extension $\widetilde{Y}_{i}:\mathbb{R}^{m}\to\mathbb{R}^{m}$ of $Y_{i}$ given by Proposition 6.10.
+We define
+$\widetilde{X}_{i}:=T\widetilde{\varphi_{i}}^{-1}\circ\widetilde{Y}_{i}\circ\widetilde{\varphi_{i}}:\widetilde{U_{i}}\to T\widetilde{U_{i}}.$
+Then $\widetilde{X}_{i}\in\mathcal{X}^{r}(\tilde{U_{i}})$ and the map
+$\Phi_{1}:\mathcal{X}_{str}^{r}(M)\to\prod_{i=1}^{k}\mathcal{X}^{r}(\widetilde{U_{i}}),\quad X\mapsto(\widetilde{X}_{i})_{i=1}^{k}$
+is continuous linear. Let us consider $\widetilde{U}_{k+1}:=M^{\circ}$, $\widetilde{U}_{k+2}:=(M^{\prime})^{\circ}$ and the open cover of $DM$
+$\mathcal{A}:=\{\widetilde{U_{1}},...,\widetilde{U_{k}},\widetilde{U}_{k+1},\widetilde{U}_{k+2}\}.$
+Then there exists a partition of the unity subordinate to $\mathcal{A}$, denoted by $\{h_{1},...,h_{k},h_{k+1},h_{k+2}\}$, such that $\text{supp}(h_{i})\subseteq\widetilde{U_{i}}$ for each $i\in\{1,...,k,k+1,k+2\}$.
+Denoting the space of all $C^{k}$-vector fields of $\widetilde{U}_{i}$ with support on $\text{supp}(h_{i})$ as $\mathcal{X}_{i}^{r}(\widetilde{U}_{i})$, we see that the map
+$\mathcal{X}^{r}(\widetilde{U}_{i})\to\mathcal{X}_{i}^{r}(\widetilde{U}_{i}),\quad Y\mapsto h_{i}Y$
+is continuous linear.
+Moreover, if $Z\in\mathcal{X}_{i}^{r}(\widetilde{U}_{i})$, we can obtain an extension $\mathcal{E}_{i}(Z)\in\mathcal{X}^{r}(DM)$ of $Z$ by extending it by 00, and a continuous linear map
+$\mathcal{E}_{i}:\mathcal{X}_{i}^{r}(\widetilde{U}_{i})\to\mathcal{X}^{r}(DM),\quad Z\mapsto\mathcal{E}_{i}(Z).$
+For $X\in\mathcal{X}_{str}^{r}(M)$, we write $\widetilde{X}_{k+1}:=X|_{M^{\circ}}$. This enables us to define the extension map
+$\alpha:\mathcal{X}_{str}^{r}(M)\to\mathcal{X}^{r}(DM),\quad X\mapsto\alpha(X):=\sum_{i=1}^{k+1}\mathcal{E}_{i}(h_{i}\widetilde{X}_{i})$
+which is continuous linear. ∎
+Remark 6.12
+Let $M$ be a compact smooth manifiold with boundary and $r\in\mathbb{N}\cup\{\infty\}$.
+By [20, Proposition 1.3], the set $\text{Diff}^{r}(M)$ is a embedded submanifold of $C^{r}(M,DM)$.
+This allows us to consider the inclusion map restricted onto its image
+$J:\text{Diff}^{r}(M)\to J(\text{Diff}^{r}(M))\subset C^{r}(M,DM),\quad\phi\mapsto\phi$ (6.5)
+then $J$ is a diffeomorphism. Since for each $g\in C^{r}(M,M)$ fixed, the right translation map
+$\rho_{C^{r}}(g):C^{r}(M,DM)\to C^{r}(M,DM),\quad\phi\mapsto\phi\circ g$
+is smooth. For each $g\in\text{Diff}^{r}(M)$ fixed, the right translation map
+$\rho(g):\text{Diff}^{r}(M)\to\text{Diff}^{r}(M),\quad\phi\mapsto\phi\circ g$
+can be written as
+$\rho(g)=J^{-1}\circ\rho_{C^{r}}(J(g))\circ J.$
+Hence $\text{Diff}^{r}(M)$ is a right half-Lie group.
+Theorem 6.13
+Let $M$ be a compact smooth manifold with boundary, $r\in\mathbb{N}\cup\{\infty\}$ and $1\leq p<\infty$.
+Then the right half-Lie group $\text{Diff}^{\text{ }r}(M)$ is $L^{p}$-semiregular and the evolution map with continuous values
+$\operatorname{Evol}_{C}:L^{p}([0,1],T_{e}\text{Diff}^{r}(M))\to C([0,1],\text{Diff}^{r}(M)),\quad[\gamma]\mapsto\operatorname{Evol}_{C}(\gamma)$
+is continuous.
+Proof.
+Let consider the map $\alpha:\mathcal{X}_{str}^{r}(M)\to\mathcal{X}^{r}(DM)$ as Proposition 6.11, we define the map
+$\tilde{\alpha}:=L^{p}([0,1],\alpha):L^{p}([0,1],\mathcal{X}_{str}^{r}(M))\to L^{p}([0,1],\mathcal{X}^{r}(DM)),\quad[\gamma]\mapsto[\alpha\circ\gamma];$
+which is linear and continuous. Since $DM$ is a compact smooth manifold without boundary, the right half-Lie group $\text{Diff}^{r}(DM)$ is $L^{p}$-semiregular with continuous evolution map denoted by
+$\operatorname{Evol}_{DM}:L^{p}([0,1],\mathcal{X}^{r}(DM))\to AC_{L^{p}}([0,1],\text{Diff}^{r}(DM)),\quad[\gamma]\mapsto\operatorname{Evol}_{DM}(\gamma).$
+Let $[\gamma]\in L^{p}([0,1],\mathcal{X}_{str}^{r}(M))$, we define the absolutely continuous function ${\xi_{\gamma}:[0,1]\to\text{Diff}^{r}(DM)}$ by $\xi_{\gamma}:=\operatorname{Evol}_{DM}(\tilde{\alpha}([\gamma]))$.
+Then $\xi_{\gamma}$ is the solution of the equation
+$\dot{\xi}_{\gamma}(t)$ $=\tilde{\alpha}([\gamma])(t).\xi_{\gamma}(t),\quad t\in[0,1]$
+$\xi_{\gamma}(0)$ $=e.$
+For $[\gamma]$ close to 00, the proof of Theorem 6.5 shows that, for each $p\in M$, the function $x_{p}:[0,1]\to DM$, given by ${x_{p}(t):=\xi_{\gamma}(t)(p)}$ is a solution of the equation
+$\dot{x_{p}}(t)$ $=\tilde{\alpha}([\gamma])(t)\circ x_{p}(t),\quad t\in[0,1]$
+$x_{p}(0)$ $=p,$
+and this ODE satisfies local uniqueness of Caratheodory solutions.
+Looking at the compact manifold $\partial M$ without boundary and the vector fields $\gamma(t)|_{\partial M}\in\mathcal{X}^{r}(\partial M)$ we likewise get a solution $y_{p}:[0,1]\to\partial M$ for each $p\in\partial M$, for the differential equation
+$\dot{y_{p}}(t)$ $=\tilde{\alpha}([\gamma])(t)\circ y_{p}(t),\quad t\in[0,1]$
+$y_{p}(0)$ $=p.$
+Then $y_{p}$ also solves the initial value problem for $x_{p}$, whence $x_{p}=y_{p}$ by local uniqueness.
+In consequence $x_{p}([0,1])\subseteq\partial M$.
+This implies that for each $t\in[0,1]$ fixed, we have $\xi_{\gamma}(t)(M)\subseteq M$, $\xi_{\gamma}(t)(M^{\prime})\subseteq M^{\prime}$ and $\xi_{\gamma}(t)(\partial M)\subseteq\partial M$.
+Therefore, we obtain ${\xi_{\gamma}(t)|_{M}\in\text{Diff}^{r}(M)}$.
+Consider the smooth embedding $\iota:\text{Diff}^{r}(DM)\to C^{r}(DM,DM)$, $\phi\mapsto\phi$. Then the map
+$\tilde{\iota}:=AC_{L^{p}}([0,1],\iota):AC_{L^{p}}([0,1],\text{Diff}^{r}(DM))\to AC_{L^{p}}([0,1],C^{r}(DM,DM)),\quad\eta\mapsto\iota\circ\eta$
+is smooth. Let $\rho:M\to DM$ be the inclusion map, which is smooth and a diffeomorphism onto its image, then the maps
+$\rho^{*}:=C^{r}(\rho,DM):C^{r}(DM,DM)\to C^{r}(M,DM),\quad\phi\mapsto\phi\circ\rho=\phi|_{M}$
+and
+$\tilde{\rho}:AC_{L^{p}}([0,1],C^{r}(DM,DM))\to AC_{L^{p}}([0,1],C^{r}(M,DM)),\quad\eta\mapsto\rho^{*}\circ\eta$
+are smooth.
+Consider the restricted inclusion map $J$ as in equation (6.5), by Lemma 2.25, we define the map
+$\tilde{j}:AC_{L^{p}}([0,1],J(\text{Diff}^{r}(M)))\to AC_{L^{p}}([0,1],\text{Diff}^{r}(M)),\quad\eta\mapsto J^{-1}\circ\eta.$
+The fact that $\xi_{\gamma}(t)|_{M}\in\text{Diff}^{r}(M)$ for each $t\in[0,1]$, enables us to define the function
+$\zeta_{\gamma}:[0,1]\to C^{r}(M,DM),\quad\zeta_{\gamma}(t):=(\tilde{j}\circ\tilde{\rho}\circ\tilde{\iota})(\xi_{\gamma})(t)$
+which is in $AC_{L^{p}}([0,1],\text{Diff}^{r}(M))$. Moreover, by definition of $\alpha$, we have
+$\tilde{\alpha}([\gamma])(t)\circ\zeta_{\gamma}(t)=\gamma(t)\circ\zeta_{\gamma}(t).$
+Hence $\zeta_{\gamma}$ verifies the equation
+$\dot{\zeta_{\gamma}}(t)$ $=\gamma(t).\zeta_{\gamma}(t),\quad t\in[0,1]$
+$\zeta_{\gamma}(0)$ $=e.$
+Therefore $\text{Diff}^{r}(M)$ is $L^{p}$-semiregular and the evolution map is given by
+$\operatorname{Evol}:L^{p}([0,1],\mathcal{X}_{str}^{r}(M))\to AC_{L^{p}}([0,1],\text{Diff}^{r}(M)),\quad[\gamma]\mapsto\Big{(}\tilde{j}\circ\tilde{\rho}\circ\tilde{\iota}\circ\operatorname{Evol}_{DM}\circ\tilde{\alpha}\Big{)}([\gamma]).$
+We consider the inclusion map
+$\omega:AC_{L^{p}}([0,1],C^{r}(M,DM))\to C([0,1],C^{r}(M,DM)),\quad\eta\mapsto\eta$
+which is smooth, and we define the smooth map
+$\widetilde{\rho}_{C}:AC_{L^{p}}([0,1],C^{r}(DM,DM))\to C([0,1],C^{r}(M,DM)),\quad\eta\mapsto\omega(\rho^{*}\circ\eta).$
+We write
+$\widetilde{j}_{C}:C([0,1],J(\text{Diff}^{r}(M)))\to C([0,1],\text{Diff}^{r}(M)),\quad\eta\to J^{-1}\circ\eta$
+then $\widetilde{j}_{C}$ is continuous. Therefore, the evolution map with continuous values $\operatorname{Evol}_{C}$ is given by
+$\operatorname{Evol}_{C}:L^{p}([0,1],\mathcal{X}_{str}^{r}(M))\to C([0,1],\text{Diff}^{r}(M)),\quad[\gamma]\mapsto\Big{(}\widetilde{j}_{C}\circ\widetilde{\rho}_{C}\circ\widetilde{\iota}\circ\operatorname{Evol}_{DM}\circ\widetilde{\alpha}\Big{)}([\gamma])$
+and is continuous since is composition of continuous maps. ∎
 Acknowledgements: The author would like to thank Helge Glöckner for his guidance during the development of this work.
-The author was financially supported by ANID and DAAD (DAAD/BecasChile 2020, ID:91762237/62190017).
+The author was partially supported by the FONDECYT Grant #1241719 and by ANID and DAAD (DAAD/BecasChile 2020, ID:91762237/62190017).